A Complete Bibliography of Publications in Isis, 1920–1929

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.05

Title word cross-reference

≫ [616, 792, 556]. ≪ [616, 792, 556]. \underline{n} [69].

-1335 [197].

1106 [459]. 1491 [416]. 1493 [561]. 1498 [562]. 1524 [103]. 1614 [41]. 1647 [662]. 1650 [104]. 1679 [42]. 1693 [43]. 16e [102]. 1721 [452]. 1724 [435]. 1726 [44]. 17e [102]. 1913 [164]. 1922 [368, 241, 284]. 1923 [332]. 1925 [593]. 1928-1 [745]. 1928-2 [781].

20th [633]. 22nd [435]. 27/11 [770].

3000 [182].

4/1 [770].

= [69].

Îles [491].

A. [544, 482, 222, 82, 104]. A.-A [18]. A.-B [18]. A.D [834]. A.D. [754]. Abû [730]. Abbé [104]. Abbildungen [837, 275, 181]. Abbott [351]. ’Abd [560, 116]. ’abd-al-Kahir [116]. Abdel [670]. Abdou [670]. Abeilles [523]. Abel [328, 678]. Abenbeder [207]. Abendlandes [271]. Aberglaube [628]. Aboriginal [627]. Abr [796]. Abu [116, 731, 577]. Abu-Mansur [116]. Abu’l [450, 441, 594, 441]. Abu’l-Jud [441, 594, 441]. Abu’l-Qasim [450]. Abu’l=Jud [594]. accidentium [835]. according [641, 290]. Account [551]. Achalme [226, 315]. Achard [570]. Achille [849]. achtzehnten [238]. Acquaintance [776]. Adams’s [171]. adaptation [311]. adh [450]. adh-dhahab [450]. Adickes [564]. Adolphe [431]. Adrien [142]. adversaires [615]. adversaries [615]. Ae [544]. Aegypter [613]. aeronautics [496]. Affinitätsproblem [512]. affinity [512]. After [435, 198]. Age [213, 397]. Ages [249, 267, 669]. Agnes [458, 459]. agnostiquer [428]. Agostino [582]. Agricola [837]. Ahmad [450, 731]. Ahmed [203]. ai [272, 138]. Aime [406]. Al [116, 836, 442, 670, 624, 203, 450, 560, 730, 247, 594, 377, 577, 450, 731, 441, 204, 732, 730, 731]. al-Andalusi [560]. al-Bîrûnî [731]. Al-Biruni [247]. al-Falasifat [732]. Al-Fark [116]. al-Firak [116]. al-Haramain [624]. al-’ilm [450]. al-’Iraqi [450, 577]. al-Jâhiz [730]. al-Jahiz [730]. Al-Khowarizmi [204]. al-Khwarizmi [203]. al-Lith [594, 441]. al-Madjriti [203]. al-Majriti [377]. al-Mas‘ûdî [731]. al-muktasab [450]. al-Qânûn [731]. al-Rahman [560]. al-Razi [442]. al-Razi’s [247]. Albert [648, 364]. Albrecht [390]. Alchemical [800]. Alchemie [310, 579, 217, 54]. Alchemist [551]. Alchemisten [448]. alchemists [198]. Alchemy [720, 445, 827, 579, 813, 664]. Alchimiques [799, 491]. Alchimistes [198]. Aldo [269, 138, 170]. Alemanus [416]. Alexander [47]. Alfred [444, 524, 99, 215, 567, 430, 539, 628]. Algazel [732]. Algebra [118, 204, 207]. Algorism [749, 805]. Algorisms [68]. ’Ali [560, 412, 638]. aliis [835]. Allbutt [174]. alle [837]. Alleged [637]. allemande [321]. Allen [816]. allesten [628]. Allier [818]. along [300]. Alpes [350]. Alsadiq [448]. Altai [529]. Altai-Iran [529]. alte [211]. alten [467, 613]. Altertum [129, 524]. Altertumer [124]. Altertums [166, 364]. Altertumskunde [114]. alteste [217]. altesten [118, 570]. altester [81]. aluminibus [785]. am [837, 647]. Ambroise [160]. Ambrose [346]. America [555, 656, 627, 205, 798]. American [356, 489, 627]. Among [239, 433]. Amori [705]. Amr [730]. ämter [837]. Anachronism [33]. Analyse [132]. Analysis [133]. Analyticae [766]. analytique [457]. Ananda [347]. Anatomic [175]. Anatomical [267, 669]. anatomie [680, 796]. Anatomy [743]. Ancien [668]. anciens [461, 218, 483]. Ancient [505, 264, 111, 613, 17, 527]. Andalusi [560]. anderen [330]. Anderson [491, 800]. Andoyer [260]. Andre [616, 422, 322, 683, 761]. Anecdota [812].

Anfangen [349]. Angabe [837]. Angelique [461]. Angle [822]. Angle-Geometry [822]. Angleterre [144]. Anglo [621]. Anglo-Norman [621]. Anianus [772]. animale [311]. Animate [125]. Annales [531]. Annie [491, 800]. Annual [802]. ans [8]. Anthiaume [104]. Anthony [311]. antica [584]. Antike [76]. Antiken [842]. Antique [668, 667]. Antiquity [742]. Antoine [665]. Anton [261]. Ants [665]. août [151]. Apercu [21]. Apert [277]. Appeal [546]. Appendice [807]. Appendix [807]. Appert [49]. Applications [97]. April [435, 147, 63, 145]. aptitudes [428]. Aquiculturist [378]. Arab [594, 788, 441, 506]. arabe [788, 845]. Araber [248]. arabes [222, 421, 740]. Arabi [412]. Arabian [586, 171]. Arabic [484, 411, 719, 804, 224, 267]. arabicae [671]. Arabische [506, 448]. arabischen [594, 118, 292, 441]. Arabs [248]. Arbeiter [522]. Archeologie [9]. Archibald [453]. Archimedes [274]. Architect [389]. Arguments [3]. Arin [117]. Aristoteles [569]. Aristotelis [34]. Arithmetic [656, 379, 507, 540, 620]. Arithmetical [722]. Arithmetique [616]. Armand [812]. Arnold [660]. Arranged [356]. Arrow [33]. Arsenal [505]. Art [180, 97, 459, 305]. Arthur [329, 14, 15, 19, 220, 467, 407, 125]. Article [644, 658, 728, 790, 809]. Artis [766]. Arts [347, 175]. Ascribed [720, 834]. Ashton [625]. Asia [452]. Asin [701]. Aspect [657]. Aspects [264]. Association [773]. Assyrian [585]. Assyrien [526]. Astrolab [638, 748]. astrolabe [748, 638]. Astrologie [563]. Astrometeorologie [103]. Astronomen [731]. Astronomia [584]. Astronomical [567, 248]. Astronomie [380, 845, 839]. astronomischen [203, 248]. Astronomy [848]. Athelhard [203]. atheniensa [812]. Atlantic [266]. atmospherique [279]. Atomen [126]. Atomic [642, 726]. atomiques [424]. Atomism [220]. Atomistik [673]. Atomists [843]. Atra [556]. Attempt [596, 388]. auch [837]. Auerbach [402]. Aufgaben [594, 441]. aufs [837]. Aufsatze [16]. Augen [837]. Augenarztes [647]. augmenter [23]. August [692]. Augustana [702]. Auguste [433]. Augustus [40]. aultres [563]. Aurelius [254]. Ausbreitung [54]. Auspices [504]. Ausschluss [261]. Auteurs [421]. Author [333, 286, 372, 854, 499, 599, 607, 631, 780, 473, 820, 685, 715, 763]. Authors [595, 436, 687, 824]. Autobiography [551, 453]. aux [144, 35]. Avalon [14, 15]. avecq [563]. Avicennae [755]. Ayer [711]. Azo [754].

B [42, 367, 164, 279, 33, 40, 43, 44, 670, 18, 403, 534, 392, 252, 759]. Bîrûnî [731]. Babcock [266]. Babylonien [526]. Back [25, 64, 91, 153, 195, 243, 287, 331, 334, 369, 373, 409, 434, 437, 471, 474, 497, 500, 543, 574, 600, 609, 632, 654, 686, 689, 855]. Background [6]. Bacon [757, 758]. Baconi [835]. Baddeley [128]. Baerwald [605]. Baghdadi [116]. Bagnell [185]. Bahr [730]. Bailey [843]. Bailly [774]. Bain [116]. Bakhshâlî [795]. Ball [554]. Ballore [536, 399]. Band [532, 733]. Banerjee [111]. Baptista [661]. Barclay [304]. Barnabas [550]. Barnacles [816]. Barnes [571]. Baron [265]. Barrow [39]. Barth [8]. Bashford [407]. Bassi [582, 582]. bassin [108]. Bath [203]. Baumstark [261]. Becker [420].

Béclard [744]. Bedeutung [202]. before [800]. Begrunder [640]. Behandlung [548]. Behaviorism [358]. Being [504, 116, 562]. Beitrag [530]. Beitrage [166, 292, 360]. Beleuchtung [794]. Belgique [216]. Belle [757]. Benedetto [756]. Benedict [345]. Benjamin [391]. Benoy [6]. Bensaude [72]. Berg [837, 776]. Berg- [837]. Bergiers [563]. Bergson [427]. Bergund [837]. Bernard [77]. Bernhard [844]. Bernoullis [274]. Berr [419]. Berriedale [220]. Berthelot [703, 411]. Berthold [739, 53]. Bertipaglia [549]. Bertrand [61, 741, 539]. Berucksichtigung [273]. beschrieben [837]. besonderer [273]. Besredka [214]. Besthorn [203]. Bestow [626]. Beta [570]. betrachtet [512]. between [751]. Bezeichnungen [837]. Beziehungen [330]. Bezold [677]. Bhandarkar [13]. Bi [717]. Bi-Centenary [717]. bibliografia [308]. Bibliographical [356]. Bibliographie [150, 151, 28, 648, 425]. Bibliographies [149, 89]. bibliographique [298]. Bibliographisher [135]. Bibliography [147, 368, 598, 333, 286, 372, 473, 175, 63, 27, 22, 194, 88, 145, 241, 332, 284, 407, 371, 472, 498, 593, 606, 630, 684, 779, 714, 762, 819, 853, 820, 854, 499, 599, 607, 631, 685, 715, 763, 780, 28]. Biblioteca [846]. Bibliotheca [137]. Bibliothek [365]. Bibliothèque [516]. Bicentenary [773]. Bidez [799, 799]. Bijdrage [569]. Bijdragen [102]. Bilancioni [324]. biochimiques [21]. Biografías [222]. Biography [146, 596, 285, 408, 306, 388]. Biologie [20, 311, 431]. biologique [21, 176, 423]. biologiques [20]. Biology [87, 815, 588]. Birth [435]. Biruni [247]. bis [673, 349, 570, 81, 238, 628]. Bjornbo [203]. Blaise [648, 694]. Blanchet [107]. Blaringhem [50, 393]. Bloch [592, 495]. Blood [280]. Blutezeit [103]. Bodensee [647]. Boerhaave [696]. Bohn [176]. Boll [283, 677]. Bolzano [769]. Book [698, 744, 775, 794, 842, 163, 164, 177, 234, 303, 316, 490, 522, 704, 737, 208, 258, 300, 540, 519, 306, 85, 132, 133, 353, 492, 664, 571, 625, 663, 673, 760, 535, 362, 405, 457, 461, 523, 536, 541, 542, 582, 584, 589, 591, 622, 629, 648, 703, 738, 774, 20, 21, 49, 50, 51, 73, 84, 86, 127, 139, 140, 141, 176, 201, 214, 226, 227, 236, 237, 259, 260, 268, 276, 277, 279, 281, 283, 298, 301, 302, 311, 315, 321, 323, 327, 328, 329, 352, 387, 393, 398, 399, 419, 423, 424, 426, 428, 431, 515, 537, 587]. Book [709, 205, 178, 217, 33, 37, 38, 39, 40, 41, 43, 44, 45, 46, 56, 60, 61, 118, 272, 798, 170, 469, 402, 741, 305, 312, 813, 743, 460, 317, 358, 180, 187, 202, 231, 117, 171, 206, 222, 223, 224, 421, 422, 531, 670, 732, 6, 7, 8, 9, 12, 13, 14, 15, 16, 17, 18, 19, 112, 113, 114, 115, 220, 221, 263, 264, 403, 433, 456, 521, 534, 583, 680, 735, 849, 190, 322, 326, 494, 518, 675, 681, 683, 713, 734, 761, 818, 527, 624, 119, 120]. Book [121, 122, 123, 705, 707, 10, 514, 605, 203, 204, 207, 261, 265, 271, 278, 310, 341, 349, 394, 395, 420, 467, 470, 524, 525, 526, 529, 530, 532, 570, 572, 48, 59, 62, 71, 175, 182, 183, 189, 72, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 87, 100, 101, 102, 103, 104, 111, 124, 125, 126, 129, 160, 167, 168, 216, 262, 266, 269, 275, 280, 282, 299, 304, 308, 309, 324, 325, 340, 346, 347, 350, 351, 354, 355, 359, 360, 361, 392, 425, 463, 466, 739, 742, 796, 839, 840, 845, 846, 847, 34, 47, 53, 54, 55, 99, 184, 185, 186, 188, 192, 193, 105, 106]. Book [107, 108, 109, 110, 116, 128, 130, 131, 135, 136, 137, 138, 142, 143, 144, 161, 162, 165, 166, 169, 172, 179, 181, 209, 210, 212, 213, 215, 218, 219, 225, 228, 229, 230, 232, 235, 238, 239, 240, 254, 255, 256, 257, 267, 270, 307, 313, 314, 318, 342, 343, 344, 345, 348, 356, 357, 363, 388, 389, 390, 391, 396, 397, 401, 404, 406, 407, 416, 417, 418, 432, 448, 449, 450, 453, 454, 455, 458, 459, 462, 468, 488, 489, 491, 493, 495, 496, 517, 520, 528, 538, 559, 560, 561, 562, 563]. Book [565, 566, 567, 568, 573, 585, 586, 620, 621, 623, 626, 627, 646, 649, 650, 660, 661, 662, 665, 667, 669, 671, 672, 674, 676, 677, 678, 679, 699, 700, 701, 702, 708, 710, 711, 730, 731, 740, 754, 755, 757, 758, 759, 772, 773, 777, 778, 793, 795, 797, 799, 800, 811, 812, 814, 815, 816, 817, 834, 836, 837, 841, 843, 844, 848, 850, 851, 191, 319, 400, 427, 429, 430, 539, 35, 36, 52, 57, 58, 134, 706, 174, 211, 233, 173, 668, 835, 852, 776, 533, 838, 273, 274, 320, 451]. Book [465, 564, 569, 666, 733, 736, 452, 464, 516, 590, 604, 756, 588, 592, 647, 652, 682, 330, 364, 365, 628, 712, 651, 792]. Books [356, 699, 767]. Bosmans [787]. Boston [347]. Botany [52]. Boubier [589]. Bourgin [405]. Boussinesq [327]. Boutaric [703]. Boutroux [132, 131]. Bouvier [302]. Bouyges [732]. Br [236]. Brahe [255]. Brahmo [74]. Brasch [773]. Bravais [727]. Breasted [307]. Brehier [321, 225]. Brett [187]. Brewster [468, 672, 817]. Brief [438, 501, 608, 688, 745, 781, 825, 546]. Briefe [390]. Britain [258, 800, 496]. britanniques [491]. British [47, 57]. Browne [171]. Bruhl [818, 317, 681]. Bruner [838]. Brunn [647]. Brunner [309]. Bruno [526]. Brunschvicg [326, 518, 713]. Bryk [209]. Bücher [837]. Bucknall [815]. Buddha [12]. Buddhism [263, 221]. Buddhist [646]. Bude [387]. Budge [109]. Buffon [161]. Builder [479]. Building [505]. Burke [757]. Bury [185]. Busco [541]. Butler [87]. Byzantins [168].

C [208, 420, 80, 392, 166, 212, 458, 660, 661, 755, 799, 400, 173, 452]. C. [544, 430]. cafe [583]. Cajal [583, 583]. Cajori [258, 38, 37, 798, 847, 814, 841]. Called [30, 831]. Caloric [199]. Campanella [107]. Campbell [585]. Canano [661]. Cannizzaro [649]. caractères [21]. Career [841]. Carl [178, 605, 677, 731, 837, 617]. Carpi [661]. Carra [265, 77]. carte [516, 604]. Carter [566]. Cartes [104]. Cartésien [148, 4]. Carus [11, 12]. Case [555]. Caspar [647, 839]. Cassirer [400]. Catalogue [491, 347, 356, 491, 711, 799, 800]. Cauchy [769]. causalisme [761]. causalité [326]. Cause [141]. Causes [121]. cavaliers [560]. ceintures [398]. Celebrate [435, 633]. Celestial [229]. Cell [534]. cellule [21]. celui [433]. centenaire [773]. Centenario [101, 649]. Centenary [435, 633, 717]. Central [452]. Centuries [348, 808, 644, 658, 728, 790, 809, 726]. Century [754, 800, 158, 68, 749, 695, 723, 750, 198, 31, 804, 805, 551, 785, 47, 105, 700, 793]. Cermenati [101]. Chaine [680]. Chambardel [35]. Chambers [116]. Champions [739]. Chandra [10]. Characteristics [706]. charbon [301]. Charity [366, 245]. Charlas [583]. Charles [698, 300, 540, 523, 743, 263, 189, 280, 742, 407, 416, 449, 562, 620, 700, 516, 604, 444, 204, 78]. Chas [46]. Chasse [740].

Chauveau [279]. Cheikh [670]. chemical [512]. Chemico [137]. Chemico-Mathematica [137]. Chemie [177, 81, 579, 316, 178, 228]. Chemische [164, 512]. chemischen [349]. Chemistries [679]. Chemistry [696, 411, 403, 82, 785, 579, 492, 10, 754]. Chequered [841]. Chesters [204]. chez [461, 283, 328, 428, 578, 818, 104, 168, 740]. Chiefly [146, 285, 408, 267]. Chiesa [533]. Child [208, 39]. chimica [269]. chimie [849, 127, 227]. chimique [73]. chimiques [344, 315]. China [305, 566, 739, 128]. Chine [218]. Chinese [306, 813, 567, 706, 459]. choses [563]. Ch’ou [752, 511]. Ch’ou-Jen [752, 511]. Choulant [175]. chrétien [225]. Christentums [467, 467]. Christiaan [257]. Christian [135]. christlich [261]. christlich-palastinensischen [261]. Christophe [516, 604]. Christopher [389, 516, 604]. Chronik [652]. Chronique [5]. Chronologically [356]. Chronometer [363]. Chuan [752, 511]. Chung [306]. chute [218]. chymeías [75]. Chymist [509]. Ciarcia [533]. cinematographe [622]. Cipher [758]. Circulation [280]. Ciruelo [722]. Civilisation [173]. civilisations [108]. Civilised [239]. civilisés [818]. Civilization [147, 368, 598, 120, 63, 22, 194, 88, 145, 241, 332, 284, 371, 472, 498, 593, 606, 630, 684, 779, 714, 762, 819, 853, 361, 850]. civique [144]. Claparede [428]. Clara [432]. Clark [627, 837]. Classical [165]. Classification [290, 142]. Claudius [639, 639]. Clements [179]. Clifford [174]. Clocks [721]. Clod [796]. Clod-Hansen [796]. Close [158]. Clues [586]. Co. [544]. Coast [841]. Coates [646]. Cohn [736]. Cohn-Vossen [736]. Coissac [622]. Colin [46]. Collections [347]. Collier [305]. Colomb [516, 604]. color [615]. Columbian [516]. Columbus [516, 516, 604]. Combes [849]. combustion [641]. Comedy [701]. Commemorating [717]. Comment [23, 428]. Commercial [507]. Common [45]. comparative [680]. comparé [433]. comparée [355]. Compendio [207]. Compendium [719]. complement [293]. completes [201, 257]. composition [641]. compost [563]. comput [772]. Conception [567, 21]. Conceptions [258]. concepts [674]. concetto [319]. Conciliation [327]. Confessions [456]. conflict [751]. conflit [751]. congelatione [755]. conglutinatione [755]. Coniques [694]. connaissance [29, 541]. Conrad [309]. conscience [713]. Conservation [829]. considered [512]. Constantyn [662]. constructions [104]. contemporains [283, 328]. Contemporary [146, 285, 408]. Contents [149, 438, 501, 608, 653, 688, 745, 781, 825, 89]. Continuous [642]. Contributions [173]. Coomaraswamy [347]. Copan [48]. Copernikanische [663]. Copte [668]. Coral [777]. Cordier [218]. Cordus [443, 478]. Corner [669]. corporis [661]. corpusculaire [751]. Correlation [727]. Correspondance [791, 5, 212]. Correspondence [597, 697, 415, 513, 603, 619, 645, 810, 753, 659, 159, 253, 729, 32, 70, 200, 339, 98, 297, 386, 581, 811, 447, 558, 833, 487, 300, 662]. cosmogonies [541]. cosmografico [533]. Cosmological [567]. Couchoud [515]. couleurs [615]. Council [718]. Courant [666, 736]. courants [683]. Courtesy [544]. Cowdry [534]. Creation [642]. Creation. [672]. Creative [808]. Cresson [322, 683]. Cricket [739]. Cristaux [84]. Critical [147, 149, 368, 598, 411, 63, 22, 194, 88, 145, 89, 241, 332, 284, 371, 472, 498, 593, 606, 630, 684, 779, 714, 762, 819, 853]. Criticism [789]. Critique [398, 150, 151]. Crookes [417]. Crusades [488]. cum [835]. Cumont [799]. Cushing [661, 565]. Cuvier [735]. Cylinder [831]. Cyril [843]. Cyrurgia [549]. Cytology [534]. Czwalina [274].

D [760, 224, 356, 755, 616]. D. [224, 458]. d’Abyssinie [667]. Dacque [469, 712]. d’Albe [417]. dall [138]. dalle [272]. Dalmatia [768, 786, 807]. Dalmatie [786, 768, 807]. Dani [255]. Daniel [31]. Danish [94]. Dannemann [136, 235, 202, 282, 359]. dans [139, 515, 190, 713, 167, 108, 35, 29, 668]. d’apres [421, 198, 641]. d’archéologie [460]. d’Archimede [201]. dargestellt [731]. Darmstaedter [837, 310]. Darstellung [273]. Darwin [444]. Dating [800]. Daudin [735]. d’Auguste [8]. David [401, 455, 517, 538, 772, 215]. Davis [777]. Davison [759]. Days [455, 432]. Dayton [391]. Dean [407]. Death [633, 717]. debut [344]. December [371]. découverte [587, 104]. déduction [494]. definition [643, 643]. Definitions [719]. degrees [101]. d’Einstein [427]. Dejust [849]. del [426, 846, 138]. Delacre [127]. Delatte [799, 812]. Delcourt [491]. d’Élie [214]. della [584, 482, 272, 101, 269, 649, 319, 533]. Delusion [432]. Demombynes [421]. Demonbynes [531]. Demonstratis [250]. denen [837]. Denis [551]. Deodat [259]. Depths [463]. depuis [216, 218]. d’Eratosthene [298, 298]. d’Ernst [302]. Desargues [274]. Descartes [517, 662, 517, 299]. Déscription [350]. descrittiva [272]. Dessoir [605]. detaillée [350]. Determinations [823]. determinisme [311, 327]. deutlichste [837]. deutschen [837]. Deutschlands [164]. Developed [116]. Development [158, 534, 121, 808, 57]. développement [225]. Devoirs [20]. d’Extreme [172]. d’Extreme-Orient [172]. dhahab [450]. d’Harcourt [744, 744]. d’histoire [531, 735, 77]. Dichtung [467]. Dickson [85, 133, 353]. Dictionary [109]. Did [30]. Diels [76]. différence [818]. Different [93]. Dijarric [849]. Dijksterhuis [569]. Dimensionen [485]. dimensions [485]. Dimier [161]. din [442]. d’Indianisme [8]. Dinge [837]. Dingler [429, 191]. Diophantine [133]. Disciples [578, 641]. Disciplines [227]. Discoverers [205]. Discoveries [601]. Discovery [443, 280, 829, 59, 110]. Disertaciones [797]. dissectio [661]. Distinctive [723]. divers [212]. Divine [701]. Divisibility [85]. Dobler [435]. Doctrine [19]. doctrines [344]. Dogme [117, 20]. Dogs [305]. Dolomieu [259]. Dondi [721]. D’Ooge [620]. d’or [542]. Dorothea [491, 800]. double [773]. Doublet [268]. Douglas [345]. down [158]. drei [485, 594, 441]. Dresden [177]. Drews [467]. Dreyer [255]. Driesch [80]. Drogen [524]. Drogenhandel [524]. Dubreuil [35, 9]. Dubreuil-Chambardel [35]. d’un [352]. d’une [21, 214, 227, 329]. durch [135, 837, 548]. Durée [427]. during [348]. Dussaud [108]. Dwight [254, 388]. dynastie [218].

E. [87, 496]. Earlier [669]. Earliest [656, 240, 562]. Early [208, 492, 664, 798, 57, 637, 485, 698, 78, 840, 404, 568, 452]. Eastman [407]. Ebene [274]. Eckert [590]. écoliers [428]. Edgar [469, 679, 712, 771]. edifices [315]. Edited [773]. Edition [509, 702]. Editions [698]. Edmond [237, 55]. Edmund [534, 570, 360, 54, 228, 830]. Edouard [428]. Eduard [316, 129]. Eduardo [426]. Education [657, 144]. Edv [81]. Edward [171, 19, 229, 239, 661, 711, 816, 516, 828]. Edwin [165, 307, 468, 672, 817]. Eecke [201]. eens [126]. eeuw [102]. Egee [108]. Egleston [620]. Egypte [668]. Egyptian [527, 109, 788]. Egyptians [613]. Égyptienne [775, 460]. égyptiens [788]. Ehrlich [522]. Eight [653]. Eighteenth [593, 599, 808]. Eighth [147, 63]. Einblattdrucken [238]. Eindringen [736]. eingefügt [837]. Einheit [384]. einiger [548]. Einstein [400, 430]. Eintritt [529]. eìs [75]. Eisenstein [760]. Eleanour [313]. Elected [612, 692]. Electrical [356]. Électricité [279]. Electricity [356, 346]. Electromagnetism [725]. élémentaires [322]. Elementar [273]. Elementar-Mathematik [273]. Elementary [695, 750, 401, 538]. Elementen [126]. Elements [618, 460, 698, 699]. Elephants [535]. Eleven [747]. Eleventh [194]. Eliot [263]. Ellen [14, 759]. Ellery [463]. Elliot [535, 462]. Ellipse [293]. Elmer [571]. Elsbach [430]. emboîtement [21]. Emil [663, 837, 303]. Emile [321, 190, 494, 675, 761, 342, 773, 466, 172, 510]. Emm [318]. Emmanuel [464]. Empire [668, 775]. Empirical [383]. Encyclopedia [306]. Encyclopédie [381]. Endicott [432]. Endo [119]. Energy [829]. Eneström [553]. Engineers [356]. England [664, 852]. English [78, 313, 528, 36, 838, 31, 637]. Enriques [319]. Enseignement [157]. Entomologiques [457]. entre [751]. Entstehung [467, 54]. entwickelt [234]. Entwicklung [316, 349, 359, 136, 235, 666, 736]. Entwicklungsgeschichte [402]. Entwicklungslinien [295]. Entwurf [274]. Epicurus [843]. èpisthmôn [75]. Epoque [668, 421, 788]. Equality [251, 413]. Equation [30]. era [788, 348]. Erganzung [293]. Ergebnisse [274]. Eric [418]. Erich [341, 837, 564]. Erik [815]. erkenntnispsychologisch [234]. Erkenntnistheorie [430]. Ernest [682]. Ernout [99]. Ernst [114, 310, 232, 837, 400, 234]. Errata [23, 24, 147, 148, 149, 150, 151, 152, 366, 367, 368, 594, 595, 596, 597, 598, 594]. Erscheinen [570]. erste [274]. ersten [660]. Erster [274]. Escorial [846]. España [426, 222]. espèce [21]. esprit [283, 419, 734]. Esquisse [329]. Essai [29, 667]. Essay [29, 694]. Essays [730, 40]. essence [406]. Etats [144]. Etats-Unis [144]. etc [392]. éternel [678]. Ether [42, 43, 44, 443]. Ethnographie [399]. Ethnologists [535]. Etienne [523]. étrangers [218]. Étude [531, 298]. Études [735, 849]. Euclid [698, 699, 637]. Eugen [353, 183, 238]. Eugene [85, 401, 407, 517, 538, 772]. Eugenio [130, 423]. Europe [484, 176, 71]. European [586, 808]. Evaluation [773]. Evangelista [256]. Evangelium [467]. Evans [621]. Ever [375]. evo [138]. Evolution [743, 231, 462, 198, 419, 236, 236, 198]. Evolutionary [672]. èvolutions [236]. exactes [167, 168, 397]. exakten [274]. Examination [411]. Exhibition [717]. existence [327]. Experimental [348]. experimentale [50]. explication [190]. Explorer [215]. Exponenten [486, 414]. exponents [414]. Exposè [236]. Expression [105, 556]. External [61]. Extractione [478]. eye [834]. Eyre [815].

F [276, 399, 37, 38, 305, 102, 124, 126, 392, 34, 110, 143, 345, 417, 754, 773, 778, 799, 57, 173, 274]. F. [128]. Faber [354]. Fabré [457]. Fachworter [273]. Facsimile [702]. Facsimiles [498]. Factorial [69]. Faerber [316]. Fagnano [293, 293]. Fahrenheit [95]. Fahs [771, 679]. Fairfield [521]. Faith [2]. Fakhr [442]. Fakhr-al-din [442]. Falasifat [732]. Fark [116]. Farmer [586]. farming [294]. Fasc. [835]. Fasciculo [561]. Fasciculus [416]. Favaro [482]. Features [723, 325]. February [606]. Federigo [319]. Feelings [576]. Feldhaus [651, 124]. Felix [402, 117, 666, 736]. Ferchault [665]. Ferdinand [841]. Fermat [30]. Fernand [536, 774]. Ferrand [845]. Ferriere [431]. Ferton [523]. Fester [349]. fi [450]. Fichte [210]. Field [61]. Fielding [230, 851]. Fifteenth [372, 371, 805]. Fifth [790, 819, 820, 802]. Fifty [346]. fin [344]. Fine [347]. Finger=Reckoning [555]. Finkel [730]. Firak [116]. First [149, 653, 509, 89, 348, 489, 644, 658, 684, 702, 728, 841, 790, 809, 379, 507, 685, 751]. Fischer [303]. fish [294]. Fishes [407]. Fishing [240]. flammae [324]. Fleming [346]. Fleury [356]. Flinders [62]. Flora [395, 532]. florecieron [222]. Florentine [614]. Florian [798, 847, 814, 841, 258]. Fluctuations [808]. Fluxions [258]. Flying [33]. Fobes [34]. followers [578, 641]. Fonahn [267]. foncière [818]. fonctions [818]. fondation [216]. Force [141]. Forel [433]. Forerunners [759]. Forke [567]. forme [226]. Forms [353]. Forscher [282]. Forschung [590]. Foscolo [756]. Foundation [477, 725]. Founders [351]. fourmis [433]. Fournier [417]. Fourteenth [333, 332]. Fourth [803, 728, 779, 780]. Fr [141]. fractional [414]. Fragepunkte [548]. francais [294, 387]. francaise [683]. francaises [350]. France [301, 344, 35, 157, 342, 493, 464]. Francis [160, 361, 520, 566, 759, 544]. Franciscan [36]. Francois [523, 668, 724]. Francon [276]. Frank [620, 626, 341, 175]. Franz [794, 79, 124, 677, 651, 182]. Frazer [542, 542]. Frederic [392, 382, 481]. Frederick [156]. French [294, 768, 786, 807, 749, 28, 198, 578, 641, 788, 615, 751, 576, 547, 29, 805]. Freshfield [345]. Freud [281]. Frick [206]. Friedrich [202, 282, 359, 235, 837, 136]. Friends [546]. Fries [391]. Fritz [186]. Front [1, 26, 65, 92, 154, 196, 244, 289, 335, 336, 370, 374, 410, 439, 440, 475, 476, 502, 503, 545, 575, 610, 611, 634, 655, 690, 691, 716, 746, 764, 783, 801, 821, 826]. Frontispiece [435, 544, 633]. Fruhgeschichte [485]. fruhmittelalterlichen [396]. Fuhrer [135]. Fulton [304]. Function [400]. Fundamental [82]. funfzehnten [238]. Furse [544]. Further [811]. fysikôn [75].

G [535, 457, 622, 176, 227, 205, 45, 171, 9, 13, 221, 456, 530, 124, 255, 462, 36, 835, 274, 516]. G. [461, 460, 103, 364]. Gabriel [845, 493]. Gabrieli [308].

Galilei [733, 235, 663]. Galton [544, 520, 520]. Garcia [426]. Garrison [230, 851]. Gaston [299, 67]. Gathorne [205]. Gathorne-Hardy [205]. Gaudefroy [421, 531]. Gaudefroy-Demombynes [421]. Gaudefroy-Demonbynes [531]. Gauranga [111]. gaz [406]. geb [830]. Geber [310]. Gebrauch [638]. gebrochenen [414, 486]. gedruckten [660]. gedurende [102]. gegen [467]. Gegenwart [202, 81, 628]. Gegonwart [129]. Gehin [294]. gehören [837]. gehörigen [837]. Geikie [453]. geistigen [529]. General [401, 457, 534]. générale [648, 398, 218]. Generalizations [383]. généraux [323]. genese [84]. Genie [518]. Geodasie [79]. geograficheskago [776]. Geographical [488, 776, 516]. Geographie [248, 298, 298]. Geography [580, 248, 464]. Geologie [536]. Géologique [24]. geologischen [83]. geologue [352]. Geology [489]. Geometria [272]. geometric [548]. Geometricae [39]. Geometrical [722]. Geometrie [295]. géometrique [350]. geometrischen [548, 274]. Geometry [822, 695, 723, 750, 517]. Georg [837, 592]. George [664, 542, 709, 60, 187, 463, 852, 838, 789]. Georges [629, 323, 587]. Georgius [837]. Gerasa [620]. Gerhardt [208]. Gerlach [320]. German [594, 512, 293, 639, 748, 385, 643, 830, 247, 248, 441, 548, 506, 638, 414, 485, 613, 579]. Germanen [525]. Gertrude [759]. Geschichte [312, 79, 275, 360, 228, 232, 396, 414, 486, 579, 275, 364, 673, 178, 114, 115, 261, 570, 572, 81, 181, 134, 273, 451, 733]. geschichtlich [512]. geschichtliche [316]. Geschichtliches [385]. Geschichtlichkeit [467]. Geschiedenis [569, 102]. Geschütz [844]. Gesetze [402]. Gesner [390]. gest [830]. gestellt [837]. Geuthner [422]. Ghazali [206]. Gibson [52]. Gift [356]. Gifte [184]. Gilbert [192]. Gillain [775]. Gino [272, 256, 676]. giorni [272, 138]. Giovanni [584, 721]. Girolamo [661]. Giuseppe [308, 256]. globe [433]. Globes [229]. Gloire [352]. Gnostizismus [467]. Goblot [237, 55]. Goguel [515]. Goldziher [117]. Gospel [12]. Gould [363]. Graebe [178]. Graeco [506]. Graesse [124]. Grandmaison [515]. grant [563]. grape [385]. Graphic [175]. Grasset [20]. Gravitation [42, 43, 44]. Great [258, 521, 496, 800, 507]. Grèce [738]. grecque [419]. grecs [491, 461, 799]. Greece [262]. Greek [110, 219, 528, 843, 506, 174]. Greek-English [528]. Greenland [100]. Greenstreet [773]. Gregor [490]. griechischen [166]. Griswold [48]. grosse [470]. grosser [282]. Grossindustrie [349]. Growth [383, 588]. Grubb [758]. Grundbegriffe [736]. Grundlagen [525, 429, 274, 590]. Grundriss [842]. Gudger [407]. Gudmund [71]. Guepes [523]. Guglielmo [324]. Guillaume [387]. Guilleminot [51]. Guimet [531]. Gulat [605]. Gulat-Wellenburg [605]. Gunther [840, 404, 568, 811]. Gustav [394, 553, 349].

H [457, 849, 123, 203, 420, 100, 266, 34, 212, 230, 396, 754, 851, 260]. H. [51, 110, 546]. Haeckel [163]. Hagemann [255]. Hakim [377]. Hakluyt [838]. Hale [463]. Hall [456]. Haller [390]. Hallers [390]. Hammer [217]. Hammer-Jensen [217]. Handbuch [275, 394]. Handschrift [647].

Hanotaux [493]. Hans [605, 572, 80]. Hansen [796]. Haramain [624]. Hardy [205]. Harlow [848]. Harold [701]. Harper [646]. Harpestræng [94]. Harriot [766]. Harrison [382]. Harry [571]. Harvey [565, 661, 850, 52]. Harvey-Gibson [52]. Hasan [527]. Haskins [449, 700]. Hassler [841]. Haughton [112]. Hauptwerk [570]. Havelock [646]. Hearnshaw [173]. Heath [699, 219]. Heawood [516]. Heiberg [397, 770, 397, 699, 710]. Heights [823]. Heilpflanzen [312]. Heinrich [163, 312, 206, 296, 731, 134, 137]. Helbronner [350]. Helen [848]. Helene [84, 344, 674]. Hellenen [525]. Hellenism [111]. Hellenismus [530]. Hellmann [103]. Henn [407]. Henri [419, 787, 218, 491, 424, 427]. Henrik [94]. Henry [521, 735, 340, 105, 254, 307, 345, 388, 390, 396, 528, 562, 586, 661, 837]. Herbal [742, 585]. Herbals [313]. Herbert [314, 837, 66]. Herdman [351]. Heredite [50, 323, 21]. Heredity [534]. Hermann [278, 76]. Hermetica [559]. Heron [816]. Heron-Allen [816]. Hice [314]. Hidâyat [754]. Hieroglyphic [109]. Higher [353, 847]. Hill [56]. Hindu [6, 7, 10]. Hinduism [263]. Hippocrates [340]. Hiram [87]. Histoire [150, 151, 329, 515, 667, 446, 214, 515, 622, 127, 268, 321, 587, 680, 72, 218, 493]. Historia [375, 426]. Historical [263]. historically [512]. historiens [734]. historisch [234]. Historische [102]. History [147, 368, 598, 557, 258, 69, 199, 540, 823, 571, 375, 709, 60, 180, 187, 221, 93, 10, 189, 80, 325, 847, 63, 27, 22, 192, 193, 116, 165, 194, 96, 88, 145, 155, 219, 230, 241, 270, 314, 332, 284, 348, 356, 401, 371, 449, 496, 472, 498, 593, 606, 665, 672, 644, 658, 630, 684, 728, 779, 714, 762, 814, 815, 851, 790, 809, 802, 819, 853, 52, 57, 381, 832, 776, 580, 482, 414, 485, 579, 612, 692, 803, 477, 635, 718, 385, 504, 224, 270, 36, 85, 133, 353, 119, 175, 74, 401]. History [538, 682]. Hjelt [81]. Hobson [41]. Hochschule [177]. Hodgson [496]. Hohenheim [343]. Hoin [646]. Holder [465]. Holland [213]. Hollander [183, 238]. Holmyard [450, 755]. Homer [709, 449, 700]. Honen [646]. Hoover [837, 837]. Hoppe [830]. Horace [345]. Horologe [831]. Houa [169]. Houssay [481, 141]. Hovelaque [172]. Howarth [848]. Hubert [405]. Hubotter [182]. Hudhail [560]. Hugo [490, 836, 191, 429]. humain [734]. humaine [20, 326]. humani [661]. Humanism [338]. humanisme [387]. Humanist [2]. humanite [23]. Hunain [602, 834]. Hunayn [376]. Hundred [489]. Husain [754]. Hüttenwesen [837]. Huygens [662, 662, 257]. Hypothese [191]. Hypotheses [42, 43, 44].

Ibn [376, 602, 560, 594, 412, 203, 116, 450, 730, 731, 834, 441, 638]. ibn-Tahir [116]. Ibrahim [624]. iconographique [225]. Idea [185]. idéalisme [20]. Ideas [37]. idée [214]. idées [142]. Idem [221]. identité [818, 675]. ìdíos [75]. Ignatius [596, 388]. Ignaz [117]. ihr [736, 330]. ihre [312, 330]. ihrem [359, 136, 235]. ihren [525]. ihrer [402, 202, 359, 136, 235]. II [133, 395, 448, 538, 736, 156, 786]. III [353, 15, 532, 167, 491, 733, 807]. Illustration [175]. Illustrations [557, 644, 658, 728, 790, 809]. illustrés [788]. ’ilm [450]. Iltis [490]. im [524, 275, 309, 181, 844, 666, 736]. Imam [448]. Immanuel [395, 532, 435]. Immigrant [454]. Impartial [596, 388]. important [293]. Impressions [521]. Incas [744]. Indern [525]. Index [333, 286, 372, 473, 436, 687, 824, 18, 820, 854, 499, 599, 607, 631, 685, 715, 763, 780, 595]. India [17, 111]. Indian [264, 347, 789, 29, 627, 220]. Indianism [28]. indienne [29]. indische [113]. indischen [114, 115]. indologique [11]. Industrial [625]. Industrie [164]. industrielle [144]. inedite [649, 806]. inférieures [818]. Influence [37, 221, 586]. Information [152, 90, 242, 288, 782, 856]. Ingeborg [217]. inizio [138]. Ink [626]. inorganique [236]. Inscriptions [48]. Insect [739]. Insect-Musicians [739]. Insects [433]. insigne [560]. Institute [356]. Instrumente [836, 837]. Integralrechnung [274]. intellectuel [23, 547]. intellectuelle [236, 547]. intellectuelles [322]. Internal [354]. Interpretation [789]. Introduction [7, 325, 188, 96, 230, 270, 851, 804, 446, 281, 709, 845, 620]. Invariantentheorie [736]. inventeurs [294]. Invention [41, 566, 213]. Inventor [454]. inventors [294]. Ioannes [661]. Ioannis [255]. Iran [529]. Iranian [221]. Iranica [53]. Iraniern [525]. iranischen [530]. ’Iraq [754]. ’Iraqi [450, 577]. Ireland [800]. Iridaceae [395]. Iridaceae-Papilionaceae [395]. Iron [625]. irrational [643]. irrationnels [643]. irréductible [818]. ’Isa [638, 638]. Isaac [717, 39, 773, 633]. Ishâq [834]. Ishaq [376, 602]. Ishizuka [646]. Isidore [738]. Isis [653, 514, 546]. Islam [223, 394, 116, 701]. Islam-Literatur [394]. islamischen [420]. Islamstudien. [420]. Islands [266]. Israel [19]. Istoria [776]. ístorían [75]. Italian [482]. italiani [482, 138]. Italien [579]. Italikern [525]. Italy [579]. IV [101, 168, 595, 436, 840]. Iwan [592]. IX [835, 688, 687, 636].

J [616, 208, 760, 457, 20, 39, 56, 422, 17, 680, 849, 103, 124, 125, 275, 209, 215, 397, 450, 491, 496, 710, 755, 773, 799, 836, 58, 173, 274, 320, 569]. J. [52]. Jâhiz [730]. Jackson [391]. Jacob [113]. Jacopo [721]. Jacques [591]. Jaeger [102, 126]. Jaennicke [124]. Ja’far [448]. Jahiz [730]. Jahr [103]. Jahre [182]. Jahrhundert [736, 666]. Jahrhunderts [238]. James [300, 542, 276, 112, 270, 307, 850, 58, 838, 300]. Jankelevitch [281]. January [612]. Japan [305, 120, 121, 123, 708, 120]. Japanese [119, 122, 123]. Japon [172]. Jean [315, 387, 774]. Jen [306, 752, 511]. Jenner [828]. Jensen [217, 572]. Jequier [460]. Jesu [467]. Jesuit [452]. Jesus [515, 515]. Jewish [788]. Joan [621]. Joannes [416, 561]. Joaquim [72]. Joh [79]. Johan [770]. Johann [490, 209, 274]. Johannes [748, 839, 390, 652, 273]. John [519, 492, 346, 185, 193, 128, 165, 488, 573, 711, 811, 748, 767, 41]. Johnson [813]. Johnston [45]. Jones [528, 340]. Jordan [455]. Jose [222, 207, 846]. Josef [529]. Joseph [327, 799, 362]. Joshua [730]. Jouguet [466]. Jourdain [42, 43, 44, 367, 33, 40, 252]. journal [629]. jours [225]. Jouveau [9]. Jouveau-Dubreuil [9]. Jr. [767]. Jud [441, 594, 441]. Juden [395, 532]. juifs [788]. Jules [86]. Julián [797]. Julius [118, 83, 448, 650, 731]. July [612, 692]. June [22]. Jura [318]. jusqu’a [668, 218]. jusqu’en [151].

Jussieu [735]. just [613].

Kahir [116]. kaì [75]. Kalendrier [563]. Kamal [527]. Kammerer [667]. Kampf [663]. Kanetsune [123]. Kant [435, 430, 564]. Karikatur [183]. Karl [275, 181, 520, 562, 660, 233]. Karpinski [540, 620]. Karplnski [204]. Kartenwissenschaft [590]. Kartographie [590]. Katalog [365]. Kate [116]. Kaye [789, 795]. Kegels [274]. Keith [18, 220]. Kelten [525]. Kennis [102]. Kenntnis [166]. Kensington [633]. Kent [758]. Kepler [839, 209]. Keplers [274]. Ketham [416, 561]. Keyserling [278]. Khayyam [709]. Khowarizmi [204]. Khwarizmi [203]. Kiai [169]. Kirtland [488]. Kishori [7]. Kitab [450]. Kittredge [664, 852]. klarste [837]. Klassiker [274]. Klebs [660]. Klein [666, 736]. Klinckowstroem [605]. Knauth [837]. Knowledge [366, 61, 245, 249, 29]. known [788]. Knud [354]. Ko [705]. Ko-an [705]. Königsberg [435]. Konrad [671]. Koren [193]. Kowalewski [274]. Kraem [776]. Krankenpflege [309]. Kritik [320]. Kubitschek [842]. Kugler [794]. Kulturhistorische [166]. Kulturptlanze [570]. Kumar [6]. Kunstgewerbliche [124]. Kuo [306]. Kurd [673]. Kuriositaten [124]. Kurzer [275].

L [133, 21, 515, 849, 275, 184, 255, 357, 397, 517, 699, 710, 776, 577, 765, 731]. l-Qasim [577]. l-Raihân [731]. L. [50, 35]. Labour [192]. Labours [520]. L’Académie [216]. Lacroix [259]. Lacy [224]. Lady [542]. Lafayette [403]. Lagercrantz [799, 491]. Lal [7]. Lamarck [735]. L’ame [681]. Landen [309]. Lands [179, 110]. l’antiquité [167]. Lapidaries [621]. lapidum [755]. Laplace [773, 260]. l’architecture [460]. L’Arithmétique [775]. L’art [225, 342]. Lasswitz [673]. l’astronomie [268]. lateinischen [837]. Latham [517]. Latin [204, 267, 800, 832]. Latitudes [249]. L’Atome [226]. Laufer [739, 53]. Laurent [774]. L’avenir [276]. Law [829, 264]. Lawrence [389, 239]. Lebegue [491]. Leben [490, 247, 303]. Lebens [529]. Lebensgebieten [330]. Lebenslinien [704]. Lebenslinien. [737]. Lecat [425, 425]. Lecat-Pierlot [425]. Leclerc [140]. Lectiones [39]. Lectures [123, 466, 47]. l’education [139]. Legacy [262]. Legendary [266]. légende [738]. Legrand [21]. Legros [457]. Lehmann [628]. Lehre [663]. Lehren [731]. Leibniz [479, 208, 640, 208]. Lemery [480]. Lenoir [734]. Leon [326, 518, 713, 662, 419, 107, 210]. Leonard [85, 353, 815, 549]. Leonardo [101, 451, 733]. Leopold [16]. Lepsius [164]. Leroy [818]. Lespieau [73]. Letter [828]. lettere [649]. Letters [520, 767]. lettre [806]. Letts [82]. leur [563]. leurs [142]. L’Évolution [236, 589]. Levy [681, 317, 738, 818]. Levy-Bruhl [317]. Lewin [184]. Lewis [188]. Lexa [668]. Lexicon [528]. L’expérience [326]. L’Heredite [49]. l’homme [433]. l’Humanité [419]. L’Hygiene [461]. Li [459]. Libby [325]. liberté [327]. Library [356, 711]. Libri [34]. Liddell [528]. L’Ideal [131]. Liége [805]. Life [300, 519, 40, 46, 433, 392, 345, 417, 453, 458, 520, 565, 551, 247, 403, 456, 160]. Light [602, 508, 751]. Like [808]. Limits [258]. Lindau [647]. l’Inde [9]. l’Indianisme [28]. Linnaeus [391]. Linné [735]. Lippmann [570, 360, 54, 228]. l’Islam [117, 265, 422]. List [612, 692, 803, 477]. Litauern [525]. Literary [221]. Literatur [261, 733, 115, 394, 135, 451]. Literature [562, 57, 221]. Lith [594, 441]. Little [835, 36]. Littré [510]. Livingstone [215, 262]. livre [629]. Locy [588]. L’oeuvre [260]. Loewenstein [592]. Logarithms [41]. Logic [220]. logica [319]. logical [643]. logique [643, 55]. Loi [117, 431]. lois [236]. Long [453]. Longitudes [249]. Lore [488]. Loria [256, 272, 676]. l’ornithologie [589]. Lorraine [804]. Lothrop [223]. Lou [837]. Louis [540, 537, 161, 225, 560, 620, 740, 393, 398, 204]. Louis-Charles [204]. Low [395, 532]. Loyola [388, 596]. Lucien [317, 681, 818]. Lucrece [99]. Ludvig [770]. Ludwig [175]. Luigi [756]. Lulu [470]. lumière [751]. Lung [459]. Lung-Mien [459]. L’unité [140]. Luther [620, 229]. Lyman [664, 852]. Lynn [348].

M [208, 49, 127, 140, 39, 56, 115, 761, 62, 124, 425, 417, 458, 560, 754, 274]. M. [205, 102, 126, 651]. maîtresses [142]. Macdonell [18]. Macfarlane [47]. Mach [234, 302]. Mackenzie [276]. Maclaurin [46]. Madjriti [203]. Maeterlinck [470]. Mager [424]. Magic [78, 348, 778, 122]. Magie [668, 362]. Magister [772]. Magnetism [356]. Magnus [391]. Mainly [847]. Maire [648, 774]. Majriti [377]. Majus [757]. Male [342]. Mamelouks [421]. Man [627, 455, 462]. mandchoue [796, 218]. Mandeville [755]. Manning [392]. Mansur [116]. Manuale [308]. manuel [772, 460]. Manuscript [549, 795]. Manuscripts [208, 800, 832]. Manuscrits [799, 491]. Map [516, 604]. Mappae [671]. Maps [71]. Marâgha [836]. Marc [495]. Marcel [283, 301]. Marcellin [703]. March [633, 241, 498, 853]. Marcia [517]. Marco [756]. Marcus [254]. Marg [744]. Margerie [318]. Marie [458, 491, 400]. Marine [363]. marines [104]. Mario [514, 101]. Markham [179]. Markus [467]. Markus-Evangelium [467]. Marquardt [522]. Marshall [110]. Martha [522, 793]. Martin [620]. Martonne [464]. Marvin [361]. Marzell [312]. Mas‘ûdî [731]. Maschinen [837]. Maslama [203, 377]. Mason [180]. Massaoth [760]. Masson [355]. Masson-Oursel [355]. Master [479, 792]. Master-Builder [479]. matemáticas [846]. matematici [319]. matemáticos [222]. Materials [146, 285, 408]. Mathematica [741, 137, 539]. Mathematical [208, 251, 383, 479, 246, 798, 93, 847, 418, 773, 814]. mathematician [594, 441]. mathematicians [548]. mathematiciens [131]. Mathematics [789, 121, 401, 538, 37, 119, 122, 847, 219, 401, 538, 706]. Mathematik [640, 273, 666, 736, 134]. Mathematiker [548]. Mathematikers [594, 441]. Mathematique [132]. mathematische [465, 736]. mathematischen [429, 320]. Matiere [51, 578]. Matter [25, 64, 1, 26, 91, 153, 195, 65, 92, 154, 243, 196, 287, 331, 334, 244, 289, 335, 369, 373, 409, 434, 336, 370, 374, 410, 437, 471, 474, 497, 500, 439, 440, 475, 476, 502, 543, 574, 600, 609, 503, 545, 575, 610, 632, 654, 686, 611, 634, 655, 689, 690, 691, 716, 746, 764, 855, 783, 801, 821, 826, 726, 578]. Maurice [589, 515, 421, 531, 732, 849, 470, 425]. Max [605, 839, 166, 834, 211, 590, 682]. Maxson [492]. Maxwell [362]. May [630, 779, 714, 332]. McKenzie [528]. Mècanique [157, 21, 327, 466]. Mécanisme [148, 4]. Mechanica [569]. Medallic [557, 644, 658, 728, 790, 809]. médecine [144, 276]. médecins [788, 35]. Mediaeval [449, 58, 173]. Medical [614]. Medici [412]. Medicina [561, 426]. Medicinae [416]. medicinalibus [835]. Medicine [325, 174, 832, 231, 171, 527, 78, 354, 230, 851, 58, 682]. Medico [719]. Medico-Philosophical [719]. Medieval [378]. medio [138]. Medizin [275, 309, 181, 396, 182, 183, 275, 232]. medizinische [135, 211]. Meeting [718, 717, 802]. Meetings [635]. mehr [485]. meinem [303]. Meissner [526]. Mekke [531]. Members [612, 692, 803, 477]. Membres [398]. mémoire [423]. Mémoires [167, 168, 623, 676, 710, 397]. Memorial [773]. Mendel [490, 403]. Mensch [522]. Menschheit [469]. mentales [818]. Mentalite [317]. mer [108]. Mercier [560, 740]. Merrill [489]. metal [198]. metallica [837]. métallique [198]. Metchnikoff [214, 214]. Meteorologicorum [34]. Method [61, 189, 143]. Methode [465]. Methodik [83]. Methods [158]. Methuen [544]. Metz [761]. Metzger [84, 344, 674]. Meunier [514]. Meyer [275, 181, 177, 459]. Meyer-Steineg [275, 181]. Meyerhof [834]. Meyerson [190, 494, 675, 761]. mi [583]. Michael [75, 454, 156, 720, 827]. Michel [622, 670, 799]. Michigan [246]. Middle [267, 669, 249]. Mieli [170, 269, 138]. Mien [459]. Miguel [701]. Mikami [121, 122, 706]. Milhaud [67, 299]. Milione [756]. Miller [671]. mineralogisch [83]. mineralogisch-geologischen [83]. mines [301]. Ming [306]. Minor [13, 614]. Mir’at [624]. Missionary [215]. Mitteilungen [836]. Mittelalter [673, 309, 844]. Modern [445, 231, 354, 173]. moderne [710]. modernen [402, 430]. modernes [541]. Mohammed [670]. Molecule [73]. monde [433]. mondiale [236]. Mongolia [128]. Montagnier [345]. Montessus [536, 399]. Moodie [357]. morale [327]. morbide [49]. Morgan [40, 591]. Morley [31, 48]. Morphologie [271, 398]. morphologique [311, 398]. Morris [777]. Mors [556]. Morte [101]. Mortimer [175]. Morton [433, 665]. Moseley [618]. Moslem [221, 116]. Most [375]. moteurs [406]. Motion [3]. Mottelay [356]. Mountains [823]. Moustapha [670]. mouvement [176]. Moyen [775, 397]. Mrs [217]. Muh [731]. Muhammad [594, 450, 441]. Muhammed [203]. muktasab [450]. Muller [79]. Musa [203]. Musculorum [661]. Musée [531]. Museum [347]. Music [123, 123]. Musical [586]. Musicians [739]. Musique [744]. Muslim [642, 548]. muslimische [548]. musulmana [308]. mystère [515, 515]. mystique [86]. Myth [816]. mythe [515]. Mythology [188].

N [741]. nach [731, 836]. Names [722, 18]. Napier [41]. Narendra [264]. Nariman [221]. nascita [649]. Nasiruddin [290]. Nath [264, 111]. Nationale [516]. Natur [712]. Natural [375, 665]. naturalis [375]. naturaliste [457]. Naturalists [521]. Nature [305, 627, 816, 125, 99]. naturelles [735]. Naturforscher [564]. Naturgeschichte [202]. naturgeschichtlicher [794]. Naturphilosophie [429]. Naturwisschenschaften [359]. Naturwissenschaften [360, 136, 235]. natuurwetenschappen [102]. nautique [72, 845]. navales [104]. Naville [142]. Nazareth [515]. Nederlanden [102]. Nehlil [560]. nel [649, 319]. Nempyo [707]. Neuburger [682, 364, 211]. Neudeck [364]. Neue [839]. Neugebauer [666]. neusprachlichen [451, 733]. Newbold [758]. Newton [258, 773, 42, 633, 717, 43, 44, 806, 615, 673, 40, 569, 43, 44, 42]. nicht [837]. Nicolas [480]. Nicomachus [620]. Nine [149, 89]. Nineteenth [47, 606, 607, 808]. Ninth [88]. No [246]. Nohl [652]. nombres [643]. Non [672, 818]. non-civilisés [818]. Non-Evolutionary [672]. Norden [129]. Nordenskiöld [815]. Nordvolker [529]. Norman [621]. Normands [104]. Norse [205]. Norsemen [100]. North [798, 539]. Northern [71]. nos [225]. Nosography [354]. nostri [138, 272]. Notable [555]. Notations [383, 479, 847, 814]. Note [95, 413, 96]. Notes [597, 697, 791, 415, 513, 603, 789, 619, 645, 810, 753, 659, 788, 159, 253, 729, 32, 70, 200, 339, 98, 297, 386, 581, 447, 558, 833, 767, 249, 556, 487, 788, 77]. Notices [146, 285, 408]. nous [818]. nouvelle [21, 424, 761]. November [593]. Nr [274]. Numbering [618]. Numbers [85, 133, 353, 643]. Nunn [516]. nur [837, 613]. Nutrition [403]. Nuzhatu [765]. Nuzhatu-l-Qulûb [765]. Nyoden [707].

O [775, 530, 570, 360, 54, 228, 417, 799, 666]. Obed [813]. Obituary [146, 285, 408]. occidentale [713]. Oceanography [351]. Octavo [233]. October [368, 194, 284, 684, 762, 819]. Oeuvre [481, 11, 214]. Oeuvres [201, 8, 648, 257]. Ogata [705, 705]. Ogden [80]. oil [385]. Okkultismus [605]. Ölbereitung [385]. Old [82, 313, 852, 679]. O’Leary [224]. Olga [214]. Olivier [818]. Olschki [451, 733]. Omar [709]. Omnia [255]. ondulatoire [751]. One [489]. Opera [255]. Opere [582, 256]. Optik [234, 274]. Optique [446]. Opus [757]. opusculis [835]. opúsculos [797]. ordine [319]. organique [236]. Organisation [150, 151, 144]. Organischen [178, 81, 228]. organisme [431]. Organization [147, 63, 22, 88, 145]. Orient [172]. orientale [591]. Origin [822, 556]. origines [387, 419, 225]. origini [272]. Ornithological [711]. Ornstein [793]. Orokaiva [778]. Orsted [212]. Orte [837]. Osborn [521, 105]. Osborne [162]. Oscar [135]. Osiris [514]. Osler [565, 231]. Ostwald [704, 737]. Ostwalds [274, 274]. Oswald [271]. Other [759]. Othman [730]. Otsuki [707]. Otto [106, 209, 491, 465]. ouest [35]. Oughtred [38, 38, 37]. Our [413, 348, 61]. Oursel [355]. Outline [314]. Outlines [52]. Oxford [840, 404, 568]. Ozar [760]. oznakomleniya [776].

P [201, 515, 422, 12, 787, 216, 166, 489]. P. [100]. Packard [160]. Padri [533]. Pagels [275]. Pagine [269]. Painting [435, 544, 459]. palastinensischen [261]. Paleopathology [357]. Palestine [738]. Pamphlet [601]. Pamphlets [356]. Panayotatou [461]. Papilionaceae [395]. Papyrus [246, 307, 418]. Paracelsus [343]. Paracelsus. [343]. Parat [849]. Pare [160]. Paris [774]. Parisini [491]. Parks [838]. Parsons [304]. Part [478]. particle [751]. particulières [236]. parure [560]. Pascal [694, 648, 648, 518]. Pasha [624]. passè [576]. Past [708, 576, 573]. Pasteur [393]. Patanjali [112]. Paucot [139]. Paul [422, 167, 168, 350, 355, 356, 397, 621, 623, 676, 710, 837, 11, 522]. Payot [563]. pays [218]. Pearson [520]. peau [849]. Pedaliaceae [532]. Pedaliaceae-Zygophyllaceae [532]. Pediatrics [573]. Peet [418]. Pelerinage [531]. Pellian [30]. Pelseneer [216]. pensatori [319]. pensée [302, 419, 683]. Penseurs [283, 265]. Perez [222, 207, 846]. perils [20]. Period [602]. Periodicals [356]. Persia [754]. persischen [731]. Pest [652]. Peste [768, 786, 807]. Pestschriften [660]. Petrie [62]. pétrole [406]. Petronievics [236]. Petrucci [169]. peuples [172]. Pfannmuller [394]. Phaëton [794]. Pharmacopoeia [702]. Philip [42, 33, 40, 43, 44, 367, 252]. Philologie [114]. Philoponos [748]. Philoponus [748]. Philosophen [278]. Philosopher [31]. Philosophic [116]. Philosophical [719, 567, 657]. Philosophie [150, 151, 578, 710, 20, 321, 713, 761, 355, 678, 29]. philosophique [683]. Philosophy [147, 368, 598, 45, 63, 22, 194, 96, 88, 145, 241, 332, 284, 371, 472, 498, 593, 606, 630, 684, 779, 714, 762, 819, 853, 578, 29, 61]. photographie [587]. physicians [788]. physiciens [328]. Physicists [47]. Physico [567, 315]. physico-chimiques [315]. Physico-Philosophical [567]. Physics [808, 7]. Physik [402, 429, 736, 191]. physikalischen [234]. physiologie [148, 4]. physiologique [849]. physique [328, 326, 678]. Phytopathology [314]. Picard [523, 773]. picturata [661]. Pieper [837]. Pierlot [425]. Pierre [132, 315, 131, 541, 352]. Piraeus [505]. pisciculture [294]. pl [498]. Place [224]. Plague [768, 786, 807]. Plana [727]. Planet [817]. Planetarium [552]. planimetric [594, 441]. planimetrische [594, 441]. Planisphaerium [639, 639]. plasma [21]. plasmas [21]. Plato [341]. Plattard [387]. Playfair [682]. Plea [580]. Plehn [274]. Plinius [202]. Pliny [375]. plus [218]. plusieurs [563]. Plutar [514]. Polity [264]. Polo [756]. Popular [375]. Portrait [435, 633]. portugaise [72]. positif [20, 283]. Positive [6]. Potonniee [587]. pour [694]. Poyer [323]. Practical [785, 613]. Prague [769]. praktisch [613]. Prapancha [15]. Praphulla [10]. pratique [616, 406]. Praxis [766]. Prearistotelici [170]. Preface [337, 504, 636, 747, 784]. prehelleniques [108]. préhistoire [591]. Preisler [135]. Preliminary [96]. premier [774, 751]. preparation [385]. Present [708]. preuves [236]. preux [560]. Primality [85]. primitive [818, 317, 681]. Principes [132]. Principia [741, 539]. principii [319]. Principle [97]. Printed [562, 379]. Printing [566, 626]. Prinzipien [234, 429]. problem [512, 643, 777]. problème [643, 724]. problèmes [50, 323]. Problems [82]. Proceedings [718]. Productivity [808]. Professor [177]. Prognose [103]. progres [431, 713]. Progress [185]. Progressions [722]. Proportion [56]. Proportions [722]. propos [427]. Prostitution [592]. Psellus [799]. psychanalyse [281]. Psychologie [130]. psychologique [323]. Psychologist [456]. Psychology [358, 187]. Ptolemaeus [639]. Ptolemy [639, 71]. publications [150, 151]. publiès [397]. Published [656, 375, 504]. Pupin [454]. Purnananda [14]. Purpose [3]. Puzzling [817]. Pythagore [738, 738]. Pythagoreer [341].

Qânûn [731]. Qasim [577, 450]. Quadratic [353]. Quadripartitum [250, 291]. Quantentheorie [186]. Quarante [8]. Quattuor [34]. que [222, 514]. Quelle [247]. Quellen [579]. quelques [788]. Quest [458]. question [431]. questions [548]. qui [788]. Qulûb [765].

R [544, 73, 139, 13, 849, 787, 203, 262, 392, 417, 754, 759, 52, 666, 736]. R. [160, 179]. rôle [793]. Rabaud [523]. Races [422, 239]. Radcliffe [240]. Raeder [255]. Rahman [560]. Raihân [731]. raison [818]. Raisonnement [130]. Rameau [542]. Ramon [583]. Raoul [744, 311, 849, 818]. Raphael [169]. rapports [142]. Rapson [17]. Rare [601]. Rasis [785]. Rat [849]. Rathgen [844]. Ratsel [470]. Raume [485]. Ray [10, 811]. Raymond [734]. Razi [442]. Razik [670]. Razi’s [247]. Re [642, 837]. Re-Creation [642]. réactions [322]. Real [426]. Realite [675]. really [613]. realm [198]. Reaumur [665]. rebus [835]. Rechenkunst [118]. Recherches [738]. Recorde [508]. Recuerdos [583]. Reden [16]. Reef [777]. Reference [27]. Reflected [459]. Reggio [550]. Reginald [585]. Region [776]. regne [198]. Reiche [186]. Reichspatentamts [365]. Reisetagebuch [278]. Relation [175, 627, 727]. relations [218]. relatives [150, 151]. relativiste [494]. Relativitatstheorie [430]. Relativite [425]. Relativity [400]. religieuse [531]. religieux [342]. Religion [467, 468]. Religionsgeschichte [530]. Religious [13]. Remy [294]. Renaissance [700, 614]. Renard [629]. rendement [23]. Rene [517, 665, 108, 144, 724]. René-Francois [724]. répertoire [457]. Report [635, 717, 802]. retardatione [835]. retour [678]. Revaluation [766]. Review [698, 744, 775, 794, 842, 163, 164, 177, 234, 303, 316, 490, 522, 704, 737, 208, 258, 300, 540, 519, 306, 85, 132, 133, 353, 492, 664, 571, 625, 663, 673, 760, 535, 362, 405, 457, 461, 523, 536, 541, 542, 582, 584, 589, 591, 622, 629, 648, 703, 738, 774, 20, 21, 49, 50, 51, 73, 84, 86, 127, 139, 140, 141, 176, 201, 214, 226, 227, 236, 237, 259, 260, 268, 276, 277, 279, 281, 283, 298, 301, 302, 311, 315, 321, 323, 327, 328, 329, 352, 387, 393, 398, 399, 419, 423, 424, 426, 428, 431, 515, 537, 587]. Review [709, 205, 178, 217, 33, 37, 38, 39, 40, 41, 43, 44, 45, 46, 56, 60, 61, 118, 272, 798, 170, 469, 402, 741, 305, 312, 813, 743, 460, 317, 358, 180, 187, 202, 231, 117, 171, 206, 222, 223, 224, 421, 422, 531, 670, 732, 6, 7, 8, 9, 12, 13, 14, 15, 16, 17, 18, 19, 112, 113, 114, 115, 220, 221, 263, 264, 403, 433, 456, 521, 534, 583, 680, 735, 849, 190, 322, 326, 494, 518, 675, 681, 683, 713, 734, 761, 818, 527, 624, 119, 120]. Review [121, 122, 123, 705, 707, 10, 514, 605, 203, 204, 207, 261, 265, 271, 278, 310, 341, 349, 394, 395, 420, 467, 470, 524, 525, 526, 529, 530, 532, 570, 572, 48, 59, 62, 71, 175, 182, 183, 189, 72, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 87, 100, 101, 102, 103, 104, 111, 124, 125, 126, 129, 160, 167, 168, 216, 262, 266, 269, 275, 280, 282, 299, 304, 308, 309, 324, 325, 340, 346, 347, 350, 351, 354, 355, 359, 360, 361, 392, 425, 463, 466, 739, 742, 796, 839, 840, 845, 846, 847, 34, 47, 53, 54, 55, 99, 184, 185, 186, 188, 192, 193, 105, 106]. Review [107, 108, 109, 110, 116, 128, 130, 131, 135, 136, 137, 138, 142, 143, 144, 161, 162, 165, 166, 169, 172, 179, 181, 209, 210, 212, 213, 215, 218, 219, 225, 228, 229, 230, 232, 235, 238, 239, 240, 254, 255, 256, 257, 267, 270, 307, 313, 314, 318, 342, 343, 344, 345, 348, 356, 357, 363, 388, 389, 390, 391, 396, 397, 401, 404, 406, 407, 416, 417, 418, 432, 448, 449, 450, 453, 454, 455, 458, 459, 462, 468, 488, 489, 491, 493, 495, 496, 517, 520, 528, 538, 559, 560, 561, 562, 563]. Review [565, 566, 567, 568, 573, 585, 586, 620, 621, 623, 626, 627, 646, 649, 650, 660, 661, 662, 665, 667, 669, 671, 672, 674, 676, 677, 678, 679, 699, 700, 701, 702, 708, 710, 711, 730, 731, 740, 754, 755, 757, 758, 759, 772, 773, 777, 778, 793, 795, 797, 799, 800, 811, 812, 814, 815, 816, 817, 834, 836, 837, 841, 843, 844, 848, 850, 851, 191, 319, 400, 427, 429, 430, 539, 35, 36, 52, 57, 58, 134, 706, 174, 211, 233, 173, 668, 835, 852, 776, 533, 838, 273, 274, 320, 451]. Review [465, 564, 569, 666, 733, 736, 452, 464, 516, 590, 604, 756, 588, 592, 647, 652, 682, 330, 364, 365, 628, 712, 651, 516]. Revolution [625]. Rey [328, 678]. Rezeptliteratur [396]. Rhind [418]. Riatinis [550]. Ribera [797]. Ricardi [250]. Richard [605, 197, 291, 838]. Richards [759]. Ridwan [412]. Rif’at [624]. Rignano [423, 130]. Rissalat [670]. Rivière [849]. Roback [358]. Robbins [620]. Robert [519, 840, 404, 528, 568, 757, 811, 330, 302, 508, 204, 304]. Robin [419]. Robinson [850]. Rochester [407]. Roderick [528]. Roger [757, 758]. Rogeri [835]. Rohde [313]. Rohr [274]. rois [495]. Roland [758]. role [139]. Romaine [758]. Rome [174]. romischen [166]. Ronciere [516, 604]. Rosenbusch [605]. Roth [662]. Rouff [301]. Rougier [537]. Rouse [554]. Route [100]. Roy [357]. royale [216]. Rube [570]. Rudolph [841]. Ruhmesblatter [651]. Ruhrah [573]. Rules [505]. Rupert [363]. Rusby [789, 795]. Ruska [731, 799, 118, 83, 448, 650]. Russel [444]. Russell [539, 61, 741, 59]. Russia [128]. Rutbatu’l [377]. Rutbatu’l-Hakim [377]. Ryugaku [646].

S [616, 625, 281, 123, 59, 392, 546, 776]. S. [546, 417]. S.J. [787]. sa [226, 216]. Sablon [140]. Saemtliche [343]. Sage [469]. Sageret [86]. Saint [759, 849, 646]. Saint-Rat [849]. S’aivism [13]. Salernitan [94]. salibus [785]. Salient [325]. salts [641]. Salvatore [533]. Samaj [74]. Samuel [113, 340]. Sanchez [222, 207, 846]. Sanctuaries [624]. Sand [144]. Sandys [165]. Sanskrit [114, 221]. Sanskrit-Philologie [114]. Santiago [583]. sara [15]. Sarkar [6]. Sarton [709, 60]. Sastri [74]. Satire [183]. Satz [293]. Saussure [345]. Savage [239]. savant [299]. savants [212]. Scale [95]. Scellinck [792]. Sceptical [509]. Schiaparelli [584]. Schiffner [837]. Schilderungen [211]. Schisms [116]. Schmidt [163, 524, 166]. Schnitt [647]. Schnitt- [647]. Scholar [389]. Scholarship [165]. Scholastic [642]. Schoy [731, 617]. Schrift [748, 572]. Schriften [247]. Schroeder [16]. Schultz [525]. Schumacher [837]. Schutte [71]. Schwalbe [232]. Schwarze [652]. Schweizerischen [309]. Science [147, 150, 151, 368, 598, 557, 612, 692, 803, 775, 477, 635, 718, 20, 84, 227, 484, 709, 60, 175, 840, 63, 27, 22, 194, 96, 88, 97, 145, 155, 241, 332, 284, 348, 404, 371, 472, 498, 568, 593, 606, 504, 644, 658, 630, 684, 728, 779, 714, 762, 790, 809, 802, 819, 853, 804, 381, 832, 657, 482, 283, 424, 72, 613, 189, 325, 449, 140, 120, 361]. Sciences [798, 167, 168, 397, 290, 139, 735, 761, 142, 493, 237, 190, 77]. Scientific [300, 61, 793, 767, 143, 708]. scientifique [260, 419, 131]. scientifiques [167, 168, 397, 710, 623, 674, 676]. Scientist [389]. Scientists [146, 285, 408]. scienza [482, 319]. scienziati [412, 138]. scolastique [537]. Scot [156, 720, 827]. Scott [528, 759, 559]. Scottish [45]. Scritti [584, 649]. se [788]. Sears [432]. sechste [448]. Second [435, 633, 644, 714, 715]. Secret [19]. Section [765]. Sects [116]. Sedgwick [254, 388]. seeds [385]. Seele [530, 712]. Seelye [116]. Seemann [836]. sein [663, 638]. seine [293, 202, 733]. Selbstbiographie [737, 206]. sélection [21]. Selections [45]. sels [641]. senectutis [835]. Sense [45]. sentiment [576]. Series [644, 658, 728, 790, 809]. serums [277]. ses [744, 578, 641, 615, 218]. Seventeenth [498, 793, 499, 695, 723, 750, 198]. Seventh [22]. Severyns [799]. Shapley [848]. Sharp [215]. Shigaku [121]. Shinsen [707]. Shotwell [270]. Shunjo [646]. Sibyllinischer [794]. Sidney [187]. siècle [157, 4, 301, 198, 342, 148, 344]. siècles [35]. Siegmund [281]. Sigerist [390, 396, 562, 661]. Sign [413]. Signs [251]. Silence [179]. Simon [813]. simultanéité [427]. Sinclair [313]. Singer [78, 491, 562, 800, 743, 189, 280, 742, 416]. Sinibus [250]. sino [272, 53]. Sino-Iranica [53]. Sir [392, 417, 565, 773, 717, 389]. Sirkar [7]. sismique [399]. sismologique [536]. sistema [533]. Sivanath [74]. Sixteenth [598, 473, 472, 551]. Sixth [809, 853, 854]. Sketch [300, 263]. Skizzen [233]. Slawen [525]. Sluse [724]. Small [233]. Smaragdina [650]. Smith [307, 517, 771, 535, 401, 462, 538, 679, 772, 57]. So-Called [30]. Social [571, 433, 431, 433]. sociale [236, 144]. socialistes [405]. sociétés [818]. Societies [793]. Society [612, 692, 803, 477, 635, 718, 120, 504, 802]. sociologie [431]. Sociology [6]. sogenannten [341]. Solar [87]. Solution [30]. Some [601, 614, 788, 548, 82]. sommaire [28]. Son [225, 616, 210]. sondern [837]. Sonnenuhren [380]. sont [788]. Source [848, 247]. sources [738, 579]. South [798]. Souvenirs [457, 352]. space [485]. spatarabischen [380]. Special [27, 538]. spécifique [21]. Speculations [567]. Spence [188]. Spencer [66]. Spengler [271]. Sperling [106]. Speyer [113]. Spinoza [693]. Spiralen [274]. Spiritualism [239]. Sports [740]. Spread [566]. Squares [122]. St [736]. Stahl [578, 641]. Stanford [698]. Stanislao [649]. Stanley [456]. Stapleton [754]. Starr [455]. States [841]. Statistics [193]. Steel [625]. Steensby [100]. Stefanidoy [75]. Steineg [275, 181]. Stellung [312]. Sterndeutung [677]. Sternglaube [677]. Sternhimmel [467]. Sternkampf [794]. Sternwarte [836]. Stevenson [229]. Stillman [492]. Stirling [300, 300]. Stoddard [223]. Stofflers [103]. Stone [192]. Storia [269, 584, 482, 319, 272]. Story [492]. Strange [135]. Strauss [470]. Streeter [661]. Stromayr [647]. structure [226]. Strutt [519, 519]. Strzygowski [529]. Stuart [528]. Studer [621]. studi [482]. Studien [102, 396, 506, 79, 396]. Studienjahre [106]. Studies [506, 482, 571, 759, 189, 449, 36]. Study [46, 813]. Subjects [18]. Submarine [304]. Substance [400]. Sud [9]. Sudhoff [275, 562, 275, 181, 660, 233]. suivie [457]. sulla [584]. Summary [28]. Sunderland [701]. Superintendent [841]. Surgeries [792]. Survey [401, 841]. survivances [744]. Suter [203, 296]. Sutherland [162]. Svami [14]. Swabey [400, 400]. Sydney [361]. Sylvain [774]. Sylvanus [48]. Symbolaì [75]. symbolischen [640]. Symbols [69]. Symmetry [97]. Synthèse [24]. Synthetic [27]. Synthetische [429]. Syphilis [562]. Syrie [421]. syrischen [261]. System [7, 112, 125]. systematischer [273]. système [237]. systèmes [405]. Systems [116, 13].

T [119, 270, 363, 418, 174]. tês [75]. tôn [75]. Table [149, 438, 501, 608, 688, 745, 781, 825, 89]. Tabula [650]. Tacquet [616]. Tafel [837]. Tafeln [203]. Tahafot [732]. Tahir [116]. Tanabe [123]. Tangentes [724]. Tannery [397, 676, 710, 167, 168, 623]. Tantra [14, 15]. Tantrik [15]. Taranatha [15]. Tarragó [797]. tasks [594, 441]. Tatsachen [402]. Tawhid [670]. Taylor [105]. Tchouan [169]. Teaching [37, 695, 750, 155]. Technik [349, 330, 364, 364, 76, 360, 651]. technique [329]. Technischen [177]. Teil [736]. temps [218, 210]. Ten [47, 562, 834, 75]. Tenney [468, 672, 817]. Tenth [145, 754, 804]. Termier [352]. Terminologie [483]. Terminology [267]. Terre [352, 536]. Terrestrial [229]. tetrapodes [398]. Text [549]. Textbook [696]. Texte [396, 261]. Texts [15, 669]. Th [124, 275, 181]. Thalamas [298]. thans [126]. thaumaturges [495]. Their [351, 567, 615]. Theodor [391, 404]. Theodore [840, 568, 811]. Theology [642]. Theophrast [343]. theorem [293]. theoretischen [79]. théorie [541, 21, 328, 641, 615, 751, 427, 29]. Theorien [402]. theories [615, 751, 672]. theorique [616, 406]. Theory [85, 133, 353, 445, 56, 80, 400, 727, 641, 29, 586]. Theosophie [113]. Thienen [792]. Third [504, 658, 762, 478, 763]. Thirteen [348, 699]. Thirteenth [286, 284, 368, 749]. Thomas [625, 171, 219, 566, 699, 792, 698]. Thomas-Stanford [698]. thomisme [537]. Thompson [213, 585]. Thomsen [796]. Thomson [125]. Thorndike [348]. Thought [224, 459, 105]. three [485, 594, 441, 730]. Tien [306]. Time [642, 488]. Times [160, 240]. Tod [652]. Todd [711]. tolerance [547, 547]. Tome [812]. Tools [62]. Topics [538]. Torney [470]. Torricelli [256]. tot [102, 569]. toutes [150, 151]. Tractates [562]. traite [616, 55, 406]. transformisme [393, 398]. transformiste [20]. Translation [204]. Traubenkernen [385]. travailleurs [629]. Travelers [831]. Travellers [452]. Treatise [550, 748]. treatises [834]. Treatment [548]. Treibhauser [529]. tremblements [536]. Treptow [837]. Treviso [379]. Trigonometric [158]. Trigonometrie [292]. trigonometrischen [731]. Tropfke [273]. Tseu [169]. Tusi [290]. Tweedie [300, 46]. Twelfth [241, 700, 68, 785]. Twelve [784]. Twentieth [630, 631]. Twenty [684, 779, 714, 762, 819, 853, 820, 854, 685, 715, 763, 780]. Twenty-Sixth [853, 854]. Two [68, 624]. Tychonis [255]. Types [93]. Tz’u [306].

u [470]. Überblick [275, 181]. übereinstimmenden [525]. Überlieferung [530]. Übersicht [402]. Umrisse [271]. Understanding [468]. Unis [144]. United [841]. Universe [463]. universelle [236]. Unpublished [828]. Unsere [312]. Untergang [271]. Unterricht [83]. Untersuchungen [529, 430]. upon [411]. Urbain [849, 227]. ‘Urdî [836]. Urgent [546]. Urkunden [605]. Urmensch [530]. Urwelt [469]. use [638].

V [457, 305, 534, 605, 595, 436]. Vaccins [277]. vague [86]. Vaisnavism [13]. Val [569]. Valerius [443, 478]. vari [649]. Various [116]. Vassura [256]. Vatican [832]. Vaux [77, 265]. Ve [150]. Vedic [18]. Venice [561]. Ver [201]. verdanken [163]. Verhaltnis [430]. veritable [327]. Vernacular [800]. Verne [849]. Vernunft [320]. Verse [749]. Version [637]. Versuchs [274]. vertebres [398]. vestigia [324]. Veterinary [57]. Veteris [324]. VI [337, 438, 595, 436]. Vialleton [398]. vibrations [424]. Vicissitudes [696]. Viclyaratna [15]. Victor [837]. vida [583]. Vie [481, 457, 523, 327, 151, 51, 214]. Vierendeel [329]. Vietor [458]. View [726, 624]. VII [501, 687]. VIII [710, 608, 687, 504]. Vilhelm [796]. Vinci [101]. Vitalism [80]. Vol [744, 85, 133, 353, 419, 15, 840, 847, 397]. volcanique [399]. Volker [467]. Volkerwanderung [529]. Volkskunde [312]. Volume [152, 90, 242, 288, 337, 438, 501, 608, 688, 745, 781, 825, 782, 856, 709, 401, 538, 504, 636, 710, 773, 747, 800, 784, 504]. Volumes [595, 653, 436, 687, 824]. Vom [420, 673, 837, 129]. vor [837]. Vorlesungen [232, 666, 736]. Vossen [736]. Voyages [104, 838].

W [544, 41, 305, 605, 62, 262, 162, 345, 669, 773, 811]. W. [143]. Waite [19]. Waley [491, 800]. Wallace [444]. Wallingford [197, 291]. Wallis [109]. Walsh [58]. Walter [325, 559, 647, 554]. Walton [378]. Walynforde [250]. Wander [529]. Wandhoff [837]. War [613]. Was [613, 163]. Wave [808, 751]. Wave-Like [808]. Weapons [62]. Weaver [356, 389]. Wein [211]. Weingarten [552]. Wellenburg [605]. Welt [420]. Welten [209]. Weltgeschichte [271, 184]. Weltordnung [525]. Werden [420, 837]. Werk [490]. Werke [343]. Werkstatt [282]. Wesen [420, 330]. Wesendonk [530]. Wessels [452]. West [808]. West-European [808]. Westaway [143]. Western [484]. Westwards [566]. Weyrauch [330]. Wheeler [433, 356, 665]. Whetzel [314]. White [392]. Whitehead [741, 539]. who [788]. Wiborg [626]. wichtiger [293]. Wieleitner [731, 134]. Wilbur [407].

Wilhelm [704, 737, 837]. William [554, 519, 38, 180, 231, 433, 266, 304, 340, 351, 392, 240, 356, 417, 491, 565, 665, 758, 777, 400, 588, 162]. Williams [778]. Williamson [838]. Willis [407]. Wilson [534]. Windisch [114]. Wineland [100]. Winternitz [115]. wir [163]. wirklich [613]. Wirkung [490]. Wissenschaft [384, 613, 590]. Wissenschaften [274]. wissenschaftlichen [451, 733]. Wissler [627]. Witchcraft [852]. Withington [835]. Witz [406]. Wm [842]. Wohlwill [663]. Wolfgang [525]. Woman [458]. Woodhouse [258]. Woods [112]. Work [411, 40, 453, 351, 773]. Works [300, 38, 614, 14]. World [61, 223, 567, 516]. World-Conception [567]. World-Map [516]. Worp [569]. Wren [389]. Wright [488]. Writing [180]. Writings [46, 247]. Wunder [238]. Wundergeburt [238]. Wundergestalt [238].

X [745, 824]. Xaver [794]. Xavier [210]. XI [781, 824]. XIe [35]. XII [825, 824]. XIIe [342, 35]. XIII [419]. XIIth [31]. XIXth [644, 658, 728, 790, 809]. XVIIe [148, 344, 157, 4, 198]. XVIIIe [301, 344]. XVIIth [726]. XVIth [105, 800, 726]. XVth [158, 726]. XXth [644, 658, 728, 790, 809].

Yakutsk [776]. Yakutskim [776]. Years [346, 489, 562]. Yoga [112]. Yoga-System [112]. Yogaku [707]. Yoshio [706]. ỳpò [75]. Yuan [169, 752, 511].

Zachaire [551]. Zacharias [274]. Zakrzewska [458]. Zasshi [121]. Zauberei [628]. Zeit [81, 733]. Zeiten [570, 628]. zeitgenossischen [211]. Zeitlinger [137]. Zeitrechnung [842, 525]. Zeittafeln [228]. Zeno [3]. Zeugnis [467]. Ziergeschichtliche [529]. Zimmer [711]. zira’at [450]. Zoological [765]. zugleich [402]. zum [570, 837]. zur [385, 530, 79, 81, 129, 360, 166, 228, 396, 292, 430, 579, 414, 485, 118]. Zuretti [799]. Zusammenhange [359, 136, 235]. Zusammentreffens [274]. Zuzammenklange [209]. Zwölf [837]. Zygophyllaceae [532].

References

Anonymous:1920:FMa

[1] Anonymous. Front matter. Isis, 3(1):1–2, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

     org/stable/224189.

Sarton:1920:FH

[2] George Sarton. The faith of a humanist. Isis, 3(1):3–6, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

     http://www.jstor.org/stable/224190.

Cajori:1920:PZA

[3] Florian Cajori. The purpose of Zeno’s arguments on motion. Isis, 3(1): 7–20, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224191.

Goblot:1920:MCP

[4] Edmond Goblot and Aug. Georges-Berthier. Le mécanisme cartésien et la physiologie au XVIIe siècle. Isis, 3(1):21–58, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224192. See errata [148].

S:1920:CC

[5] G. S. Chronique et correspondance. Isis, 3(1):59–62, January 1920. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/224193.

Masson-Oursel:1920:BRBa

[6] P. Masson-Oursel. Book review: The Positive Background of Hindu Sociology by Benoy Kumar Sarkar. Isis, 3(1):63–64, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224194.

Masson-Oursel:1920:BRBb

[7] P. Masson-Oursel. Book review: An Introduction to the Hindu System of Physics by Kishori Lal Sirkar. Isis, 3(1):64–65, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224195.

Masson-Oursel:1920:BRBc

[8] P. Masson-Oursel. Book review: Quarante ans d’Indianisme by Oeuvres d’Auguste Barth. Isis, 3(1):65–66, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224196.

Masson-Oursel:1920:BRBd

[9] P. Masson-Oursel. Book review: Archeologie du Sud de l’Inde by G. Jouveau-Dubreuil. Isis, 3(1):66–67, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224197.

Oursel:1920:BRB

[10] P. Masson Oursel. Book review: A History of Hindu Chemistry by Praphulla Chandra Ray. Isis, 3(1):68–73, January 1920. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224198.

Anonymous:1920:LID

[11] Anonymous. L’oeuvre indologique du by Paul carus. Isis, 3(1):73–74, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224199.

Masson-Oursel:1920:BRBe

[12] P. Masson-Oursel. Book review: The Gospel of Buddha by P. Carus. Isis, 3(1):74–76, January 1920. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224200.

Masson-Oursel:1920:BRBf

[13] P. Masson-Oursel. Book review: Vaisnavism, S’aivism and Minor Religious Systems by R. G. Bhandarkar. Isis, 3(1):76–77, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224201.

Masson-Oursel:1920:BRBg

[14] P. Masson-Oursel. Book review: Works on Tantra by Arthur Avalon; Ellen Avalon; Purnananda Svami. Isis, 3(1):77–81, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224202.

Masson-Oursel:1920:BRBh

[15] P. Masson-Oursel. Book review: Tantrik Texts: Vol. III. — Prapancha sara tantra by Arthur Avalon; Taranatha Viclyaratna. Isis, 3(1):81, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224203.

Masson-Oursel:1920:BRBi

[16] P. Masson-Oursel. Book review: Reden und Aufsatze by Leopold von Schroeder. Isis, 3(1):81–82, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224204.

Masson-Oursel:1920:BRBj

[17] P. Masson-Oursel. Book review: Ancient India by E. J. Rapson. Isis, 3(1): 83–84, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224205.

Masson-Oursel:1920:BRBk

[18] P. Masson-Oursel. Book review: Vedic Index of Names and Subjects by A.-A. Macdonell; A.-B. Keith. Isis, 3(1):84–87, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224206.

Masson-Oursel:1920:BRBl

[19] P. Masson-Oursel. Book review: The Secret Doctrine in Israel by Arthur Edward Waite. Isis, 3(1):87–88, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224207.

Guinet:1920:BRBa

[20] L. Guinet. Book review: Devoirs et perils biologiques by Grasset, J.; La Biologie humaine; L’idéalisme positif; Science et philosophie; Le dogme transformiste. Isis, 3(1):88, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224208.

Guinet:1920:BRBb

[21] L. Guinet. Book review: La sélection du plasma spécifique by Legrand, L.; L’emboîtement des plasmas. Aperçu d’une théorie mécanique de l’hérédité; Les caractères biochimiques de l’espèce; Une conception biologique nouvelle de la cellule. Isis, 3(1):88–89, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224209.

Sarton:1920:SCB

[22] George Sarton. Seventh critical bibliography of the history, philosophy and organization of science and of the history of civilization (to June 1919). Isis, 3(1):90–154, January 1920. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224210.

Anonymous:1920:ECA

[23] Anonymous. Errata: Comment augmenter le rendement intellectuel de l’humanite? Isis, 3(1):155, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224211.

Anonymous:1920:ESG

[24] Anonymous. Errata: La synthèse géologique de 1775 à 1918. Isis, 3(1): 155, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224212.

Anonymous:1920:BMa

[25] Anonymous. Back matter. Isis, 3(1):??, January 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224213.

Anonymous:1920:FMb

[26] Anonymous. Front matter. Isis, 3(2):157–158, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224001.

Sarton:1920:SBS

[27] George Sarton. Synthetic bibliography, with special reference to the history of science. Isis, 3(2):159–170, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224002.

Masson-Oursel:1920:BSI

[28] P. Masson-Oursel. Bibliographie sommaire de l’Indianisme. (French) [Summary bibliography of Indianism]. Isis, 3(2):171–218, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224003.

Suali:1920:ETC

[29] L. Suali. Essai sur la théorie de la connaissance dans la philosophie indienne. (French) [Essay on the theory of knowledge in Indian philosophy]. Isis, 3(2):219–254, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224004.

Child:1920:DFS

[30] J. M. Child. Did Fermat have a solution of the so-called Pellian equation? Isis, 3(2):255–262, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224005.

Singer:1920:DME

[31] Charles Singer. Daniel of Morley. An English philosopher of the XIIth Century. Isis, 3(2):263–269, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224006.

S:1920:NC

[32] G. S. and Dorothea Waley Singer. Notes and correspondence. Isis, 3 (2):270–276, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224007.

J:1920:BRBa

[33] J. Book review: The Flying Arrow; An Anachronism by Philip E. B. Jourdain. Isis, 3(2):277–278, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224008.

Sarton:1920:BRBa

[34] George Sarton. Book review: Meteorologicorum Libri Quattuor by Aristotelis; F. H. Fobes. Isis, 3(2):278–279, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224009.

Singer:1920:BRBa

[35] Charles Singer. Book review: Les médecins dans l’ouest de la France aux XIe et XIIe siècles by Dubreuil-Chambardel, L. Isis, 3(2):279–280, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224010.

Singer:1920:BRBb

[36] Charles Singer. Book review: Studies in English Franciscan History by A. G. Little. Isis, 3(2):280–282, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224011.

J:1920:BRBb

[37] J. Book review: Oughtred’s Ideas and Influence on the Teaching of Mathematics by F. Cajori. Isis, 3(2):282, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224012.

J:1920:BRBc

[38] J. Book review: The Works of William Oughtred by F. Cajori; William Oughtred. Isis, 3(2):282–283, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224013.

J:1920:BRBd

[39] J. Book review: The Lectiones Geometricae of Isaac Barrow by J. M. Child. Isis, 3(2):283, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224014.

J:1920:BRBe

[40] J. Book review: Essays on the Life and Work of Newton by Augustus De Morgan; Philip E. B. Jourdain. Isis, 3(2):283–285, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224015.

J:1920:BRBf

[41] J. Book review: John Napier and the Invention of Logarithms, 1614 by E. W. Hobson. Isis, 3(2):285–286, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224016.

Anonymous:1920:BHE

[42] Anonymous. Newton’s Hypotheses of Ether and of Gravitation from 1672 to 1679 by Philip E. B. Jourdain; Newton. Isis, 3(2):286–287, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224017.

J:1920:BRBg

[43] J. Book review: Newton’s Hypotheses of Ether and of Gravitation from 1679 to 1693 by Philip E. B. Jourdain; Newton. Isis, 3(2):287–288, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224018.

J:1920:BRBh

[44] J. Book review: Newton’s Hypotheses of Ether and of Gravitation from 1693 to 1726 by Philip E. B. Jourdain; Newton. Isis, 3(2):288–289, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224019.

J:1920:BRBi

[45] J. Book review: Selections from the Scottish Philosophy of Common Sense by G. A. Johnston. Isis, 3(2):289, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224020.

J:1920:BRBj

[46] J. Book review: A Study of the Life and Writings of Colin Maclaurin by Chas. Tweedie. Isis, 3(2):289–290, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224021.

Sarton:1920:BRBb

[47] George Sarton. Book review: Lectures on Ten British Physicists of the Nineteenth Century by Alexander Macfarlane. Isis, 3(2):291, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224022.

S:1920:BRBa

[48] G. S. Book review: The Inscriptions at Copan by Sylvanus Griswold Morley. Isis, 3(2):292–294, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224023.

Guinet:1920:BRBc

[49] L. Guinet. Book review: L’Heredite morbide by M. Appert. Isis, 3(2): 295, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224024.

Guinet:1920:BRBd

[50] L. Guinet. Book review: Les problèmes de l’Heredite experimentale by Blaringhem, L. Isis, 3(2):295–296, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224025.

Guinet:1920:BRBe

[51] L. Guinet. Book review: La Matiere et la Vie by Guilleminot, H. Isis, 3(2):296–297, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224026.

Singer:1920:BRBc

[52] Charles Singer. Book review: Outlines of the History of Botany by Harvey-Gibson, R. J. Isis, 3(2):297–299, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224027.

Sarton:1920:BRBc

[53] George Sarton. Book review: Sino-Iranica by Berthold Laufer. Isis, 3 (2):299–302, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224028.

Sarton:1920:BRBd

[54] George Sarton. Book review: Entstehung und Ausbreitung der Alchemie by Edmund O. von Lippmann. Isis, 3(2):302–305, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224029.

Sarton:1920:BRBe

[55] George Sarton. Book review: Traite de Logique by Edmond Goblot. Isis, 3(2):306–307, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224030.

J:1920:BRBk

[56] J. Book review: The Theory of Proportion by M. J. M. Hill. Isis, 3(2): 307, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224031.

Singer:1920:BRBd

[57] Charles Singer. Book review: The Early History of Veterinary Literature and Its British Development by F. Smith. Isis, 3(2):307–308, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224032.

Singer:1920:BRBe

[58] Charles Singer. Book review: Mediaeval Medicine by James J. Walsh. Isis, 3(2):308–309, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224033.

S:1920:BRBb

[59] G. S. Book review: Discovery by A. S. Russell. Isis, 3(2):309–310, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224034.

Jourdain:1920:BRBa

[60] Ph. E. B. Jourdain. Book review: The History of Science by George Sarton. Isis, 3(2):310–311, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224035.

Jourdain:1920:BRBb

[61] Ph. E. B. Jourdain. Book review: Our Knowledge of the External World as a Field for Scientific Method in Philosophy by Bertrand Russell. Isis, 3(2):311–314, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224036.

S:1920:BRBc

[62] G. S. Book review: Tools and Weapons by W. M. Flinders Petrie. Isis, 3(2):314–315, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224037.

Sarton:1920:ECB

[63] George Sarton. Eighth critical bibliography of the history, philosophy and organization of science and of the history of civilization (to April 1920). Isis, 3(2):316–371, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224038.

Anonymous:1920:BMb

[64] Anonymous. Back matter. Isis, 3(2):372, Autumn 1920. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224039.

Anonymous:1921:FMa

[65] Anonymous. Front matter. Isis, 3(3):373–374, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224045.

Sarton:1921:HS

[66] George Sarton. Herbert Spencer 1820–1903. Isis, 3(3):375–390, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224046.

Goblot:1921:GM

[67] Edmond Goblot. Gaston Milhaud. Isis, 3(3):391–395, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224047.

Karpinski:1921:TTC

[68] Louis C. Karpinski. Two Twelfth Century algorisms. Isis, 3(3):396–413, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224048.

Cajori:1921:HSF

[69] Florian Cajori. History of symbols for \underline{n} = factorial. Isis, 3(3):414–418, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224049.

S:1921:NC

[70] C. S. and G. Sarton. Notes and correspondence. Isis, 3(3):419–421, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224050.

S:1921:BRBa

[71] G. S. Book review: Ptolemy’s Maps of Northern Europe by Schutte, Gudmund. Isis, 3(3):422–424, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224051.

S:1921:BRBb

[72] G. S. Book review: Histoire de la science nautique portugaise by Bensaude, Joaquim. Isis, 3(3):424–426, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224052.

Guinet:1921:BRBa

[73] L. Guinet. Book review: La Molecule chimique by R. Lespieau. Isis, 3 (3):426–428, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224053.

S:1921:BRBc

[74] G S. Book review: History of the Brahmo Samaj by Sivanath Sastri. Isis, 3(3):428–429, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224054.

S:1921:BRBd

[75] G. S. Book review: Symbolaì eìs tèn ístorían tôn fysikôn èpisthmôn kaì ìdíos tês chymeías ỳpò, by Michael K. Stefanidoy. Isis, 3(3):430–433, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224055.

S:1921:BRBe

[76] G. S. Book review: Antike Technik by Hermann Diels. Isis, 3(3):433– 434, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224056.

S:1921:BRBf

[77] G. S. Book review: Notes d’histoire des sciences by Carra de Vaux, Bernard. Isis, 3(3):434–435, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224057.

S:1921:BRBg

[78] G. S. Book review: Early English Magic and Medicine by Singer, Charles. Isis, 3(3):435–437, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224058.

S:1921:BRBh

[79] G. S. Book review: Studien zur Geschichte der theoretischen Geodasie by Franz Joh. Muller. Isis, 3(3):438–439, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224059.

S:1921:BRBi

[80] G. S. Book review: The History and Theory of Vitalism by Hans Driesch; C. K. Ogden. Isis, 3(3):439–440, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224060.

S:1921:BRBj

[81] G. S. Book review: Geschichte der organischen Chemie von altester Zeit bis zur Gegenwart by Edv. Hjelt. Isis, 3(3):440–443, Summer 1921. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/224061.

S:1921:BRBk

[82] G. S. Book review: Some Fundamental Problems in Chemistry Old and New by Letts, E. A. Isis, 3(3):443, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224062.

S:1921:BRBl

[83] G. S. Book review: Methodik des mineralogisch-geologischen Unterricht by Julius Ruska. Isis, 3(3):443–444, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224063.

Guinet:1921:BRBb

[84] Leon Guinet. Book review: La genese de la Science des Cristaux by Helene Metzger. Isis, 3(3):445–446, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224064.

Child:1921:BRBa

[85] J. M. Child. Book review: History of the Theory of Numbers, Vol. I: Divisibility and Primality by Leonard Eugene Dickson. Isis, 3(3):446– 448, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224065.

Guinet:1921:BRBc

[86] Leon Guinet. Book review: La vague mystique by Jules Sageret. Isis, 3(3):448–449, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224066.

S:1921:BRBm

[87] G. S. Book review: Solar Biology by Butler, Hiram E. Isis, 3(3):449– 450, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224067.

Sarton:1921:NCB

[88] George Sarton. Ninth critical bibliography of the history, philosophy and organization of science and of the history of civilization. Isis, 3(3):451– 497, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224068.

Sarton:1921:TCF

[89] George Sarton. Table of contents of the first nine critical bibliographies. Isis, 3(3):498–503, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224069.

Anonymous:1921:VI

[90] Anonymous. Volume information. Isis, 3(3):504–569, Summer 1921. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/224044.

Anonymous:1921:BMa

[91] Anonymous. Back matter. Isis, 3(3):570, Summer 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224070.

Anonymous:1921:FMb

[92] Anonymous. Front matter. Isis, 4(1):1–4, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224071.

Miller:1921:DTM

[93] G. A. Miller. Different types of mathematical history. Isis, 4(1):5–12, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224072.

Johnsson:1921:HHD

[94] J. W. S. Johnsson. Henrik Harpestræng. a Danish Salernitan. Isis, 4(1): 13–16, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224073.

Cajori:1921:NFS

[95] Florian Cajori. Note on the Fahrenheit scale. Isis, 4(1):17–22, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224074.

Sarton:1921:IHP

[96] George Sarton. Introduction to the history and philosophy of science (preliminary note). Isis, 4(1):23–31, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224075.

Sarton:1921:PSA

[97] George Sarton. The principle of symmetry and its applications to science and to art. Isis, 4(1):32–38, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224076.

Sarton:1921:NC

[98] G. Sarton and J. W. S. Johnsson. Notes and correspondence. Isis, 4(1): 39–46, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224077.

Sarton:1921:BRBa

[99] George Sarton. Book review: De la Nature by Lucrece; Alfred Ernout. Isis, 4(1):47–48, May 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224078.

S:1921:BRBn

[100] G. S. Book review: The Norsemen’s Route from Greenland to Wineland by Steensby, H. P. Isis, 4(1):48, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224079.

S:1921:BRBo

[101] G. S. Book review: Per il IV degrees Centenario della Morte di Leonardo da Vinci by Mario Cermenati. Isis, 4(1):48–49, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224080.

S:1921:BRBp

[102] G. S. Book review: Historische Studien. Bijdragen tot de Kennis van de geschiedenis der natuurwetenschappen in de Nederlanden gedurende de 16e en 17e eeuw by Jaeger, F. M. Isis, 4(1):50–51, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224081.

S:1921:BRBq

[103] G. S. Book review: Aus der Blutezeit der Astrometeorologie. J. Stofflers Prognose für des Jahr 1524 by Hellmann, G. Isis, 4(1):51–52, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224082.

S:1921:BRBr

[104] G. S. Book review: Cartes marines, constructions navales. Voyages de découverte chez les Normands, 1500–1650 by Anthiaume, L’Abbé A. Isis, 4(1):52, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224083.

Sarton:1921:BRBb

[105] George Sarton. Book review: Thought and Expression in the XVIth Century by Taylor, Henry Osborn. Isis, 4(1):52–54, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224084.

Sarton:1921:BRBc

[106] G. Sarton. Book review: Studienjahre by Otto Sperling. Isis, 4(1):54, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224085.

Sarton:1921:BRBd

[107] George Sarton. Book review: Campanella by Blanchet, Leon. Isis, 4(1): 55–56, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224086.

Sarton:1921:BRBe

[108] G. Sarton. Book review: Les civilisations prehelleniques dans le bassin de la mer Egee by Dussaud, Rene. Isis, 4(1):56–57, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224087.

Sarton:1921:BRBf

[109] George Sarton. Book review: An Egyptian Hieroglyphic Dictionary by Budge, E. A. Wallis. Isis, 4(1):57–58, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224088.

Sarton:1921:BRBg

[110] George Sarton. Book review: Discovery in Greek Lands by Marshall, F. H. Isis, 4(1):59, May 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224089.

S:1921:BRBs

[111] G. S. Book review: Hellenism in Ancient India by Banerjee, Gauranga Nath. Isis, 4(1):59–60, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224090.

Masson-Oursel:1921:BRBa

[112] P. Masson-Oursel. Book review: The Yoga-System of Patanjali by James Haughton Woods. Isis, 4(1):60–61, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224091.

Masson-Oursel:1921:BRBb

[113] P. Masson-Oursel. Book review: Die indische Theosophie by Jacob Samuel Speyer. Isis, 4(1):61, May 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224092.

Masson-Oursel:1921:BRBc

[114] P. Masson-Oursel. Book review: Geschichte der Sanskrit-Philologie und indischen Altertumskunde by Ernst Windisch. Isis, 4(1):61–62, May 1921.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224093.

Masson-Oursel:1921:BRBd

[115] P. Masson-Oursel. Book review: Geschichte der indischen Literatur by M. Winternitz. Isis, 4(1):62, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224094.

Sarton:1921:BRBh

[116] George Sarton. Book review: Moslem Schisms and Sects (Al-Fark Bain al-Firak) being the History of the Various Philosophic Systems Developed in Islam by Abu-Mansur ’abd-al-Kahir ibn-Tahir al Baghdadi; Kate Chambers Seelye. Isis, 4(1):63–64, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224095.

Macdonald:1921:BRBa

[117] Duncan B. Macdonald. Book review: Le Dogme et la Loi de l’Islam by Goldziher, Ignaz; Felix Arin. Isis, 4(1):64–66, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224096.

Karpinski:1921:BRB

[118] Louis C. Karpinski. Book review: Zur altesten arabischen Algebra und Rechenkunst by Julius Ruska. Isis, 4(1):67–70, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224097.

Mikami:1921:BRBa

[119] Yoshio Mikami. Book review: History of Japanese Mathematics by T. Endo. Isis, 4(1):70–72, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224098.

Mikami:1921:BRBb

[120] Y. Mikami. Book review: Science in Japan by Japan Civilization Society. Isis, 4(1):72–74, May 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224099.

Mikami:1921:BRBc

[121] Y. Mikami. Book review: The Causes of the Development of Mathematics in Japan, The Shigaku Zasshi by Y. Mikami. Isis, 4(1):75–76, May 1921.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224100.

Mikami:1921:BRBd

[122] Y. Mikami. Book review: Magic Squares in Japanese Mathematics by Y. Mikami. Isis, 4(1):76–77, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224101.

Mikami:1921:BRBe

[123] Yoshio Mikami. Book review: Music in Japan by S. Kanetsune; Lectures on Japanese Music by H. Tanabe. Isis, 4(1):77–81, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224102.

S:1921:BRBt

[124] G. S. Book review: Kunstgewerbliche Altertumer und Kuriositaten by Graesse, J. G. Th.; F. Jaennicke; Franz M. Feldhaus. Isis, 4(1):82, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224103.

S:1921:BRBu

[125] G. S. Book review: The System of Animate Nature by Thomson J. Arthur. Isis, 4(1):82–83, May 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224104.

S:1921:BRBv

[126] G. S. Book review: Elementen en Atomen eens en thans by Jaeger, F. M. Isis, 4(1):83, May 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224105.

Guinet:1921:BRBd

[127] L. Guinet. Book review: Histoire de la Chimie by M. Delacre. Isis, 4 (1):84, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224106.

Sarton:1921:BRBi

[128] George Sarton. Book review: Russia, Mongolia, China by Baddeley, John F. Isis, 4(1):85–88, May 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224107.

S:1921:BRBw

[129] G. S. Book review: Vom Altertum zur Gegonwart by Eduard Norden. Isis, 4(1):89, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224108.

Sarton:1921:BRBj

[130] G. Sarton. Book review: Psychologie du Raisonnement by Eugenio Rignano. Isis, 4(1):90–93, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224109.

Sarton:1921:BRBk

[131] G. Sarton. Book review: L’Ideal scientifique des mathematiciens by Pierre Boutroux. Isis, 4(1):93–96, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224110.

Child:1921:BRBb

[132] J. M. Child. Book review: Les Principes de l’Analyse Mathematique by Pierre Boutroux. Isis, 4(1):96–107, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224111.

Child:1921:BRBc

[133] J. M. Child. Book review: History of the Theory of Numbers: Vol. II “Diophantine Analysis” by L. E. Dickson. Isis, 4(1):107–108, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224112.

Smith:1921:BRB

[134] David Eugene Smith. Book review: Geschichte der Mathematik by Heinrich Wieleitner. Isis, 4(1):109, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224113.

Sarton:1921:BRBl

[135] G. Sarton. Book review: Bibliographisher Fuhrer durch die medizinische Literatur by Oscar [Christian Strange] Preisler. Isis, 4(1):109–110, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224114.

Sarton:1921:BRBm

[136] G. Sarton. Book review: Die Naturwissenschaften in ihrer Entwicklung und in ihrem Zusammenhange by Dannemann, Friedrich. Isis, 4(1):110– 111, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224115.

Sarton:1921:BRBn

[137] George Sarton. Book review: Bibliotheca Chemico-Mathematica by Zeitlinger, Heinrich. Isis, 4(1):111–112, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224116.

Sarton:1921:BRBo

[138] Sarton. Book review: Gli scienziati italiani dall’ inizio del medio evo ai nostri giorni by Aldo Mieli. Isis, 4(1):112–114, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224117.

Guinet:1921:BRBe

[139] L. Guinet. Book review: Le role des sciences dans l’education by R. Paucot. Isis, 4(1):114–115, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224118.

Guinet:1921:BRBf

[140] L. Guinet. Book review: L’unité de la science by M. Leclerc du Sablon. Isis, 4(1):115, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224119.

Guinet:1921:BRBg

[141] L. Guinet. Book review: Force et Cause by Fr. Houssay. Isis, 4(1): 115–117, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224120.

Sarton:1921:BRBp

[142] George Sarton. Book review: Classification des sciences. Les idées maîtresses des sciences et leurs rapports by Naville, Adrien. Isis, 4(1): 118, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224121.

Sarton:1921:BRBq

[143] George Sarton. Book review: Scientific Method by Westaway, F. W. Isis, 4(1):119–122, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224122.

Sarton:1921:BRBr

[144] George Sarton. Book review: Organisation industrielle, médecine sociale et education civique en Angleterre et aux Etats-Unis by Sand, Rene. Isis, 4(1):122–123, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224123.

Sarton:1921:TCB

[145] George Sarton. Tenth critical bibliography of the history, philosophy and organization of science and of the history of civilization (to April 1921). Isis, 4(1):124–160, May 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224124.

Anonymous:1921:MBC

[146] Anonymous. Materials for the biography of contemporary scientists (chiefly obituary notices). Isis, 4(1):161–219, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224125.

Anonymous:1921:EEC

[147] Anonymous. Errata: Eighth critical bibliography of the history, philosophy and organization of science and of the history of civilization (to April 1920). Isis, 4(1):220, May 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224129.

Anonymous:1921:EMC

[148] Anonymous. Errata: Le Mécanisme Cartésien et la physiologie au XVIIe siècle. Isis, 4(1):220, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/ 224128. See [4].

Anonymous:1921:ETC

[149] Anonymous. Errata: Table of contents of the first nine critical bibliographies. Isis, 4(1):220, May 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224130.

Anonymous:1921:EVBa

[150] Anonymous. Errata: Ve bibliographie critique de toutes les publications relatives à l’histoire, á la philosophie et à l’organisation de la science. Isis, 4(1):220, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224126.

Anonymous:1921:EVBb

[151] Anonymous. Errata: VIe bibliographie critique de toutes les publications relatives à l’histoire, la philosophie et l’organisation de la science (jusqu’en août 1914). Isis, 4(1):220, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224127.

Anonymous:1921:EVI

[152] Anonymous. Errata: Volume information. Isis, 4(1):220, May 1921. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/224131.

Anonymous:1921:BMb

[153] Anonymous. Back matter. Isis, 4(1):??, May 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224132.

Anonymous:1921:FMc

[154] Anonymous. Front matter. Isis, 4(2):221–223, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224246.

Sarton:1921:THS

[155] George Sarton. The teaching of the history of science. Isis, 4(2):225– 249, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224247.

Haskins:1921:MSF

[156] Charles H. Haskins. Michael Scot and Frederick II. Isis, 4(2):250–275, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224248.

Boutroux:1921:LMF

[157] Pierre Boutroux. L’enseignement de la mècanique en France au XVIIe siècle. Isis, 4(2):276–294, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224249.

Bond:1921:DTM

[158] John David Bond. The development of trigonometric methods down to the close of the XVth Century. Isis, 4(2):295–323, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224250.

Millikan:1921:NC

[159] R. A. Millikan and L. Guinet. Notes and correspondence. Isis, 4(2):324– 325, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224251.

S:1921:BRBx

[160] G. S. Book review: Life and Times of Ambroise Pare (1510–1590) by Packard, Francis R. Isis, 4(2):326–327, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224252.

Sarton:1921:BRBs

[161] George Sarton. Book review: Buffon by Louis Dimier. Isis, 4(2):327– 328, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224253.

Sarton:1921:BRBt

[162] George Sarton. Book review: William Sutherland by W. A. Osborne. Isis, 4(2):328–330, October 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224254.

Bloch:1921:BRBa

[163] Ernst Bloch. Book review: Was wir Ernst Haeckel verdanken by Heinrich Schmidt. Isis, 4(2):330–335, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224255.

Bloch:1921:BRBb

[164] Ernst Bloch. Book review: Deutschlands Chemische Industrie 1888–1913 by B. Lepsius. Isis, 4(2):335, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224256.

Sarton:1921:BRBu

[165] G. Sarton. Book review: A History of Classical Scholarship by John Edwin Sandys. Isis, 4(2):335–336, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224257.

Sarton:1921:BRBv

[166] George Sarton. Book review: Kulturhistorische Beitrage zur Kenntnis des griechischen und romischen Altertums by Max C. P. Schmidt. Isis, 4(2):336–338, October 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224258.

S:1921:BRBy

[167] G. S. Book review: Mémoires scientifiques: III. Sciences exactes dans l’antiquité by Paul Tannery. Isis, 4(2):338–341, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224259.

S:1921:BRBz

[168] G. S. Book review: Mémoires scientifiques: IV. Sciences exactes chez les Byzantins (1884–1919) by Paul Tannery. Isis, 4(2):342–345, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224260.

Sarton:1921:BRBw

[169] G. Sarton. Book review: Kiai Tseu Yuan Houa Tchouan by Raphael Petrucci. Isis, 4(2):345–347, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224261.

Karpinskl:1921:BRB

[170] Louis C. Karpinskl. Book review: I Prearistotelici by Mieli, Aldo. Isis, 4(2):347–349, October 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224262.

Macdonald:1921:BRBb

[171] D. B. Macdonald. Book review: Arabian Medicine by Edward G. Browne; Thomas Adams’s. Isis, 4(2):349–350, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224263.

Sarton:1921:BRBx

[172] George Sarton. Book review: Les peuples d’Extreme-Orient. — Le Japon by Hovelaque, Emile. Isis, 4(2):351–352, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224264.

Thorndike:1921:BRB

[173] Lynn Thorndike. Book review: Mediaeval Contributions to Modern Civilisation by F. J. C. Hearnshaw. Isis, 4(2):352–354, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224265.

Streeter:1921:BRB

[174] Edw. C. Streeter. Book review: Greek Medicine in Rome by T. Clifford Allbutt. Isis, 4(2):355–357, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224266.

S:1921:BRBaa

[175] G. S. Book review: History and Bibliography of Anatomic Illustration in Its Relation to Anatomic Science and the Graphic Arts by Ludwig Choulant; Mortimer Frank. Isis, 4(2):357–359, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224267.

Guinet:1921:BRBh

[176] L. Guinet. Book review: Le mouvement biologique en Europe by G. Bohn. Isis, 4(2):359–360, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224268.

Bloch:1921:BRBc

[177] Ernst Bloch. Book review: Professor der Chemie an der Technischen Hochschule in Dresden by Ernst; Meyer. Isis, 4(2):360–361, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224269.

Hopkins:1921:BRB

[178] Arthur John Hopkins. Book review: Geschichte der Organischen Chemie by Carl Graebe. Isis, 4(2):361–365, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224270.

Sarton:1921:BRBy

[179] G. Sarton. Book review: The Lands of Silence by Markham, Clements R. Isis, 4(2):365–367, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224271.

Libby:1921:BRBa

[180] Walter Libby. Book review: A History of the Art of Writing by William A. Mason. Isis, 4(2):367–368, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224272.

Sarton:1921:BRBz

[181] George Sarton. Book review: Geschichte der Medizin im Überblick mit Abbildungen by Th. Meyer-Steineg; Karl Sudhoff. Isis, 4(2):368–369, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224273.

S:1921:BRBab

[182] G. S. Book review: 3000 Jahre Medizin by Hubotter, Franz. Isis, 4(2): 369–370, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224274.

S:1921:BRBac

[183] G. S. Book review: Die Karikatur und Satire in der Medizin by Hollander, Eugen. Isis, 4(2):370–371, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224275.

Sarton:1921:BRBaa

[184] George Sarton. Book review: Die Gifte in der Weltgeschichte by L. Lewin. Isis, 4(2):371–373, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224276.

Sarton:1921:BRBab

[185] G. Sarton. Book review: The Idea of Progress by John Bagnell Bury. Isis, 4(2):373–375, October 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224277.

Sarton:1921:BRBac

[186] George Sarton. Book review: Die Quantentheorie by Reiche, Fritz. Isis, 4(2):375–376, October 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224278.

Libby:1921:BRBb

[187] Walter Libby. Book review: A History of Psychology by George Sidney Brett. Isis, 4(2):376–378, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224279.

Sarton:1921:BRBad

[188] George Sarton. Book review: An Introduction to Mythology by Lewis Spence. Isis, 4(2):378–380, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224280.

S:1921:BRBad

[189] G. S. Book review: Studies in the History and Method of Science by Charles Singer. Isis, 4(2):380–382, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224281.

Metzger:1921:BRB

[190] Helene Metzger. Book review: De l’explication dans les sciences by Emile Meyerson. Isis, 4(2):382–385, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224282.

Sheffer:1921:BRB

[191] H.-M. Sheffer. Book review: Physik und Hypothese by Hugo Dingler. Isis, 4(2):385, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224283.

Sarton:1921:BRBae

[192] George Sarton. Book review: A History of Labour by Stone Gilbert. Isis, 4(2):385–386, October 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224284.

Sarton:1921:BRBaf

[193] G. Sarton. Book review: The History of Statistics by John Koren. Isis, 4(2):387–389, October 1921. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224285.

Sarton:1921:ECB

[194] George Sarton and Elizabeth Gilpatrick. Eleventh critical bibliography of the history and philosophy of science and of the history of civilization. (to October 1921). Isis, 4(2):390–453, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224286.

Anonymous:1921:BMc

[195] Anonymous. Back matter. Isis, 4(2):??, October 1921. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224287.

Anonymous:1922:FM

[196] Anonymous. Front matter. Isis, 4(3):455–458, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/223822.

Bond:1922:RW

[197] John David Bond. Richard Wallingford: (1292?-1335). Isis, 4(3):459– 465, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223823.

Metzger:1922:LDR

[198] Helene Metzger. L’Évolution du regne métallique d’après les alchimistes du XVIIe siècle. (French) [The evolution of the realm of metal after the alchemists of the Seventeenth Century]. Isis, 4(3):466–482, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223824.

Cajori:1922:HC

[199] Florian Cajori. On the history of caloric. Isis, 4(3):483–492, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223825.

S:1922:NC

[200] G. S., L. Guinet, and Florian Cajori. Notes and correspondence. Isis, 4(3): 493–498, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223826.

Guinet:1922:BRBa

[201] L. Guinet. Book review: Les Oeuvres completes d’Archimede by P. Ver Eecke. Isis, 4(3):499–500, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223827.

Libby:1922:BRBa

[202] Walter Libby. Book review: Plinius und seine Naturgeschichte in ihrer Bedeutung für die Gegenwart by Friedrich Dannemann. Isis, 4(3):501– 502, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223828.

Ruska:1922:BRBa

[203] Julius Ruska. Book review: Die astronomischen Tafeln by H. Suter; Muhammed ibn Musa al-Khwarizmi; Maslama ibn Ahmed al-Madjriti; Athelhard von Bath; A. Bjornbo; R. Besthorn. Isis, 4(3):502–503, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223829.

Ruska:1922:BRBb

[204] Julius Ruska. Book review: Robert of Chesters Latin Translation of the Algebra of Al-Khowarizmi by Karplnski, Louis-Charles. Isis, 4(3):504– 505, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223830.

Hermannsson:1922:BRB

[205] Halldor Hermannsson. Book review: The Norse Discoverers of America by Gathorne-Hardy, G. M. Isis, 4(3):505–508, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223831.

Macdonald:1922:BRBa

[206] D. B. Macdonald. Book review: Ghazali’s Selbstbiographie by Heinrich Frick. Isis, 4(3):508, April 1922. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223832.

Ruska:1922:BRBc

[207] Julius Ruska. Book review: Compendio de Algebra de Abenbeder by Jose A. Sanchez Perez. Isis, 4(3):509, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223833.

Boutroux:1922:BRB

[208] P. Boutroux. Book review: The Early Mathematical Manuscripts of Leibniz by C. I. Gerhardt; J. M. Child; Leibniz. Isis, 4(3):510–512, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223834.

Sarton:1922:BRBa

[209] Sarton. Book review: Die Zuzammenklange der Welten by Johann Kepler; Otto J. Bryk. Isis, 4(3):512–514, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223835.

Sarton:1922:BRBb

[210] Sarton. Book review: Fichte et son temps by Xavier Leon. Isis, 4(3): 514–515, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223836.

Streeter:1922:BRBa

[211] Edward C. Streeter. Book review: Das alte medizinische Wein in zeitgenossischen Schilderungen by Max Neuburger. Isis, 4(3):515–516, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/223837.

Sarton:1922:BRBc

[212] Sarton. Book review: Correspondance avec divers savants by H. C. Orsted. Isis, 4(3):516–517, April 1922. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223838.

Sarton:1922:BRBd

[213] G. Sarton. Book review: The Age of Invention by Holland Thompson. Isis, 4(3):517–519, April 1922. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223839.

Guinet:1922:BRBb

[214] L. Guinet. Book review: Vie d’Élie Metchnikoff (1845–1916) by Olga Metchnikoff; Histoire d’une idée. L’oeuvre de Metchnikoff by A. Besredka. Isis, 4(3):519–521, April 1922. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223840.

Sarton:1922:BRBe

[215] G. Sarton. Book review: David Livingstone, Missionary and Explorer by J. Alfred Sharp. Isis, 4(3):521–522, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223841.

S:1922:BRB

[216] G. S. Book review: L’Académie royale de Belgique depuis sa fondation (1772–1922) by P. Pelseneer. Isis, 4(3):522–523, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223842.

Hopkins:1922:BRB

[217] A. J. Hopkins. Book review: Die alteste Alchemie by Mrs. Ingeborg Hammer-Jensen. Isis, 4(3):523–530, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223843.

Sarton:1922:BRBf

[218] George Sarton. Book review: Histoire générale de la Chine et de ses relations avec les pays étrangers depuis les temps les plus anciens jusqu’à la chute de la dynastie mandchoue by Henri Cordier. Isis, 4(3):530–532, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223844.

Sarton:1922:BRBg

[219] George Sarton. Book review: A History of Greek Mathematics by Thomas Heath. Isis, 4(3):532–535, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223845.

Masson-Oursel:1922:BRBa

[220] P. Masson-Oursel. Book review: Indian Logic and Atomism by Arthur Berriedale Keith. Isis, 4(3):535–536, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223846.

Masson-Oursel:1922:BRBb

[221] P. Masson-Oursel. Book review: Iranian Influence on Moslem Literature by G. K. Nariman; Literary History of Sanskrit Buddhism by Idem. Isis, 4(3):537, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223847.

Macdonald:1922:BRBb

[222] D. B. Macdonald. Book review: Biografías de matemáticos arabes que florecieron en España by Sanchez Perez, Jose A. Isis, 4(3):537–538, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223848.

Macdonald:1922:BRBc

[223] D. B. Macdonald. Book review: The New World of Islam by Lothrop Stoddard. Isis, 4(3):538–539, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223849.

Macdonald:1922:BRBd

[224] D. B. Macdonald. Book review: Arabic Thought and Its Place in History by De Lacy O’Leary D. D. Isis, 4(3):539–540, April 1922. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223850.

Sarton:1922:BRBh

[225] George Sarton. Book review: L’art chrétien. Son développement iconographique des origines à nos jours by Louis Brehier. Isis, 4(3):540–544, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/223851.

Guinet:1922:BRBc

[226] L. Guinet. Book review: L’Atome, sa structure, sa forme by Achalme. Isis, 4(3):544–545, April 1922. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223852.

Guinet:1922:BRBd

[227] L. Guinet. Book review: Les Disciplines d’une Science: La Chimie by G. Urbain. Isis, 4(3):545–548, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223853.

Sarton:1922:BRBi

[228] George Sarton. Book review: Zeittafeln zur Geschichte der organischen Chemie by Edmund O. von Lippmann. Isis, 4(3):548–549, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223854.

Sarton:1922:BRBj

[229] George Sarton. Book review: Terrestrial and Celestial Globes by Stevenson, Edward Luther. Isis, 4(3):549–553, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223855.

Sarton:1922:BRBk

[230] George Sarton. Book review: An Introduction to the History of Medicine by Fielding H. Garrison. Isis, 4(3):554–556, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223856.

Libby:1922:BRBb

[231] Walter Libby. Book review: The Evolution of Modern Medicine by William Osler. Isis, 4(3):556–557, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223857.

Sarton:1922:BRBl

[232] George Sarton. Book review: Vorlesungen über Geschichte der Medizin by Ernst Schwalbe. Isis, 4(3):557–558, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223858.

Streeter:1922:BRBb

[233] Edward C. Streeter. Book review: Skizzen, Small Octavo by Karl Sudhoff. Isis, 4(3):558–560, April 1922. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223859.

Bloch:1922:BRB

[234] Ernst Bloch. Book review: Die Prinzipien der physikalischen Optik, historisch und erkenntnispsychologisch entwickelt by Mach, Ernst. Isis, 4(3): 560–562, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223860.

Sarton:1922:BRBm

[235] George Sarton. Book review: Die Naturwissenschaften in ihrer Entwicklung und in ihrem Zusammenhange by Friedrich Dannemann; Galilei. Isis, 4(3):563–564, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223861.

Guinet:1922:BRBe

[236] L. Guinet. Book review: L’Évolution universelle. Exposè des preuves et des lois de l’èvolution mondiale et des èvolutions particulières (inorganique, intellectuelle et sociale) I. L’Évolution mondiale, inorganique et organique by Br. Petronievics. Isis, 4(3):564–565, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223862.

Guinet:1922:BRBf

[237] L. Guinet. Book review: Le système des sciences by Edmond Goblot. Isis, 4(3):565–566, April 1922. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223863.

Sarton:1922:BRBn

[238] George Sarton. Book review: Wunder, Wundergeburt und Wundergestalt in Einblattdrucken des funfzehnten bis achtzehnten Jahrhunderts by Hollander, Eugen. Isis, 4(3):566–567, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223864.

Sarton:1922:BRBo

[239] George Sarton. Book review: Spiritualism Among Civilised and Savage Races by Edward Lawrence. Isis, 4(3):567–568, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223865.

Sarton:1922:BRBp

[240] G. Sarton. Book review: Fishing from the Earliest Times by William Radcliffe. Isis, 4(3):568–571, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223866.

Sarton:1922:TCB

[241] George Sarton and Elizabeth Gilpatrick. Twelfth critical bibliography of the history and philosophy of science and of the history of civilization (to March 1922). Isis, 4(3):572–647, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223867.

Anonymous:1922:VI

[242] Anonymous. Volume information. Isis, 4(3):648–654, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223821.

Anonymous:1922:BM

[243] Anonymous. Back matter. Isis, 4(3):??, April 1922. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223868.

Anonymous:1923:FMa

[244] Anonymous. Front matter. Isis, 5(1):1–3, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223594.

Sarton:1923:KC

[245] George Sarton. Knowledge and charity. Isis, 5(1):5–19, ???? 1923. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/223595.

Karpinski:1923:MMP

[246] Louis-C. Karpinski. Michigan mathematical papyrus no. 621. Isis, 5(1): 20–25, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223596.

Ruska:1923:ABQ

[247] Julius Ruska. Al-Biruni als Quelle für das Leben und die Schriften al-Razi’s. (German) [Al-Biruni as a source for the life and the writings of al-Razi’s]. Isis, 5(1):26–50, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223597.

Schoy:1923:AGA

[248] Carl Schoy. Aus der astronomischen Geographie der Araber. (German) [From the astronomical geography of the Arabs]. Isis, 5(1):51–74, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223598.

Wright:1923:NKL

[249] John Kirtland Wright. Notes on the knowledge of latitudes and longitudes in the Middle Ages. Isis, 5(1):75–98, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223599.

Bond:1923:BRW

[250] Joannes David Bond. Quadripartitum Ricardi Walynforde de Sinibus Demonstratis. Isis, 5(1):99–115, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223600.

Cajori:1923:MSE

[251] Florian Cajori. Mathematical signs of equality. Isis, 5(1):116–125, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223601.

Sarton:1923:PBJ

[252] George Sarton and Laura Jourdain. Philip E. B. Jourdain (1879–1919). Isis, 5(1):126–136, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223602.

See errata [367].

Milner:1923:NC

[253] Viscount Milner, Leopold de Saussure, G. S., and Searcher. Notes and correspondence. Isis, 5(1):137–147, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223603.

Sarton:1923:BRBa

[254] Sarton. Book review: Marcus Aurelius by Henry Dwight Sedgwick. Isis, 5(1):148–149, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223604.

Sarton:1923:BRBb

[255] Sarton. Book review: Dani Opera Omnia by Tychonis Brahe; I. L. E. Dreyer; Ioannis Raeder; G. A. Hagemann. Isis, 5(1):149–151, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223605.

Sarton:1923:BRBc

[256] Sarton. Book review: Opere di Evangelista Torricelli by Gino Loria; Giuseppe Vassura. Isis, 5(1):151–154, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223606.

Sarton:1923:BRBd

[257] Sarton. Book review: Oeuvres complètes de Christiaan Huygens. Isis, 5 (1):154–156, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223607.

Boutroux:1923:BRB

[258] Pierre Boutroux. Book review: A History of the Conceptions of Limits and Fluxions in Great Britain, from Newton to Woodhouse by Cajori, Florian. Isis, 5(1):156–157, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223608.

Guinet:1923:BRBa

[259] L. Guinet. Book review: Deodat Dolomieu by A. Lacroix. Isis, 5(1): 158–159, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223609.

Guinet:1923:BRBb

[260] L Guinet. Book review: L’oeuvre scientifique de Laplace by Andoyer, H. Isis, 5(1):159–160, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223610.

Ruska:1923:BRBa

[261] Julius Ruska. Book review: Geschichte der syrischen Literatur mit Ausschluss der christlich-palastinensischen Texte by Anton Baumstark. Isis, 5 (1):160–162, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223611.

S:1923:BRBa

[262] G. S. Book review: The Legacy of Greece by R. W. Livingstone. Isis, 5(1): 162–163, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223612.

Masson-Oursel:1923:BRBa

[263] P. Masson-Oursel. Book review: Hinduism and Buddhism, an Historical Sketch by Charles Eliot. Isis, 5(1):163–164, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223613.

Masson-Oursel:1923:BRBb

[264] P. Masson-Oursel. Book review: Aspects of Ancient Indian Polity by Narendra Nath Law. Isis, 5(1):164–165, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223614.

Ruska:1923:BRBb

[265] Julius Ruska. Book review: Les Penseurs de l’Islam by Baron Carra de Vaux. Isis, 5(1):165–167, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223615.

S:1923:BRBb

[266] G. S. Book review: Legendary Islands of the Atlantic by William H. Babcock. Isis, 5(1):167–170, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223616.

Sarton:1923:BRBe

[267] George Sarton. Book review: Arabic and Latin Anatomical Terminology Chiefly from the Middle Ages by A. Fonahn. Isis, 5(1):170–172, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223617.

Guinet:1923:BRBc

[268] L. Guinet. Book review: Histoire de l’astronomie by E. Doublet. Isis, 5(1): 172–173, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223618.

S:1923:BRBc

[269] G. S. Book review: Pagine di Storia della chimica by Aldo Mieli. Isis, 5(1): 173–174, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223619.

Sarton:1923:BRBf

[270] George Sarton. Book review: An Introduction to the History of History by James T. Shotwell. Isis, 5(1):174–176, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223620.

Ruska:1923:BRBc

[271] J. Ruska. Book review: Der Untergang des Abendlandes. Umrisse einer Morphologie der Weltgeschichte by Oswald Spengler. Isis, 5(1):176–181, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/223621.

Karpinski:1923:BRB

[272] Louis C. Karpinski. Book review: Storia della Geometria descrittiva dalle origini sino ai giorni nostri by Gino Loria. Isis, 5(1):181–182, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223622.

Wieleitner:1923:BRBa

[273] H. Wieleitner. Book review: Geschichte der Elementar-Mathematik in systematischer Darstellung mit besonderer Berucksichtigung der Fachworter by Tropfke, Johannes. Isis, 5(1):182–186, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223623.

Wieleitner:1923:BRBb

[274] H. Wieleitner. Book review: Ostwalds Klassiker der exakten Wissenschaften. Nr 194. Die erste Integralrechnung by Johann Bernoullis; G. Kowalewski; Ostwalds Klassiker der exakten Wissenschaften. Nr 197. Erster Entwurf eines Versuchs über die Ergebnisse des Zusammentreffens eines Kegels mit einer Ebene by G. Desargues; M. Zacharias; Ostwalds Klassiker der exakten Wissenschaften. Nr 198. J. Keplers Grundlagen der geometrischen Optik by F. Plehn; M. von Rohr; Ostwalds Klassiker der exakten Wissenschaften. Nr 201. Über Spiralen by Archimedes; A. Czwalina. Isis, 5(1):186–187, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223624.

S:1923:BRBd

[275] G. S. Book review: Kurzer Handbuch der Geschichte der Medizin by Karl Sudhoff; J. L. Pagels; Geschichte der Medizin im Überblick mit 215 Abbildungen by Th. Meyer-Steineg; Karl Sudhoff. Isis, 5(1):188, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223625.

Guinet:1923:BRBd

[276] L. Guinet. Book review: L’avenir de la médecine by Mackenzie, James; F. Francon. Isis, 5(1):189–190, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223626.

Guinet:1923:BRBe

[277] L. Guinet. Book review: Vaccins et serums by Apert. Isis, 5(1):190– 191, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223627.

Ruska:1923:BRBd

[278] Julius Ruska. Book review: Das Reisetagebuch eines Philosophen by Hermann Keyserling. Isis, 5(1):191–193, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223628.

Guinet:1923:BRBf

[279] L. Guinet. Book review: Électricité atmospherique by B. Chauveau. Isis, 5(1):193–194, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223629.

S:1923:BRBe

[280] G. S. Book review: The Discovery of the Circulation of the Blood by Charles Singer. Isis, 5(1):194–196, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223630.

Guinet:1923:BRBg

[281] L. Guinet. Book review: Introduction á la psychanalyse by Freud Siegmund; S. Jankelevitch. Isis, 5(1):196–198, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223631.

S:1923:BRBf

[282] G. S. Book review: Aus der Werkstatt grosser Forscher by Friedrich Dannemann. Isis, 5(1):198–200, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223632.

Guinet:1923:BRBh

[283] L. Guinet. Book review: La science et l’esprit positif chez les Penseurs contemporains by Boll, Marcel. Isis, 5(1):200–201, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223633.

Sarton:1923:TCB

[284] George Sarton. Thirteenth critical bibliography of the history and philosophy of science and of the history of civilization. (to October 1922). Isis, 5(1):202–244+262–313, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/ 223634. See errata [368].

Brasch:1923:MBC

[285] Frederick E. Brasch. Materials for the biography of contemporary scientists (chiefly obituary notices). Isis, 5(1):244–262, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223635.

Gilpatrick:1923:AIT

[286] Elizabeth Gilpatrick. Authors’ index to the thirteenth bibliography. Isis, 5(1):314–324, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223636.

Anonymous:1923:BMa

[287] Anonymous. Back matter. Isis, 5(1):??, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223637.

Anonymous:1923:VI

[288] Anonymous. Volume information. Isis, 5(1):??, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223593.

Anonymous:1923:FMb

[289] Anonymous. Front matter. Isis, 5(2):325–328, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/223731.

Shephenson:1923:CSA

[290] J. Shephenson. The classification of the sciences according to nasiruddin tusi. Isis, 5(2):329–338, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223732.

Bond:1923:RWB

[291] John David Bond. Richard Wallingford’s Quadripartitum. Isis, 5(2):339– 363, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223733.

Schoy:1923:BAT

[292] Carl Schoy. Beitrage zur arabischen trigonometrie. Isis, 5(2):364–399, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223734.

Bopp:1923:WSE

[293] K. Bopp. Ein wichtiger Satz über die Ellipse des Fagnano und seine Erganzung. (German) [An important theorem on the ellipse of the Fagnano and its complement]. Isis, 5(2):400–402, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223735.

Guinet:1923:RGI

[294] L. Guinet. Remy et Gehin, les inventeurs français de la pisciculture. (French) [Remy and Gehin, the French inventors of fish farming]. Isis, 5(2):403–405, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223736.

Bopp:1923:EG

[295] K. Bopp. Entwicklungslinien in der geometrie. Isis, 5(2):406–408, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223737.

Ruska:1923:HS

[296] Julius Ruska. Heinrich suter. Isis, 5(2):409–417, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223738.

Sarton:1923:NC

[297] George Sarton. Notes and correspondence. Isis, 5(2):418–421, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223739.

Guinet:1923:BRBi

[298] L. Guinet. Book review: Étude bibliographique de la geographie d’Eratosthene by Thalamas, A.; La Géographie d’Eratosthene. Isis, 5(2): 422–426, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223740.

S:1923:BRBg

[299] G. S. Book review: Descartes savant by Gaston Milhaud. Isis, 5(2): 426–429, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223741.

Cajori:1923:BRB

[300] Florian Cajori. Book review: James Stirling. A Sketch of His Life and Works, along with His Scientific Correspondence by Charles Tweedie; James Stirling. Isis, 5(2):429–432, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223742.

Guinet:1923:BRBj

[301] L. Guinet. Book review: Les mines de charbon en France au XVIIIe siècle (1744–1791) by Rouff, Marcel. Isis, 5(2):432–437, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223743.

Guinet:1923:BRBk

[302] L. Guinet. Book review: La pensée d’Ernst Mach by Bouvier, Robert. Isis, 5(2):437–440, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223744.

Bloch:1923:BRBa

[303] Ernst Bloch. Book review: Aus meinem Leben by Fischer, Emil. Isis, 5(2): 440–441, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223745.

S:1923:BRBh

[304] G. S. Book review: Robert Fulton and the Submarine by William Barclay Parsons. Isis, 5(2):441–444, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223746.

Laufer:1923:BRBa

[305] B. Laufer. Book review: Dogs of China and Japan in Nature and Art by V. W. F. Collier. Isis, 5(2):444–446, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223747.

Chao:1923:BRB

[306] Yuen Ren Chao. Book review: Encyclopedia of Chinese Biography by Chung Kuo Jen Ming Ta Tz’u Tien. Isis, 5(2):446–447, ???? 1923. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/223748.

Sarton:1923:BRBg

[307] George Sarton. Book review: The Edwin Smith Papyrus by James Henry Breasted. Isis, 5(2):447–449, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223749.

S:1923:BRBi

[308] G. S. Book review: Manuale di bibliografia musulmana by Giuseppe Gabrieli. Isis, 5(2):449–450, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223750.

S:1923:BRBj

[309] G. S. Book review: Über Medizin und Krankenpflege im Mittelalter in Schweizerischen Landen by Conrad Brunner. Isis, 5(2):450–451, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223751.

Ruska:1923:BRBe

[310] Julius Ruska. Book review: Die Alchemie des Geber by Ernst Darmstaedter. Isis, 5(2):451–455, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223752.

Guinet:1923:BRBl

[311] L. Guinet. Book review: Le determinisme et l’adaptation morphologique en biologie animale by Raoul Anthony. Isis, 5(2):455–456, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223753.

Laufer:1923:BRBb

[312] B. Laufer. Book review: Unsere Heilpflanzen, ihre Geschichte und ihre Stellung in der Volkskunde by Heinrich Marzell. Isis, 5(2):456–457, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223754.

Sarton:1923:BRBh

[313] G. Sarton. Book review: The Old English Herbals by Eleanour Sinclair Rohde. Isis, 5(2):457–461, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223755.

Sarton:1923:BRBi

[314] G. Sarton. Book review: An Outline of the History of Phytopathology by Whetzel, Herbert Hice. Isis, 5(2):461–464, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223756.

Guinet:1923:BRBm

[315] L. Guinet. Book review: Les edifices physico-chimiques by Pierre Jean Achalme. Isis, 5(2):464–465, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223757.

Bloch:1923:BRBb

[316] Ernst Bloch. Book review: Die geschichtliche Entwicklung der Chemie by Eduard Faerber. Isis, 5(2):465–466, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223758.

Lenoir:1923:BRB

[317] Raymond Lenoir. Book review: La Mentalite primitive by Lucien Levy-Bruhl. Isis, 5(2):467–468, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223759.

Sarton:1923:BRBj

[318] George Sarton. Book review: Le Jura by Emm. de Margerie. Isis, 5(2): 468–469, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223760.

Sheffer:1923:BRB

[319] Henry M. Sheffer. Book review: Per la storia della logica: I principii e l’ordine della scienza nel concetto dei pensatori matematici by Federigo Enriques. Isis, 5(2):469–470, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223761.

Wieleitner:1923:BRBc

[320] H. Wieleitner. Book review: Kritik der mathematischen Vernunft by J. E. Gerlach. Isis, 5(2):470–471, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223762.

Guinet:1923:BRBn

[321] L. Guinet. Book review: Histoire de la philosophie allemande by Emile Brehier. Isis, 5(2):471–473, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223763.

Metzger:1923:BRBa

[322] Helene Metzger. Book review: Les réactions intellectuelles élémentaires by Andre Cresson. Isis, 5(2):473–474, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223764.

Guinet:1923:BRBo

[323] L. Guinet. Book review: Les problèmes généraux de l’héredité psychologique by Georges Poyer. Isis, 5(2):474–475, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223765.

S:1923:BRBk

[324] G. S. Book review: Veteris vestigia flammae by Guglielmo Bilancioni. Isis, 5(2):475–477, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223766.

S:1923:BRBl

[325] G. S. Book review: An Introduction to the History of Science by Walter Libby; The History of Medicine in Its Salient Features. Isis, 5(2):478– 479, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223767.

Metzger:1923:BRBb

[326] H. Metzger. Book review: L’expérience humaine et la causalité physique by Leon Brunschvicg. Isis, 5(2):479–483, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223768.

Guinet:1923:BRBp

[327] L. Guinet. Book review: Conciliation du veritable déterminisme mécanique avec l’existence de la vie et de la liberté morale by Joseph Boussinesq. Isis, 5(2):483–484, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223769.

Guinet:1923:BRBq

[328] L. Guinet. Book review: La théorie de la physique chez les physiciens contemporains by Abel Rey. Isis, 5(2):484–485, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223770.

Guinet:1923:BRBr

[329] L. Guinet. Book review: Esquisse d’une histoire de la technique by Arthur Vierendeel. Isis, 5(2):486–487, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223771.

vKlinckowstroem:1923:BRB

[330] Carl v. Klinckowstroem. Book review: Die Technik, ihr Wesen und ihre Beziehungen zu anderen Lebensgebieten by Robert Weyrauch. Isis, 5(2): 487–488, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223772.

Anonymous:1923:BMb

[331] Anonymous. Back matter. Isis, 5(2):??, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223773.

Sarton:1923:FCB

[332] George Sarton. Fourteenth critical bibliography of the history and philosophy of science and of the history of civilization. (to May 1923). Isis, 5(3): 489–569, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224041.

Gilpatrick:1923:AIF

[333] Elizabeth Gilpatrick. Authors’ index to the fourteenth bibliography. Isis, 5(3):570–576, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224042.

Anonymous:1923:BMc

[334] Anonymous. Back matter. Isis, 5(3):??, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224043.

Anonymous:1923:FMc

[335] Anonymous. Front matter. Isis, 5(3):??, ???? 1923. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224040.

Anonymous:1924:FMa

[336] Anonymous. Front matter. Isis, 6(1):1–3, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223967.

Anonymous:1924:PVV

[337] Anonymous. Preface to volume VI. Isis, 6(1):4–8, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223968.

Sarton:1924:NH

[338] George Sarton. The new humanism. Isis, 6(1):9–42, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223969.

S:1924:NC

[339] G. S. and L. G. Notes and correspondence. Isis, 6(1):43–46, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223970.

S:1924:BRBa

[340] G. S. Book review: Hippocrates by William Henry Samuel Jones. Isis, 6 (1):47–48, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223971.

Ruska:1924:BRBa

[341] Julius Ruska. Book review: Plato und die sogenannten Pythagoreer by Erich Frank. Isis, 6(1):48–52, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223972.

Sarton:1924:BRBa

[342] George Sarton. Book review: L’art religieux du XIIe siècle en France by Emile Male. Isis, 6(1):52–56, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223973.

Sarton:1924:BRBb

[343] George Sarton. Book review: Paracelsus. — Saemtliche Werke by Theophrast von Hohenheim; Paracelsus. Isis, 6(1):56–57, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223974.

Sarton:1924:BRBc

[344] George Sarton. Book review: Les doctrines chimiques en France du debut du XVIIe á la fin du XVIIIe siècle by Helene Metzger. Isis, 6(1):57– 64, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223975.

Sarton:1924:BRBd

[345] George Sarton. Book review: The Life of Horace Benedict de Saussure by Douglas W. Freshfield; Henry F. Montagnier. Isis, 6(1):64–71, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223976.

S:1924:BRBb

[346] G. S. Book review: Fifty Years of Electricity by John Ambrose Fleming. Isis, 6(1):71–73, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223977.

S:1924:BRBc

[347] G. S. Book review: Catalogue of the Indian Collections in the Museum of Fine Arts, Boston by Ananda K. Coomaraswamy. Isis, 6(1):73–74, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223978.

Sarton:1924:BRBe

[348] George Sarton. Book review: A History of Magic and Experimental Science during the First Thirteen Centuries of Our Era by Lynn Thorndike. Isis, 6(1):74–89, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223979.

Ruska:1924:BRBb

[349] Julius Ruska. Book review: Die Entwicklung der chemischen Technik bis zu den Anfangen der Grossindustrie by Fester, Gustav. Isis, 6(1): 89–90, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223980.

S:1924:BRBd

[350] G. S. Book review: Déscription géometrique detaillée des Alpes françaises by Paul Helbronner. Isis, 6(1):90–91, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223981.

S:1924:BRBe

[351] G. S. Book review: Founders of Oceanography and Their Work by William Abbott Herdman. Isis, 6(1):91–95, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223982.

Guinet:1924:BRBa

[352] L. Guinet. Book review: A la Gloire de la Terre, souvenirs d’un geologue by Termier, Pierre. Isis, 6(1):95–96, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223983.

Child:1924:BRB

[353] J. M. Child. Book review: History of the Theory of Numbers, Vol. III Quadratic and Higher Forms by Leonard Eugen Dickson. Isis, 6(1):96– 98, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223984.

S:1924:BRBf

[354] G. S. Book review: Nosography in Modern Internal Medicine by Knud Faber. Isis, 6(1):98–99, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223985.

S:1924:BRBg

[355] G. S. Book review: La philosophie comparée by Paul Masson-Oursel. Isis, 6(1):99–104, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223986.

Sarton:1924:BRBf

[356] George Sarton. Book review: Catalogue of the Wheeler Gift of Books, Pamphlets and Periodicals in the Library of the American Institute of Electrical Engineers by William D. Weaver; Bibliographical History of Electricity and Magnetism Chronologically Arranged by Paul Fleury Mottelay. Isis, 6(1):104–107, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223987.

Sarton:1924:BRBg

[357] George Sarton. Book review: Paleopathology by Roy L. Moodie. Isis, 6(1): 107–112, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223988.

Lenoir:1924:BRB

[358] Raymond Lenoir. Book review: Behaviorism and Psychology by A. A. Roback. Isis, 6(1):112–115, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223989.

S:1924:BRBh

[359] G. S. Book review: Die Naturwisschenschaften in ihrer Entwicklung und in ihrem Zusammenhange by Friedrich Dannemann. Isis, 6(1):115–116, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/223990.

S:1924:BRBi

[360] G. S. Book review: Beitrage zur Geschichte der Naturwissenschaften und Technik by Edmund O. von Lippmann. Isis, 6(1):116–119, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223991.

S:1924:BRBj

[361] G. S. Book review: Science and Civilization by Francis Sydney Marvin. Isis, 6(1):119–121, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223992.

G:1924:BRBa

[362] L. G. Book review: La Magie by Maxwell, Joseph. Isis, 6(1):121–122, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223993.

Sarton:1924:BRBh

[363] George Sarton. Book review: The Marine Chronometer by Rupert T. Gould. Isis, 6(1):122–129, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223994.

vKlinckowstroem:1924:BRBa

[364] Carl v. Klinckowstroem. Book review: Die Technik des Altertums by Neuburger, Albert; Geschichte der Technik by Neudeck, G. Isis, 6(1): 129–131, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223995.

vKlinckowstroem:1924:BRBb

[365] Carl v. Klinckowstroem. Book review: Katalog der Bibliothek des Reichspatentamts. Isis, 6(1):131, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223996.

Anonymous:1924:EKC

[366] Anonymous. Errata: Knowledge and charity. Isis, 6(1):132, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223997.

Anonymous:1924:EPB

[367] Anonymous. Errata: Philip E. B. Jourdain (1879–1919). Isis, 6(1):132, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/223998. See [252].

Anonymous:1924:ETC

[368] Anonymous. Errata: Thirteenth critical bibliography of the history and philosophy of science and of the history of civilization. (to October 1922). Isis, 6(1):132, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223999. See

[284].

Anonymous:1924:BMa

[369] Anonymous. Back matter. Isis, 6(1):??, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224000.

Anonymous:1924:FMb

[370] Anonymous. Front matter. Isis, 6(2):133, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223906.

Sarton:1924:FCB

[371] George Sarton. Fifteenth critical bibliography of the history and philosophy of science and of the history of civilization. (to December 1923). Isis, 6(2):135–251, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223907.

Gilpatrick:1924:AIF

[372] Elizabeth Gilpatrick. Authors’ index to the fifteenth bibliography. Isis, 6 (2):253–264, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223908.

Anonymous:1924:BMb

[373] Anonymous. Back matter. Isis, 6(2):??, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223909.

Anonymous:1924:FMc

[374] Anonymous. Front matter. Isis, 6(3):265–267, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224310.

Gudger:1924:PHN

[375] E. W. Gudger. Pliny’s historia naturalis. the most popular natural history ever published. Isis, 6(3):269–281, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224311.

Gabrieli:1924:HII

[376] Giuseppe Gabrieli. Hunayn ibn ishaq. Isis, 6(3):282–292, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224312.

Holmyard:1924:MAM

[377] E. J. Holmyard. Maslama al-Majriti and the Rutbatu’l-Hakim. Isis, 6(3): 293–305, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224313.

Sarton:1924:WMA

[378] George Sarton. Walton. A Medieval aquiculturist. Isis, 6(3):306–310, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224314.

Smith:1924:FPA

[379] David Eugene Smith. The first printed arithmetic (Treviso, 1478). Isis, 6 (3):311–331, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224315.

Schoy:1924:SSA

[380] Karl Schoy. Sonnenuhren der spatarabischen astronomie. Isis, 6(3):332– 360, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224316.

Thorndike:1924:LHS

[381] Lynn Thorndike. L’Encyclopédie and the history of science. Isis, 6(3): 361–386, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224317.

Marvin:1924:FH

[382] F. S. Marvin. Frederic Harrison. Isis, 6(3):387–390, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224318.

Cajori:1924:EGG

[383] Florian Cajori. Empirical generalizations on the growth of mathematical notations. Isis, 6(3):391–394, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224319.

Dannemann:1924:WE

[384] Friedrich Dannemann. Wissenschaft als einheit. Isis, 6(3):395–399, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224320.

Menghin:1924:GOT

[385] Oswald Menghin. Geschichtliches zur Ölbereitung aus Traubenkernen. (German) [History for the preparation of oil from grape seeds]. Isis, 6(3): 400–403, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224321.

Sarton:1924:NC

[386] George Sarton and L. G. Notes and correspondence. Isis, 6(3):404–412, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224322.

Guinet:1924:BRBb

[387] L. Guinet. Book review: Guillaume Bude (1468–1540) et les origines de l’humanisme français by Jean Plattard. Isis, 6(3):413–415, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224323.

Sarton:1924:BRBi

[388] George Sarton. Book review: Ignatius Loyola. An Attempt at an Impartial Biography by Henry Dwight Sedgwick. Isis, 6(3):415–416, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224324. See errata [596].

Sarton:1924:BRBj

[389] G. Sarton. Book review: Sir Christopher Wren. Scientist, Scholar, and Architect by Lawrence Weaver. Isis, 6(3):416–419, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224325.

Sarton:1924:BRBk

[390] George Sarton. Book review: Hallers Briefe an Johannes Gesner (1728– 1777) by Haller, Albrecht von; Henry E. Sigerist. Isis, 6(3):419–420, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224326.

Sarton:1924:BRBl

[391] George Sarton. Book review: Linnaeus by Benjamin Dayton Jackson; Theodor Magnus Fries. Isis, 6(3):420–423, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224327.

S:1924:BRBk

[392] G. S. Book review: The Life of Sir William White, K. C. B., F. R. S., etc by Frederic Manning. Isis, 6(3):423, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224328.

Guinet:1924:BRBc

[393] L. Guinet. Book review: Pasteur et le transformisme by Blaringhem Louis. Isis, 6(3):424–425, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224329.

Ruska:1924:BRBc

[394] Julius Ruska. Book review: Handbuch der Islam-Literatur by Gustav Pfannmuller. Isis, 6(3):425–427, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224330.

Ruska:1924:BRBd

[395] Julius Ruska. Book review: Die Flora der Juden. II. Iridaceae-Papilionaceae by Immanuel Low. Isis, 6(3):428–429, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224331.

Sarton:1924:BRBm

[396] George Sarton. Book review: Studien und Texte zur fruhmittelalterlichen Rezeptliteratur. (Studien zur Geschichte der Medizin, H. 13) by Henry E. Sigerist. Isis, 6(3):429–430, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224332.

Sarton:1924:BRBn

[397] George Sarton. Book review: Mèmoires scientifiques publiès par J. L. Heiberg, Vol. 5. Sciences exactes au Moyen Age (1887–1921) by Paul Tannery; J. L. Heiberg. Isis, 6(3):431–435, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224333.

Guinet:1924:BRBd

[398] L. Guinet. Book review: Morphologie générale; Membres et ceintures des vertebres tetrapodes; Critique morphologique du transformisme by Vialleton, Louis. Isis, 6(3):435–437, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224334.

Guinet:1924:BRBe

[399] L. Guinet. Book review: Ethnographie sismique et volcanique by F. de Montessus de Ballore. Isis, 6(3):437–439, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224335.

Sheffer:1924:BRBa

[400] Henry M. Sheffer. Book review: Substance and Function and Einstein’s Theory of Relativity by Ernst Cassirer; Marie C. Swabey; William C. Swabey. Isis, 6(3):439–440, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224336.

Sarton:1924:BRBo

[401] George Sarton. Book review: History of Mathematics. Volume 1. General Survey of the History of Elementary Mathematics by David Eugene Smith. Isis, 6(3):440–444, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224337.

Koenig:1924:BRB

[402] F. Otto Koenig, Jr. and F. Dannemann. Book review: Entwicklungsgeschichte der modernen Physik, zugleich eine Übersicht ihrer Tatsachen, Gesetze und Theorien by Felix Auerbach. Isis, 6(3):444–447, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224338.

May:1924:BRBa

[403] R. M. May. Book review: Nutrition: The Chemistry of Life by Lafayette B. Mendel. Isis, 6(3):447–448, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224339.

Sarton:1924:BRBp

[404] George Sarton. Book review: Early Science in Oxford by Robert Theodor Gunther. Isis, 6(3):449–453, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224340.

G:1924:BRBb

[405] L. G. Book review: Les systèmes socialistes by Bourgin, Hubert. Isis, 6(3): 453–454, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224341.

Sarton:1924:BRBq

[406] George Sarton. Book review: Traité théorique et pratique des moteurs à gaz, à essence et à pétrole by Aime Witz. Isis, 6(3):454–456, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224342.

Sarton:1924:BRBr

[407] George Sarton. Book review: A Bibliography of Fishes by Bashford Dean; Charles Rochester Eastman; Eugene Willis Gudger; Arthur Wilbur Henn. Isis, 6(3):456–459, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224343.

Brasch:1924:MBC

[408] Frederick E. Brasch. Materials for the biography of contemporary scientists (chiefly obituary notices). Isis, 6(3):460–474, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224344.

Anonymous:1924:BMc

[409] Anonymous. Back matter. Isis, 6(3):??, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224345.

Anonymous:1924:FMd

[410] Anonymous. Front matter. Isis, 6(4):477–478, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/223526.

Holmyard:1924:CEB

[411] E. J. Holmyard. A critical examination of Berthelot’s work upon Arabic chemistry. Isis, 6(4):479–499, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223528.

Gabrieli:1924:MSA

[412] Giuseppe Gabrieli. Medici e scienziati arabi: Ali ibn ridwan. Isis, 6(4): 500–506, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223529.

Cajori:1924:NOS

[413] Florian Cajori. Note on our sign of equality. Isis, 6(4):507–508, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223530.

Wieleitner:1924:GGE

[414] H. Wieleitner. Zur Geschichte der gebrochenen Exponenten. (German) [On the history of fractional exponents]. Isis, 6(4):509–520, ???? 1924.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223531.

Brasch:1924:NC

[415] Frederick E. Brasch, R. C. Archibald, Julius Ruska, Victor Goldschmidt, and George Sarton. Notes and correspondence. Isis, 6(4):521–546, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223532.

Sarton:1924:BRBs

[416] George Sarton. Book review: Fasciculus Medicinae, 1491 by Joannes de Ketham Alemanus; Charles Singer. Isis, 6(4):547–549, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223533.

Sarton:1924:BRBt

[417] G. Sarton. Book review: The Life of Sir William Crookes, O. M., F. R. S. by E. E. Fournier d’Albe. Isis, 6(4):549–553, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223534.

Sarton:1924:BRBu

[418] George Sarton. Book review: The Rhind Mathematical Papyrus by T. Eric Peet. Isis, 6(4):553–557, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223535.

Guinet:1924:BRBf

[419] L. Guinet. Book review: La pensée grecque et les origines de l’esprit scientifique. Vol. XIII de l’Ëvolution de l’Humanité by Robin, Léon; Henri Berr. Isis, 6(4):557–559, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223536.

Ruska:1924:BRBe

[420] Julius Ruska. Book review: Islamstudien. Vom Werden und Wesen der islamischen Welt by C. H. Becker. Isis, 6(4):559–561, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223537.

Macdonald:1924:BRBa

[421] D. B. Macdonald. Book review: La Syrie a l’epoque des Mamelouks d’apres les Auteurs Arabes by Maurice Gaudefroy-Demombynes. Isis, 6(4):561– 563, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223538.

Macdonald:1924:BRBb

[422] D. B. Macdonald. Book review: L’Islam et les Races by P. J. Andre; Paul Geuthner. Isis, 6(4):563–564, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223539.

Guinet:1924:BRBg

[423] L. Guinet. Book review: La mémoire biologique by Rignano, Eugenio. Isis, 6(4):564–565, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223540.

Guinet:1924:BRBh

[424] L. Guinet. Book review: Une science nouvelle: La science des vibrations atomiques by Mager, Henri. Isis, 6(4):565–566, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223541.

S:1924:BRBl

[425] G. S. Book review: Bibliographie de la Relativite by Lecat, Maurice; M. Lecat-Pierlot. Isis, 6(4):567–568, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223542.

Guinet:1924:BRBi

[426] L. Guinet. Book review: Historia de la medicina en España by Garcia del Real, Eduardo. Isis, 6(4):568–569, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223543.

Sheffer:1924:BRBb

[427] Henry M. Sheffer. Book review: Durée et simultanéité, à propos de la théorie d’Einstein by Bergson, Henri. Isis, 6(4):570–571, ???? 1924. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/223544.

Guinet:1924:BRBj

[428] L. Guinet. Book review: Comment agnostiquer les aptitudes chez les écoliers by Claparede, Edouard. Isis, 6(4):571–572, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223545.

Sheffer:1924:BRBc

[429] Henry M. Sheffer. Book review: Die Grundlagen der Physik. Synthetische Prinzipien der mathematischen Naturphilosophie by Dingler, Hugo. Isis, 6(4):572–573, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223546.

Sheffer:1924:BRBd

[430] Henry M. Sheffer. Book review: Kant und Einstein. Untersuchungen über das Verhaltnis der modernen Erkenntnistheorie zur Relativitatstheorie by Elsbach, Alfred C. Isis, 6(4):573–574, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223547.

Guinet:1924:BRBk

[431] L. Guinet. Book review: La loi du progres en biologie et en sociologie et la question de l’organisme social by Ferriere, Adolphe. Isis, 6(4):574– 576, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223548.

Sarton:1924:BRBv

[432] George Sarton. Book review: Days of Delusion by Clara Endicott Sears. Isis, 6(4):576–578, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223549.

May:1924:BRBb

[433] R. M. May. Book review: Social Life among the Insects by Wheeler, William Morton; Le monde social des fourmis du globe comparé à celui de l’homme by Forel, Auguste. Isis, 6(4):578–580, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223550.

Anonymous:1924:BMd

[434] Anonymous. Back matter. Isis, 6(4):??, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223551.

Anonymous:1924:FPI

[435] Anonymous. Frontispiece: Portrait of Immanuel Kant to celebrate the second centenary of his birth at Königsberg on the 22nd of April 1724 [after the painting of Dobler (1791)]. Isis, 6(4):??, ???? 1924. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223527.

Gilpatrick:1925:IAV

[436] Elizabeth Gilpatrick. Index of authors to volumes IV, V, and VI. Isis, 6(5): 583–664, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223523.

Anonymous:1925:BMa

[437] Anonymous. Back matter. Isis, 6(5):??, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223525.

Anonymous:1925:BTCa

[438] Anonymous. Brief table of contents of volume VI, (1924). Isis, 6(5): ??, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223524.

Anonymous:1925:FMa

[439] Anonymous. Front matter. Isis, 6(5):??, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223522.

Anonymous:1925:FMb

[440] Anonymous. Front matter. Isis, 7(1):1–3, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224214.

Schoy:1925:DPA

[441] C. Schoy. Drei planimetrische Aufgaben des arabischen Mathematikers Abu’l-Jud Muhammad ibn al-Lith. (German) [Three planimetric tasks of the Arab mathematician Abu’l-Jud Muhammad ibn al-Lith]. Isis, 7(1): 5–8, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224215. See errata

[594].

Gabrieli:1925:FAD

[442] Giuseppe Gabrieli. Fakhr-al-din al-razi. Isis, 7(1):9–13, ???? 1925. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/224216.

Leake:1925:VCD

[443] Chauncey D. Leake. Valerius Cordus and the discovery of ether. Isis, 7(1): 14–24, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224217.

Petronievics:1925:CDA

[444] Branislav Petronievics. Charles Darwin und Alfred Russel Wallace. Isis, 7 (1):25–57, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224218.

Hopkins:1925:MTA

[445] Arthur John Hopkins. A modern theory of alchemy. Isis, 7(1):58–76, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224219.

Turriere:1925:IHO

[446] Emile Turriere. Introduction a l’histoire de l’optique. Isis, 7(1):77–104, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224220.

Schmidt:1925:NC

[447] Georg C. Schmidt, J. Wuerschmidt, Joseph Larmor, George Sarton, and A. L. Notes and correspondence. Isis, 7(1):105–118, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224221.

Sarton:1925:BRBa

[448] George Sarton. Book review: Arabische Alchemisten. II. Ja’far Alsadiq, der sechste Imam by Julius Ruska. Isis, 7(1):119–121, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224222.

Sarton:1925:BRBb

[449] G. Sarton. Book review: Studies in the History of Mediaeval Science by Charles Homer Haskins. Isis, 7(1):121–124, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224223.

Sarton:1925:BRBc

[450] George Sarton. Book review: Kitab al-’ilm al-muktasab fi zira’at adhdhahab by E. J. Holmyard; Abu’l-Qasim Muhammad ibn Ahmad al-’Iraqi. Isis, 7(1):124–128, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224224.

Wieleitner:1925:BRBa

[451] H. Wieleitner. Book review: Geschichte der neusprachlichen wissenschaftlichen Literatur by Leonardo Olschki. Isis, 7(1):128–129, ????

1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224225.

Wright:1925:BRBa

[452] J. K. Wright. Book review: Early Jesuit Travellers in Central Asia: 1603– 1721 by C. Wessels. Isis, 7(1):129–132, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224226.

Sarton:1925:BRBd

[453] George Sarton. Book review: A Long Life’s Work. An Autobiography by Archibald Geikie. Isis, 7(1):132–135, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224227.

Sarton:1925:BRBe

[454] George Sarton. Book review: From Immigrant to Inventor by Michael Pupin. Isis, 7(1):135–138, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224228.

Sarton:1925:BRBf

[455] George Sarton. Book review: The Days of a Man by David Starr Jordan. Isis, 7(1):138–141, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224229.

May:1925:BRB

[456] R. M. May. Book review: Life and Confessions of a Psychologist by G. Stanley Hall. Isis, 7(1):141–143, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224230.

G:1925:BRBa

[457] L. G. Book review: La vie de J. H. Fabré naturaliste, suivie du répertoire général analytique des Souvenirs Entomologiques by G. V. Legros. Isis, 7 (1):143–144, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224231.

Sarton:1925:BRBg

[458] George Sarton. Book review: A Woman’s Quest. The Life of Marie E. Zakrzewska, M. D. by Agnes C. Vietor. Isis, 7(1):144, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224232.

Sarton:1925:BRBh

[459] George Sarton. Book review: Chinese Painting as Reflected in the Thought and Art of Li Lung-Mien, 1070–1106 by Agnes E. Meyer. Isis, 7(1):145– 147, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224233.

Lefrancq:1925:BRB

[460] J. Lefrancq. Book review: Manuel d’archéologie égyptienne. — I. Les éléments de l’architecture by Jequier, G. Isis, 7(1):147–148, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224234.

G:1925:BRBb

[461] L. G. Book review: L’Hygiene chez les anciens Grecs by Panayotatou, Angelique, G. Isis, 7(1):148–149, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224235.

Sarton:1925:BRBi

[462] George Sarton. Book review: The Evolution of Man by G. Elliot Smith. Isis, 7(1):149–152, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224236.

S:1925:BRBa

[463] G. S. Book review: The Depths of the Universe by George Ellery Hale. Isis, 7(1):152–153, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224237.

Wright:1925:BRBb

[464] J. K. Wright. Book review: Geography in France by Emmanuel de Martonne. Isis, 7(1):153–155, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224238.

Wieleitner:1925:BRBb

[465] H. Wieleitner. Book review: Die mathematische Methode by Otto Holder. Isis, 7(1):155–156, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224239.

S:1925:BRBb

[466] G. S. Book review: Lectures de Mècanique by Jouguet, Emile. Isis, 7(1): 156–158, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224240.

Ruska:1925:BRBa

[467] Julius Ruska. Book review: Der Sternhimmel in der Dichtung und Religion der alten Volker und des Christentums by Arthur Drews; Das Markus-Evangelium als Zeugnis gegen die Geschichtlichkeit Jesu; Die Entstehung des Christentums aus dem Gnostizismus. Isis, 7(1):158–162, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224241.

Sarton:1925:BRBj

[468] George Sarton. Book review: The Understanding of Religion by Edwin Tenney Brewster. Isis, 7(1):162–165, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224242.

Klinckowstroem:1925:BRB

[469] Graf Klinckowstroem. Book review: Urwelt, Sage und Menschheit by Edgar Dacque. Isis, 7(1):165–167, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224243.

Ruska:1925:BRBb

[470] Julius Ruska. Book review: Das grosse Ratsel by Maurice Maeterlinck; Lulu von Strauss u. Torney. Isis, 7(1):167–168, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224244.

Anonymous:1925:BMb

[471] Anonymous. Back matter. Isis, 7(1):??, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224245.

Sarton:1925:SCBa

[472] George Sarton. Sixteenth critical bibliography of the history and philosophy of science and of the history of civilization. Isis, 7(2):171–352, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223800.

Gilpatrick:1925:AIS

[473] Elizabeth Gilpatrick. Author’s index to the sixteenth bibliography. Isis, 7(2):353–366, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223801.

Anonymous:1925:BMc

[474] Anonymous. Back matter. Isis, 7(2):??, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223802.

Anonymous:1925:FMc

[475] Anonymous. Front matter. Isis, 7(2):??, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223799.

Anonymous:1925:FMd

[476] Anonymous. Front matter. Isis, 7(3):369–370, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224288.

Brasch:1925:LFM

[477] Frederick E. Brasch. List of the foundation members of the History of Science Society. Isis, 7(3):371–393, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224289.

Tallmadge:1925:TPB

[478] Guy K. Tallmadge. The third part of the De Extractione of Valerius Cordus. Isis, 7(3):394–411, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224290.

Cajori:1925:LMB

[479] Florian Cajori. Leibniz, the master-builder of mathematical notations. Isis, 7(3):412–429, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224291.

Leroux:1925:NL

[480] Lucien Leroux. Nicolas Lemery. Isis, 7(3):430–432, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224292.

Bounoure:1925:VOF

[481] Louis Bounoure. La vie et l’oeuvre de Frédéric Houssay (1860–1920). Isis, 7(3):433–455, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224293.

Gabrieli:1925:FSI

[482] Giuseppe Gabrieli. A. Favaro e gli studi italiani di storia della scienza. (Italian) [A. Favaro and Italian studies in the history of science]. Isis, 7(3): 456–467, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224294.

Stephanides:1925:TA

[483] Michel Stephanides. La terminologie des anciens. Isis, 7(3):468–477, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224295.

Haskins:1925:ASW

[484] Charles H. Haskins. Arabic science in Western Europe. Isis, 7(3):478– 485, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224296.

Wieleitner:1925:FRM

[485] H. Wieleitner. Zur Fruhgeschichte der Raume von mehr als drei Dimensionen. (German) [Toward an early history of space of more than three dimensions]. Isis, 7(3):486–489, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224297.

Wieleitner:1925:GGE

[486] H. Wieleitner. Geschichte der gebrochenen exponenten. Isis, 7(3):490– 491, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224298.

vonLippmann:1925:NC

[487] Edmund O. von Lippmann and E. J. Holmyard. Notes and correspondence. Isis, 7(3):492–493, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224299.

Sarton:1925:BRBk

[488] George Sarton. Book review: The Geographical Lore of the Time of the Crusades by John Kirtland Wright. Isis, 7(3):495–498, ???? 1925. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/224300.

Sarton:1925:BRBl

[489] George Sarton. Book review: The First One Hundred Years of American Geology by George P. Merrill. Isis, 7(3):498–501, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224301.

Bloch:1925:BRB

[490] Ernst Bloch. Book review: Gregor Johann Mendel. Leben, Werk und Wirkung by Hugo Iltis. Isis, 7(3):501–507, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224302.

Sarton:1925:BRBm

[491] George Sarton. Book review: Catalogue des manuscrits alchimiques grecs. I. Les Parisini by Henri Lebegue; Marie Delcourt; Catalogue des manuscrits alchimiques grecs. III. Les manuscrits des Îles britanniques by Dorothea Waley Singer; Annie Anderson; William J. Anderson; Otto Lagercrantz. Isis, 7(3):507–511, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224303.

Davis:1925:BRB

[492] Tenney L. Davis. Book review: The Story of Early Chemistry by John Maxson Stillman. Isis, 7(3):511–514, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224304.

Sarton:1925:BRBn

[493] George Sarton. Book review: Histoire des sciences en France by Gabriel Hanotaux. Isis, 7(3):514–516, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224305.

Metzger:1925:BRB

[494] Helene Metzger. Book review: La déduction relativiste by Emile Meyerson. Isis, 7(3):517–520, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224306.

Sarton:1925:BRBo

[495] George Sarton. Book review: Les rois thaumaturges by Marc Bloch. Isis, 7(3):520–521, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224307.

Sarton:1925:BRBp

[496] George Sarton. Book review: The History of aeronautics in Great Britain by Hodgson, J. E. Isis, 7(3):521–528, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224308.

Anonymous:1925:BMd

[497] Anonymous. Back matter. Isis, 7(3):??, ???? 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224309.

Sarton:1925:SCBb

[498] George Sarton. Seventeenth critical bibliography of the history and philosophy of science and of the history of civilization (to March 1925, with 4 facsimiles on 2 pl.). Isis, 7(4):531–606, December 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223775.

Stewart:1925:AIS

[499] Elizabeth Gilpatrick Stewart. Authors’ index to the seventeenth bibliography. Isis, 7(4):607–614, December 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223776.

Anonymous:1925:BMe

[500] Anonymous. Back matter. Isis, 7(4):??, December 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223778.

Anonymous:1925:BTCb

[501] Anonymous. Brief table of contents of volume VII, (1925). Isis, 7(4): ??, December 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223777.

Anonymous:1925:FMe

[502] Anonymous. Front matter. Isis, 7(4):??, December 1925. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223774.

Anonymous:1926:FMa

[503] Anonymous. Front matter. Isis, 8(1):3–6, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223668.

Sarton:1926:PVV

[504] George Sarton. Preface to Volume VIII (being the third volume published under the auspices of the History of Science Society). Isis, 8(1): 7–11, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223670.

Hovgaard:1926:APA

[505] William Hovgaard. The arsenal in Piraeus and the ancient building rules. Isis, 8(1):12–20, February 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223671.

Schoy:1926:GAS

[506] C. Schoy. Graeco–Arabische studien. (German) [Greek–Arab studies]. Isis, 8(1):21–40, February 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223672.

Smith:1926:FGC

[507] David Eugene Smith. The first great commercial arithmetic. Isis, 8(1):41– 49, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223673.

Smith:1926:NLR

[508] David Eugene Smith and Frances Marguerite Clarke. New light on Robert recorde. Isis, 8(1):50–70, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223674.

Davis:1926:FES

[509] Tenney L. Davis. The first edition of the sceptical chymist. Isis, 8(1):71– 76, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223675.

Guinet:1926:EL

[510] L. Guinet. Émile Littré (1801–1881). Isis, 8(1):77–102, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223676.

VanHee:1926:BJC

[511] Pere Louis Van Hee, Frances Marguerite Clarke, and Anna Elizabeth Houghtaling. The Ch’ou-Jen Chuan of Yüan Yüan. Isis, 8(1):103–118, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/223677.

Bloch:1926:CAG

[512] Ernst Bloch. Das chemische Affinitätsproblem, geschichtlich betrachtet. (German) [The chemical affinity problem, considered historically]. Isis, 8 (1):119–157, February 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223678.

Brasch:1926:NC

[513] F. Brasch and George Sarton. Notes and correspondence. Isis, 8(1):158– 164, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223679.

Petit:1926:BRB

[514] Raymond Petit. Book review: Isis et Osiris by Plutar que; Mario Meunier. Isis, 8(1):165–167, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223680.

Guinet:1926:BRBa

[515] L. Guinet. Book review: Le mystère de Jesus by P. L. Couchoud; Jesus de Nazareth, mythe ou histoire by Maurice Goguel; Jesus dans l’histoire et dans le mystère by L. de Grandmaison. Isis, 8(1):167–168, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223681.

Wright:1926:BRBa

[516] J. K. Wright. Book review: La carte de Christophe Colomb: The Map of Christopher Columbus by Charles de la Ronciere; The “Columbian” World-Map in the Bibliothèque Nationale by Edward Heawood; Review in Geographical Review by G. E. Nunn; Christopher Columbus. Isis, 8 (1):168–173, February 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223682.

Sarton:1926:BRBa

[517] George Sarton. Book review: The Geometry of Rene Descartes by Descartes Rene; David Eugene Smith; Marcia L. Latham. Isis, 8(1):173– 175, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223683.

Metzger:1926:BRB

[518] Helene Metzger. Book review: Le Genie de Pascal by Leon Brunschvicg. Isis, 8(1):175–176, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223684.

Campbell:1926:BRB

[519] Norman R. Campbell. Book review: The Life of John William Strutt by Robert John Strutt. Isis, 8(1):177–181, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223685.

Sarton:1926:BRBb

[520] George Sarton. Book review: The Life, Letters and Labours of Francis Galton by Karl Pearson; Francis Galton. Isis, 8(1):181–188, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223686.

May:1926:BRBa

[521] Raoul M. May. Book review: Impressions of Great Naturalists by Henry Fairfield Osborn. Isis, 8(1):188–189, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223687.

Bloch:1926:BRB

[522] Ernst Bloch. Book review: Paul Ehrlich als Mensch und Arbeiter by Martha Marquardt. Isis, 8(1):189–191, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223688.

G:1926:BRBa

[523] L. G. Book review: La vie des Abeilles et des Guepes by Charles Ferton; Etienne Rabaud; Francois Picard. Isis, 8(1):191–192, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223689.

Ruska:1926:BRBa

[524] Julius Ruska. Book review: Drogen und Drogenhandel im Altertum by Alfred Schmidt. Isis, 8(1):192–193, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223690.

Ruska:1926:BRBb

[525] Julius Ruska. Book review: Zeitrechnung und Weltordnung in ihren übereinstimmenden Grundlagen bei den Indern, Iraniern, Hellenen, Italikern, Germanen, Kelten, Litauern, Slawen by Wolfgang Schultz. Isis, 8(1):193–195, February 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223691.

Ruska:1926:BRBc

[526] Julius Ruska. Book review: Babylonien und Assyrien by Bruno Meissner. Isis, 8(1):195–198, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223692.

Meyerhof:1926:BRB

[527] Max Meyerhof. Book review: Ancient Egyptian Medicine by Hasan Kamal. Isis, 8(1):198–200, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223693.

Sarton:1926:BRBc

[528] George Sarton. Book review: A Greek-English Lexicon by Henry George Liddell; Robert Scott; Henry Stuart Jones; Roderick McKenzie. Isis, 8 (1):200–202, February 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223694.

Ruska:1926:BRBd

[529] Julius Ruska. Book review: Altai-Iran und Volkerwanderung. Ziergeschichtliche Untersuchungen über den Eintritt der Wander — und Nordvolker in die Treibhauser geistigen Lebens by Josef Strzygowski. Isis, 8(1):203–207, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223695.

Ruska:1926:BRBe

[530] Julius Ruska. Book review: Urmensch und Seele in der iranischen Überlieferung. Ein Beitrag zur Religionsgeschichte des Hellenismus by O. G. von Wesendonk. Isis, 8(1):207–209, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223696.

Macdonald:1926:BRB

[531] D. B. Macdonald. Book review: Le Pelerinage á la Mekke; Étude d’histoire religieuse. Annales du Musée Guimet by Maurice Gaudefroy-Demonbynes. Isis, 8(1):209–210, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223697.

Ruska:1926:BRBf

[532] Julius Ruska. Book review: Die Flora der Juden. Band III, Pedaliaceae-Zygophyllaceae by Immanuel Low. Isis, 8(1):210–211, February 1926. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/223698.

W:1926:BRB

[533] J. K. W. Book review: Il sistema cosmografico e i Padri della Chiesa by Salvatore Ciarcia. Isis, 8(1):212–213, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223699.

May:1926:BRBb

[534] Raoul M. May. Book review: General Cytology by E. V. Cowdry; The Cell in Development and Heredity by Edmund B. Wilson. Isis, 8(1):213– 215, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223700.

Follin:1926:BRB

[535] Maynard D. Follin. Book review: Elephants and Ethnologists by G. G. Elliot Smith. Isis, 8(1):215–217, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223701.

G:1926:BRBb

[536] L. G. Book review: La Geologie sismologique: Les tremblements de terre by Fernand de Montessus de Ballore. Isis, 8(1):217–219, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223702.

Guinet:1926:BRBb

[537] L. Guinet. Book review: La scolastique et le thomisme by Louis Rougier. Isis, 8(1):219–221, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223703.

Sarton:1926:BRBd

[538] George Sarton. Book review: History of Mathematics. Volume II. Special Topics of Elementary Mathematics by David Eugene Smith. Isis, 8(1): 221–225, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223704.

Sheffer:1926:BRB

[539] Henry M. Sheffer. Book review: Principia Mathematica by Whitehead, Alfred North; Russell, Bertrand. Isis, 8(1):226–231, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223705.

Cajori:1926:BRB

[540] Florian Cajori. Book review: The History of Arithmetic by Louis Charles Karpinski. Isis, 8(1):231–232, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223706.

G:1926:BRBc

[541] L. G. Book review: Les cosmogonies modernes et la théorie de la connaissance by Busco, Pierre. Isis, 8(1):233–234, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223707.

G:1926:BRBd

[542] L. G. Book review: Le Rameau d’or by James George Frazer; Lady Frazer. Isis, 8(1):234–236, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223708.

Anonymous:1926:BMa

[543] Anonymous. Back matter. Isis, 8(1):??, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223709.

Anonymous:1926:FFG

[544] Anonymous. Frontispiece. Francis Galton in 1903 (Ae. 81). from a painting by C. W. Furse, A. R. A. (courtesy of Methuen and Co.). Isis, 8(1): ??, February 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223669.

Anonymous:1926:FMb

[545] Anonymous. Front matter. Isis, 8(2):237–239, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223638.

Sarton:1926:BUA

[546] George Sarton. A brief but urgent appeal to the friends of Isis and of the H. S. S. Isis, 8(2):240, May 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223639.

Sarton:1926:TIF

[547] George Sarton. Sur la tolèrance intellectuelle. (French) [On intellectuel tolerance]. Isis, 8(2):241–253, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223640.

Schoy:1926:BEG

[548] C. Schoy. Behandlung einiger geometrischen Fragepunkte durch muslimische Mathematiker. (German) [Treatment of some geometric questions by Muslim mathematicians]. Isis, 8(2):254–263, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223641.

Thorndike:1926:MTC

[549] Lynn Thorndike. The manuscript text of the cyrurgia of Leonard of Bertipaglia. Isis, 8(2):264–284, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223642.

Thorndike:1926:ATB

[550] Lynn Thorndike. Another treatise by Barnabas de Riatinis of Reggio. Isis, 8(2):285–286, May 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223643.

Zachaire:1926:ADZ

[551] Denis Zachaire and Tenney L. Davis. The autobiography of Denis Zachaire: An account of an Alchemist’s life in the Sixteenth Century. Isis, 8(2):287–299, May 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223644.

Bevenot:1926:WP

[552] Hugh G. Bevenot. The Weingarten planetarium. Isis, 8(2):300–312, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223645.

Lorey:1926:GE

[553] W. Lorey. Gustav Eneström. Isis, 8(2):313–320, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223646.

Cajori:1926:WWR

[554] Florian Cajori. Walter William Rouse Ball. Isis, 8(2):321–324, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223647.

Cajori:1926:NCF

[555] Florian Cajori. A notable case of finger=reckoning in America. Isis, 8(2): 325–327, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223648.

dIrsay:1926:NOE

[556] Stephen d’Irsay. Notes to the origin of the expression: ≪Atra Mors≫. Isis, 8(2):328–332, May 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223649.

Anonymous:1926:MIH

[557] Anonymous. Medallic illustrations of the history of science. Isis, 8(2): 333–335, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223650.

Thorndike:1926:NC

[558] Lynn Thorndike, Comte Klinckowstroem, K. W. D., Abel Rey, Florian Znaniecki, and Walter Libby. Notes and correspondence. Isis, 8(2):336– 342, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223651.

Sarton:1926:BRBe

[559] George Sarton. Book review: Hermetica by Walter Scott. Isis, 8(2): 343–346, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223652.

Sarton:1926:BRBf

[560] George Sarton. Book review: La parure des cavaliers et l’insigne des preux by Ibn Hudhail,’Ali ibn ’Abd al-Rahman al-Andalusi; M. Nehlil; Louis Mercier. Isis, 8(2):346–349, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223653.

Sarton:1926:BRBg

[561] George Sarton. Book review: The Fasciculo di Medicina, Venice, 1493 by Joannes de Ketham. Isis, 8(2):350, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223654.

Sarton:1926:BRBh

[562] George Sarton. Book review: The Earliest Printed Literature on Syphilis, being Ten Tractates from the Years 1495–1498 by Karl Sudhoff; Charles Singer; Henry E. Sigerist. Isis, 8(2):351–354, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223655.

Sarton:1926:BRBi

[563] George Sarton. Book review: Le grant Kalendrier et compost des Bergiers avecq leur Astrologie et plusieurs aultres choses by Payot. Isis, 8(2):354– 355, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223656.

Wieleitner:1926:BRBa

[564] H. Wieleitner. Book review: Kant als Naturforscher by Erich Adickes. Isis, 8(2):355–358, May 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223657.

Sarton:1926:BRBj

[565] George Sarton. Book review: The Life of Sir William Osler (1849–1919) by Harvey Cushing. Isis, 8(2):358–361, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223658.

Sarton:1926:BRBk

[566] George Sarton. Book review: The Invention of Printing in China and Its Spread Westwards by Thomas Francis Carter. Isis, 8(2):361–373, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223659.

Sarton:1926:BRBl

[567] George Sarton. Book review: The World-Conception of the Chinese. Their Astronomical, Cosmological and Physico-Philosophical Speculations by Alfred Forke. Isis, 8(2):373–375, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223660.

Sarton:1926:BRBm

[568] George Sarton. Book review: Early Science in Oxford by Robert Theodore Gunther. Isis, 8(2):375–377, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223661.

Wieleitner:1926:BRBb

[569] H. Wieleitner. Book review: Val en Worp. Een Bijdrage tot de Geschiedenis der Mechanica van Aristoteles tot Newton by E. J. Dijksterhuis. Isis, 8(2):378–379, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223662.

Ruska:1926:BRBg

[570] Julius Ruska. Book review: Geschichte der Rube (Beta) als Kulturptlanze von den altesten Zeiten an bis zum Erscheinen von Achard’s Hauptwerk (1809) by Edmund O. von Lippmann. Isis, 8(2):379–380, May 1926. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/223663.

Dept:1926:BRB

[571] Gaston G. Dept. Book review: The New History and the Social Studies by Harry Elmer Barnes. Isis, 8(2):380–382, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223664.

Ruska:1926:BRBh

[572] Julius Ruska. Book review: Geschichte der Schrift by Hans Jensen. Isis, 8(2):382–385, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223665.

Sarton:1926:BRBn

[573] George Sarton. Book review: Pediatrics of the Past by John Ruhrah. Isis, 8(2):386–388, May 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223666.

Anonymous:1926:BMb

[574] Anonymous. Back matter. Isis, 8(2):??, May 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223667.

Anonymous:1926:FMc

[575] Anonymous. Front matter. Isis, 8(3):389–390, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223910.

Sarton:1926:SDP

[576] George Sarton. Le sentiment du passè. (French) [Feelings of the past]. Isis, 8(3):391–402, July 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223911.

Holmyard:1926:AQA

[577] E. J. Holmyard and E. J. H. Abu’ l-Qasim al-Iraqi. Isis, 8(3):403–426, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223912.

Metzger:1926:PMC

[578] Helene Metzger. La philosophie de la matière chez Stahl et ses disciples. (French) [The philosophy of matter of Stahl and his followers]. Isis, 8(3): 427–464, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223913.

vonLippmann:1926:QGC

[579] Edmund O. von Lippmann. Quellen zur Geschichte der Chemie und Alchemie in Italien. (German) [Sources on the history of chemistry and alchemy in Italy]. Isis, 8(3):465–476, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223914.

Wright:1926:PHG

[580] John Kirtland Wright. A plea for the history of geography. Isis, 8(3): 477–491, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223915.

Sarton:1926:NC

[581] George Sarton and L. Le Bondidier. Notes and correspondence. Isis, 8(3): 492–496, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223916.

G:1926:BRBe

[582] L. G. Book review: Opere di Agostino Bassi by Agostino Bassi. Isis, 8(3): 497–498, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223917.

May:1926:BRBc

[583] Raoul M. May. Book review: Recuerdos de mi vida by Santiago Ramon y Cajal; Charlas de cafe by Santiago Ramon y Cajal. Isis, 8(3):498–503, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223918.

G:1926:BRBf

[584] L. G. Book review: Scritti sulla storia della Astronomia antica by Giovanni Schiaparelli. Isis, 8(3):503–506, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223919.

Sarton:1926:BRBo

[585] George Sarton. Book review: The Assyrian Herbal by Reginald Campbell Thompson. Isis, 8(3):506–508, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223920.

Sarton:1926:BRBp

[586] George Sarton. Book review: Clues for the Arabian Influence on European Musical Theory by Henry George Farmer. Isis, 8(3):508–511, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223921.

Guinet:1926:BRBc

[587] L. Guinet. Book review: Histoire de la découverte de la photographie by Georges Potonniee. Isis, 8(3):511–513, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223922.

dIrsay:1926:BRBa

[588] Stephen d’Irsay. Book review: The Growth of Biology by William A. Locy. Isis, 8(3):513–514, July 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223923.

G:1926:BRBg

[589] L. G. Book review: L’évolution de l’ornithologie by Maurice Boubier. Isis, 8(3):515–517, July 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223924.

Wright:1926:BRBb

[590] J. K. Wright. Book review: Die Kartenwissenschaft: Forschung und Grundlagen zu einer Kartographie als Wissenschaft by Max Eckert. Isis, 8(3):517–519, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223925.

G:1926:BRBh

[591] L. G. Book review: La préhistoire orientale by Jacques de Morgan. Isis, 8(3):519–521, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223926.

dIrsay:1926:BRBb

[592] Stephen d’Irsay. Book review: Die Prostitution by Iwan Bloch; Georg Loewenstein. Isis, 8(3):521–525, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223927.

Sarton:1926:ECB

[593] George Sarton. Eighteenth critical bibliography of the history and philosophy of science and the history of civilization. (to November 1925). Isis, 8(3):526–654, July 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223928.

Anonymous:1926:EDP

[594] Anonymous. Errata: Drei planimetrische Aufgaben des arabischen Mathematikers Abu’l=Jud Muhammad ibn al-Lith. (German) [Errata: Three planimetric tasks of the Arab mathematician Abu’l-Jud Muhammad ibn al-Lith]. Isis, 8(3):655, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/ 223931. See [441].

Anonymous:1926:EIA

[595] Anonymous. Errata: Index of authors to volumes IV, V, and VI. Isis, 8 (3):655, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223930.

Anonymous:1926:EIL

[596] Anonymous. Errata: Ignatius Loyola. an attempt at an impartial biography. Isis, 8(3):655, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/ 223929. See [388].

Anonymous:1926:ENC

[597] Anonymous. Errata: Notes and correspondence. Isis, 8(3):655, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223932.

Anonymous:1926:ESC

[598] Anonymous. Errata: Sixteenth critical bibliography of the history and philosophy of science and of the history of civilization. Isis, 8(3):655, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223933.

Stewart:1926:AIE

[599] Elizabeth Gilpatrick Stewart. Authors’ index to the eighteenth bibliography. Isis, 8(3):656–667, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223934.

Anonymous:1926:BMc

[600] Anonymous. Back matter. Isis, 8(3):??, July 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223935.

Pearson:1926:RPM

[601] Karl Pearson, Abraham de Moivre, and R. C. Archibald. A rare pamphlet of moivre and some of his discoveries. Isis, 8(4):671–683, October 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223870.

Meyerhof:1926:NLH

[602] Max Meyerhof. New light on Hunain Ibn Ishaq and his period. Isis, 8 (4):685–724, October 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223871.

G:1926:NC

[603] L. G. and G. S. Notes and correspondence. Isis, 8(4):725, October 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223872.

Wright:1926:BRBc

[604] J. K. Wright and Ch. de la Ronciere. Book review: La carte de Christophe Colomb: The Map of Christopher Columbus by Charles de la Ronciere. Isis, 8(4):726–728, October 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223873.

Roback:1926:BRB

[605] A. A. Roback. Book review: Der Okkultismus in Urkunden by Max Dessoir; W. von Gulat-Wellenburg; Carl v. Klinckowstroem; Hans Rosenbusch; Richard Baerwald. Isis, 8(4):728–731, October 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223874.

Sarton:1926:NCB

[606] George Sarton. Nineteenth critical bibliography of the history and philosophy of science and of the history of civilization (to February 1926). Isis, 8(4):732–823, October 1926. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223875.

Stewart:1926:AIN

[607] Elizabeth Gilpatrick Stewart. Authors’ index to the nineteenth bibliography. Isis, 8(4):824–831, October 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223876.

Anonymous:1926:BTC

[608] Anonymous. Brief table of contents of volume VIII, (1926). Isis, 8(4):833– 834, October 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223877.

Anonymous:1926:BMd

[609] Anonymous. Back matter. Isis, 8(4):??, October 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223878.

Anonymous:1926:FMd

[610] Anonymous. Front matter. Isis, 8(4):??, October 1926. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223869.

Anonymous:1927:FMa

[611] Anonymous. Front matter. Isis, 9(1):3–5, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224166.

Anonymous:1927:LMH

[612] Anonymous. List of members of the History of Science Society elected from January 30, 1925 to July 31, 1926. Isis, 9(1):7–10, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224168.

Wieleitner:1927:WWA

[613] Heinrich Wieleitner. War die Wissenschaft der alten Aegypter wirklich nur praktisch?. (German) [Was the science of the Ancient Egyptians really just practical?]. Isis, 9(1):11–28, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224169.

Thorndike:1927:SMM

[614] Lynn Thorndike. Some minor medical works of the Florentine Renaissance. Isis, 9(1):29–43, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224170.

Rosenfeld:1927:TCN

[615] L. Rosenfeld. La théorie des couleurs de Newton et ses adversaires. (French) [Newton’s color theories and their adversaries]. Isis, 9(1):44– 65, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224171.

Bosmans:1927:ATJ

[616] H. Bosmans. Andre tacquet (s. j.) et son traite d’ ≪arithmetique theorique et pratique≫. Isis, 9(1):66–82, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224172.

Ruska:1927:CS

[617] Julius Ruska. Carl Schoy. Isis, 9(1):83–95, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224173.

Sarton:1927:MNE

[618] George Sarton. Moseley: The numbering of the elements. Isis, 9(1):96– 111, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224174.

Green-Armytage:1927:NC

[619] R. N. Green-Armytage, George Sarton, and H. E. B. Notes and correspondence. Isis, 9(1):112–119, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224175.

Sarton:1927:BRBa

[620] George Sarton. Book review: Introduction to Arithmetic by Nicomachus of Gerasa; Martin Luther D’Ooge; Frank Egleston Robbins; Louis Charles Karpinski. Isis, 9(1):120–123, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224176.

Sarton:1927:BRBb

[621] George Sarton. Book review: Anglo-Norman Lapidaries by Paul Studer; Joan Evans. Isis, 9(1):123–124, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224177.

G:1927:BRBa

[622] L. G. Book review: Histoire du cinematographe by G. Michel Coissac. Isis, 9(1):125–127, February 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224178.

Sarton:1927:BRBc

[623] George Sarton. Book review: Mémoires scientifiques by Paul Tannery. Isis, 9(1):127–130, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224179.

Meyerhof:1927:BRB

[624] Max Meyerhof. Book review: View of the Two Sanctuaries (Mir’at al-Haramain) by Rif’at Pasha, Ibrahim. Isis, 9(1):130–133, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224180.

Desch:1927:BRB

[625] Cecil H. Desch. Book review: Iron and Steel in the Industrial Revolution by Thomas S. Ashton. Isis, 9(1):133–134, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224181.

Sarton:1927:BRBd

[626] George Sarton. Book review: Printing Ink by Frank Bestow Wiborg. Isis, 9(1):134–138, February 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224182.

Sarton:1927:BRBe

[627] George Sarton. Book review: The American Indian by Clark, Wissler; The Relation of Nature to Man in Aboriginal America. Isis, 9(1):138– 142, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224183.

vKlinckowstroem:1927:BRB

[628] Carl v. Klinckowstroem. Book review: Aberglaube und Zauberei, von den allesten Zeiten bis in die Gegenwart by Alfred Lehmann. Isis, 9(1):142– 143, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224184.

G:1927:BRBb

[629] L. G. Book review: Les travailleurs du livre et du journal by Georges Renard. Isis, 9(1):143–144, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224185.

Sarton:1927:TCB

[630] George Sarton. Twentieth critical bibliography of the history and philosophy of science and of the history of civilization. (to may 1926). Isis, 9 (1):145–214, February 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224186.

Waterman:1927:AITa

[631] Dorothy Waterman. Authors’ index to the twentieth bibliography. Isis, 9(1):215–220, February 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224187.

Anonymous:1927:BMa

[632] Anonymous. Back matter. Isis, 9(1):??, February 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224188.

Anonymous:1927:FPI

[633] Anonymous. Frontispiece: Portrait of Isaac Newton to celebrate the second centenary of his death at Kensington on the 20th of March 1727. Isis, 9(1):??, February 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224167.

Anonymous:1927:FMb

[634] Anonymous. Front matter. Isis, 9(2):221–222, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/223710.

Brasch:1927:HSS

[635] Frederick E. Brasch. The History of Science Society: Report of the meetings for 1925 and 1926. Isis, 9(2):223–225, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/223711.

Sarton:1927:PVI

[636] George Sarton. Preface to volume IX. Isis, 9(2):226–233, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223712.

Yeldham:1927:AEE

[637] Florence A. Yeldham. The alleged early English version of Euclid. Isis, 9 (2):234–238, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223713.

Schoy:1927:AII

[638] C. Schoy and J. Drecker. ’Ali ibn ’Isa, Das Astrolab und sein Gebrauch. (German) [’Ali ibn ’Isa, the astrolabe and its use]. Isis, 9(2):239–254, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223714.

Drecker:1927:PCP

[639] J. Drecker. Das Planisphaerium des Claudius Ptolemaeus. (German) [The Planisphaerium of Claudius Ptolemy]. Isis, 9(2):255–278, June 1927. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/223715.

Mahnke:1927:LBS

[640] Dietrich Mahnke. Leibniz als begrunder der symbolischen mathematik. Isis, 9(2):279–293, June 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223716.

Metzger:1927:TCS

[641] Helene Metzger. La théorie de la composition des sels et la théorie de la combustion d’après Stahl et ses disciples. (French) [The theory of composition of salts and the theory of combustion according to Stahl and his followers]. Isis, 9(2):294–325, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223717.

Macdonald:1927:CRC

[642] D. B. Macdonald. Continuous re-creation and atomic time in Muslim scholastic theology. Isis, 9(2):326–344, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223718.

Rosenfeld:1927:PLD

[643] L. Rosenfeld. Le problème logique de la définition des nombres irrationnels. (German) [The logical problem of the definition of irrational numbers]. Isis, 9(2):345–358, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

223719.

Sarton:1927:MIHa

[644] George Sarton. Medallic illustrations of the history of science (first series: XIXth and XXth Centuries.) second article. Isis, 9(2):359–362, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223720.

Keen:1927:NC

[645] W. W. Keen and G. S. Notes and correspondence. Isis, 9(2):363–364, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/223721.

Sarton:1927:BRBf

[646] George Sarton. Book review: Honen the Buddhist Saint by Shunjo Hoin; Harper Havelock Coates; Ryugaku Ishizuka. Isis, 9(2):365–367, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/223722.

dIrsay:1927:BRBa

[647] Stephen d’Irsay. Book review: Die Handschrift des Schnitt- und Augenarztes Caspar Stromayr in Lindau am Bodensee by Walter von Brunn. Isis, 9(2):367–370, June 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/223723.

G:1927:BRBc

[648] L. G. Book review: Bibliographie générale des oeuvres de Blaise Pascal by Albert Maire; Blaise Pascal. Isis, 9(2):370–372, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/223724.

Sarton:1927:BRBg

[649] George Sarton. Book review: Scritti vari e lettere inedite nel centenario della nascita by Stanislao Cannizzaro. Isis, 9(2):372–374, June 1927. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/223725.

Sarton:1927:BRBh

[650] George Sarton. Book review: Tabula Smaragdina by Julius Ruska. Isis, 9 (2):375–377, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/223726.

vonKlinckowstroem:1927:BRB

[651] Carl von Klinckowstroem. Book review: Ruhmesblatter der Technik by Feldhaus, Franz M. Isis, 9(2):378–380, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223727.

dIrsay:1927:BRBb

[652] Stephen d’Irsay. Book review: Der Schwarze Tod, eine Chronik der Pest by Johannes Nohl. Isis, 9(2):380–382, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223728.

Anonymous:1927:CFE

[653] Anonymous. Contents of the first eight volumes of Isis. Isis, 9(2):383– 386, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/223729.

Anonymous:1927:BMb

[654] Anonymous. Back matter. Isis, 9(2):??, June 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/223730.

Anonymous:1927:FMc

[655] Anonymous. Front matter. Isis, 9(3):387–389, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330805.

Cajori:1927:EAP

[656] Florian Cajori. The earliest arithmetic published in America. Isis, 9 (3):391–401, December 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330806.

Turner:1927:PAE

[657] D. M. Turner. The philosophical aspect of education in science. Isis, 9 (3):402–419, December 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330807.

Sarton:1927:MIHb

[658] George Sarton. Medallic illustrations of the history of science (first series: XIXth and XXth Centuries third article). Isis, 9(3):420–423, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/330808.

Metzger:1927:NC

[659] Helene Metzger, Erich Adickes, Tenney L. Davis, Lynn Thorndike, Frederick E. Brasch, G. S., and L. Guinet. Notes and correspondence. Isis, 9 (3):424–429, December 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330809.

Sarton:1927:BRBi

[660] George Sarton. Book review: Die ersten gedruckten Pestschriften by Arnold C. Klebs; Karl Sudhoff. Isis, 9(3):430–433, December 1927. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/330810.

Sarton:1927:BRBj

[661] George Sarton. Book review: Musculorum humani corporis picturata dissectio by Ioannes Baptista Canano; Girolamo da Carpi; Harvey Cushing; Edward C. Streeter; Henry E. Sigerist. Isis, 9(3):433–434, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330811.

Sarton:1927:BRBk

[662] George Sarton. Book review: Correspondence of Descartes and Constantyn Huygens, 1635–1647 by Leon Roth; Descartes; Constantyn Huygens. Isis, 9(3):434–437, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330812.

Dingler:1927:BRBa

[663] Hugo Dingler. Book review: Galilei und sein Kampf für die Copernikanische Lehre by Emil Wohlwill. Isis, 9(3):437–439, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330813.

Davis:1927:BRB

[664] Tenney L. Davis. Book review: Alchemy in Early New England by George Lyman Kittredge. Isis, 9(3):440–445, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330814.

Sarton:1927:BRBl

[665] George Sarton. Book review: The Natural History of Ants by Rene Antoine Ferchault de Reaumur; William Morton Wheeler. Isis, 9(3):445– 447, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330815.

Wieleitner:1927:BRB

[666] H. Wieleitner. Book review: Vorlesungen über die Entwicklung der Mathematik im 19. Jahrhundert by Felix Klein; R. Courant; O. Neugebauer. Isis, 9(3):447–449, December 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330816.

Sarton:1927:BRBm

[667] George Sarton. Book review: Essai sur l’histoire antique d’Abyssinie by A. Kammerer. Isis, 9(3):449–450, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330817.

Thorndike:1927:BRB

[668] Lynn Thorndike. Book review: La Magie dans l’Egypte Antique de l’Ancien Empire jusqu’a l’Epoque Copte by Francois Lexa. Isis, 9(3):450– 452, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330818.

Sarton:1927:BRBn

[669] George Sarton. Book review: Anatomical Texts of the Earlier Middle Ages by George W. Corner. Isis, 9(3):452–456, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330819.

Macdonald:1927:BRB

[670] D. B. Macdonald. Book review: Rissalat al Tawhid by Cheikh Mohammed Abdou; B. Michel; Cheikh Moustapha Abdel Razik. Isis, 9(3):456–457, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330820.

Sarton:1927:BRBo

[671] George Sarton. Book review: Mappae arabicae by Konrad Miller. Isis, 9 (3):458–462, December 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330821.

Sarton:1927:BRBp

[672] George Sarton. Book review: Creation. A History of Non-Evolutionary Theories by Edwin Tenney Brewster. Isis, 9(3):462–465, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/330822.

Dingler:1927:BRBb

[673] Hugo Dingler. Book review: Geschichte der Atomistik vom Mittelalter bis Newton by Kurd Lasswitz. Isis, 9(3):465–467, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330823.

Sarton:1927:BRBq

[674] George Sarton. Book review: Les concepts scientifiques by Helene Metzger. Isis, 9(3):467–470, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

330824.

Metzger:1927:BRBa

[675] Helene Metzger. Book review: Identite et Realite by Emile Meyerson. Isis, 9(3):470–472, December 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330825.

Sarton:1927:BRBr

[676] George Sarton. Book review: Mémoires scientifiques by Paul Tannery; Gino Loria. Isis, 9(3):472–476, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330826.

Sarton:1927:BRBs

[677] George Sarton. Book review: Sternglaube und Sterndeutung by Franz Boll; Carl Bezold. Isis, 9(3):476–477, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330827.

Sarton:1927:BRBt

[678] George Sarton. Book review: Le retour éternel et la philosophie de la physique by Abel Rey. Isis, 9(3):477–479, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/330828.

Sarton:1927:BRBu

[679] George Sarton. Book review: Old Chemistries by Edgar Fahs Smith. Isis, 9(3):479–480, December 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330829.

May:1927:BRB

[680] Raoul M. May. Book review: Histoire de l’anatomie comparative by J. Chaine. Isis, 9(3):480–482, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/330830.

Metzger:1927:BRBb

[681] Helene Metzger. Book review: L’ame primitive by Lucien Levy Bruhl. Isis, 9(3):482–486, December 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330831.

dIrsay:1927:BRBc

[682] Stephen d’Irsay. Book review: History of Medicine by Max Neuburger; Ernest Playfair. Isis, 9(3):486–489, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330832.

Metzger:1927:BRBc

[683] Helene Metzger. Book review: Les courants de la pensée philosophique française by Andre Cresson. Isis, 9(3):489–490, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/330833.

Sarton:1927:TFC

[684] George Sarton. Twenty first critical bibliography of the history and philosophy of science and of the history of civilization (to October 1926). Isis, 9(3):491–604, December 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330834.

Waterman:1927:AITb

[685] Dorothy Waterman. Author’s index to the twenty first bibliography. Isis, 9(3):605–613, December 1927. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/330835.

Anonymous:1927:BMc

[686] Anonymous. Back matter. Isis, 9(3):??, December 1927. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330836.

Guinet:1928:IAV

[687] Leon Guinet. Index of authors to volumes VII, VIII and IX. Isis, 9(4): 617–714, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/330802.

Anonymous:1928:BTCa

[688] Anonymous. Brief table of contents of volume IX, (1927). Isis, 9(4):715– 716, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/330803.

Anonymous:1928:BM

[689] Anonymous. Back matter. Isis, 9(4):??, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330804.

Anonymous:1928:FMa

[690] Anonymous. Front matter. Isis, 9(4):??, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/330801.

Anonymous:1928:FMb

[691] Anonymous. Front matter. Isis, 10(1):1–5, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224733.

Anonymous:1928:LMH

[692] Anonymous. List of members of the History of Science Society elected from August 1, 1926 to July 31, 1927. Isis, 10(1):7–10, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224734.

Sarton:1928:S

[693] George Sarton. Spinoza 1632–1677–1927. Isis, 10(1):11–15, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224735.

Clarke:1928:BPC

[694] Frances Marguerite Clarke and David Eugene Smith. “Essay pour les Coniques” of Blaise Pascal. Isis, 10(1):16–20, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224736.

Karpinski:1928:TEG

[695] L. C. Karpinski and F. W. Kokomoor. The teaching of elementary geometry in the Seventeenth Century. Isis, 10(1):21–32, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224737.

Davis:1928:VBT

[696] Tenney L. Davis. The vicissitudes of Boerhaave’s textbook of chemistry. Isis, 10(1):33–46, March 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224738.

Barnes:1928:NC

[697] Harry Elmer Barnes, Edward Kremers, George Sarton, H. B. E., T. L. Davis, Lynn Thorndike, and L. G. Notes and correspondence. Isis, 10(1): 47–58, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224739.

A:1928:BRBa

[698] R. C. A. Book review: Early Editions of Euclid’s Elements by Charles Thomas-Stanford. Isis, 10(1):59–60, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224740.

Sarton:1928:BRBa

[699] George Sarton. Book review: The Thirteen Books of Euclid’s Elements by Thomas L. Heath; Heiberg. Isis, 10(1):60–62, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224741.

Sarton:1928:BRBb

[700] George Sarton. Book review: The Renaissance of the Twelfth Century by Charles Homer Haskins. Isis, 10(1):62–65, March 1928. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224742.

Sarton:1928:BRBc

[701] George Sarton. Book review: Islam and the Divine Comedy by Miguel Asin; Harold Sunderland. Isis, 10(1):65–69, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224743.

Sarton:1928:BRBd

[702] George Sarton. Book review: A Facsimile of the First Edition of the Pharmacopoeia Augustana. Isis, 10(1):69–71, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224744.

G:1928:BRBa

[703] L. G. Book review: Marcellin Berthelot (1827–1907) by A. Boutaric. Isis, 10(1):71–73, March 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224745.

Bloch:1928:BRBa

[704] Ernst Bloch. Book review: Lebenslinien by Wilhelm Ostwald. Isis, 10(1): 74–80, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224746.

Mikami:1928:BRBa

[705] Yoshio Mikami. Book review: Ogata Ko-an and Amori by K. Ogata. Isis, 10(1):80, March 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224747.

Smith:1928:BRB

[706] David Eugene Smith. Book review: Characteristics of Chinese Mathematics by Yoshio Mikami. Isis, 10(1):80–82, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224748.

Mikami:1928:BRBb

[707] Yoshio Mikami. Book review: Shinsen Yogaku Nempyo by Nyoden Otsuki. Isis, 10(1):82–83, March 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224749.

Sarton:1928:BRBe

[708] George Sarton. Book review: Scientific Japan, Past and Present. Isis, 10 (1):83–88, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224750.

Haskins:1928:BRB

[709] Charles H. Haskins. Book review: Introduction to the History of Science. Volume I, from Homer to Omar Khayyam by George Sarton. Isis, 10(1): 88–92, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224751.

Sarton:1928:BRBf

[710] George Sarton. Book review: Mémoires scientifiques. Volume VIII. Philosophie moderne. 1876–1903 by Paul Tannery; J. L. Heiberg. Isis, 10(1):92–94, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224752.

Sarton:1928:BRBg

[711] George Sarton. Book review: Catalogue of the Edward E. Ayer Ornithological Library by John Todd Zimmer. Isis, 10(1):94–95, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224753.

vKlinckowstroem:1928:BRB

[712] Graf Carl v. Klinckowstroem. Book review: Natur und Seele by Edgar Dacqué. Isis, 10(1):95–98, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224754.

Metzger:1928:BRBa

[713] Hélène Metzger. Book review: Le progrès de la conscience dans la philosophie occidentale by Leon Brunschvicg. Isis, 10(1):98–102, March 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224755.

Sarton:1928:TSC

[714] George Sarton. Twenty second critical bibliography of the history and philosophy of science and of the history of civilization (to may 1927). Isis, 10(1):103–312, March 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224756.

Waterman:1928:AITa

[715] Dorothy Waterman. Author’s index to the twenty second bibliography. Isis, 10(1):313–327, March 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224757.

Anonymous:1928:FMc

[716] Anonymous. Front matter. Isis, 10(2):329–331, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224500.

Brasch:1928:RME

[717] Frederick E. Brasch. Report of the meeting and exhibition commemorating the bi-centenary of the death of Sir Isaac Newton. Isis, 10(2):333–337, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224501.

Brasch:1928:PCM

[718] Frederick E. Brasch. Proceedings of the council meeting of the History of Science Society. Isis, 10(2):338–339, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224502.

Meyerhof:1928:ACM

[719] Max Meyerhof. An Arabic compendium of medico-philosophical definitions. Isis, 10(2):340–349, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224503.

Haskins:1928:AAM

[720] Charles H. Haskins. The “alchemy” ascribed to Michael Scot. Isis, 10(2): 350–359, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224504.

Thorndike:1928:CJG

[721] Lynn Thorndike. The clocks of Jacopo and Giovanni de’ Dondi. Isis, 10(2): 360–362, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224505.

Cajori:1928:CNA

[722] Florian Cajori. Ciruelo on the names “arithmetical” and “geometrical” proportions and progressions. Isis, 10(2):363–366, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224506.

Kokomoor:1928:DFS

[723] F. W. Kokomoor. The distinctive features of Seventeenth Century geometry. Isis, 10(2):367–415, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224507.

Rosenfeld:1928:RFS

[724] L. Rosenfeld. René-François de sluse et le problème des tangentes. Isis, 10 (2):416–434, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224508.

Sarton:1928:FE

[725] George Sarton and John Christian Oersted. The foundation of electromagnetism (1820). Isis, 10(2):435+437–444, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224509.

Stones:1928:AVM

[726] G. B. Stones. The atomic view of matter in the XVth, XVIth, and XVI-Ith Centuries. Isis, 10(2):445–465, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224510.

Walker:1928:RPB

[727] Helen M. Walker. The relation of Plana and Bravais to theory of correlation. Isis, 10(2):466–484, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224511.

Sarton:1928:MIH

[728] George Sarton. Medallic illustrations of the history of science (first series: XIXth and XXth Centuries) fourth article. Isis, 10(2):485–488, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224512.

Mourgue:1928:NC

[729] R. Mourgue, Lynn Thorndike, and George Sarton. Notes and correspondence. Isis, 10(2):489–493, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224513.

Sarton:1928:BRBh

[730] George Sarton. Book review: Three Essays of Abû ‘Othman ‘ Amr ibn Bah.r al-Jâh.iz. by al-Jah.iz; Joshua Finkel. Isis, 10(2):494–495, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224514.

Sarton:1928:BRBi

[731] George Sarton. Book review: Die trigonometrischen Lehren des persischen Astronomen Abu l-Raih.ân Muh.. ibn Ah.mad al-Bîrûnî dargestellt nach al-Qânûn al-Mas‘ûdî by Carl Schoy; Julius Ruska; Heinrich Wieleitner. Isis, 10(2):495–497, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224515.

Macdonald:1928:BRB

[732] Duncan B. Macdonald. Book review: Tahafot al-Falasifat by Algazel; Maurice Bouyges. Isis, 10(2):497–499, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224516.

Wieleitner:1928:BRBa

[733] H. Wieleitner. Book review: Geschichte der neusprachlichen wissenschaftlichen Literatur. III. Band: Galilei und seine Zeit by Leonardo Olschki. Isis, 10(2):499–501, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224517.

Metzger:1928:BRBb

[734] Hélène Metzger. Book review: Les historiens de l’esprit humain by Raymond Lenoir. Isis, 10(2):501–502, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224518.

May:1928:BRB

[735] Raoul M. May. Book review: Études d’histoire des sciences naturelles by Henry Daudin; De Linné à Jussieu; Cuvier; Lamarck. Isis, 10(2): 502–505, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224519.

Wieleitner:1928:BRBb

[736] H. Wieleitner. Book review: Vorlesungen über die Entwicklung der Mathematik im 19. Jahrhundert. Teil II. Die Grundbegriffe der Invariantentheorie und ihr Eindringen in die mathematische Physik by Felix Klein; R. Courant; St. Cohn-Vossen. Isis, 10(2):505–506, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224520.

Bloch:1928:BRBb

[737] Ernst Bloch. Book review: Lebenslinien. Eine Selbstbiographie by Wilhelm Ostwald. Isis, 10(2):506–509, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224521.

G:1928:BRBb

[738] L. G. Book review: Recherches sur les sources de la légende de Pythagore by Isidore Lévy; La légende de Pythagore de Grèce en Palestine. Isis, 10 (2):509–510, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224522.

S:1928:BRBa

[739] G. S. Book review: Insect-Musicians and Cricket Champions of China by Berthold Laufer. Isis, 10(2):510–511, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224523.

Sarton:1928:BRBj

[740] George Sarton. Book review: La Chasse et les Sports chez les Arabes by Louis Mercier. Isis, 10(2):511–513, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224524.

Langford:1928:BRB

[741] C. H. Langford. Book review: Principia Mathematica by A. N. Whitehead; Bertrand Russell. Isis, 10(2):513–519, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224525.

S:1928:BRBb

[742] G. S. Book review: The Herbal in Antiquity by Charles Singer. Isis, 10(2): 519–521, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224526.

Leake:1928:BRB

[743] Chauncey D. Leake. Book review: The Evolution of Anatomy by Charles Singer. Isis, 10(2):521–524, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224527.

A:1928:BRBb

[744] R. C. A. Book review: La Musique des Incas et ses survivances, Vol. I by Raoul d’Harcourt; Marg. Béclard d’Harcourt. Isis, 10(2):524–526, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224528.

Anonymous:1928:BTCb

[745] Anonymous. Brief table of contents of volume X (1928-1). Isis, 10(2): 527–528, June 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224529.

Anonymous:1928:FMd

[746] Anonymous. Front matter. Isis, 11(1):1–4, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224758.

Sarton:1928:PVE

[747] George Sarton. Preface to volume eleven. Isis, 11(1):5–14, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224759.

Drecker:1928:JPS

[748] J. Drecker. Des Johannes Philoponos Schrift über das Astrolab. (German) [John Philoponus’ treatise on the astrolabe]. Isis, 11(1):15–44, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224760.

Karpinski:1928:TCA

[749] Louis C. Karpinski and E. G. R. Waters. A Thirteenth Century algorism in French verse. Isis, 11(1):45–84, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224761.

Kokomoor:1928:TEG

[750] F. W. Kokomoor. The teaching of elementary geometry in the Seventeenth Century. Isis, 11(1):85–110, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224762.

Rosenfeld:1928:PCE

[751] L. Rosenfeld. Le premier conflit entre la théorie ondulatoire et la théorie corpusculaire de la lumière. (French) [The first conflict between wave and particle theories of light]. Isis, 11(1):111–122, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224763.

Mikami:1928:CJC

[752] Yoshio Mikami. The ch’ou-jen chuan of Yüan Yüan. Isis, 11(1):123–126, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224764.

Macdonald:1928:NC

[753] Duncan B. Macdonald and George Sarton. Notes and correspondence. Isis, 11(1):127–128, September 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224765.

Sarton:1928:BRBk

[754] George Sarton. Book review: Chemistry in ’Iraq and Persia in the Tenth Century A.D. by H. E. Stapleton; R. F. Azo; M. Hidâyat H .usain. Isis, 11 (1):129–134, September 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224766.

Sarton:1928:BRBl

[755] George Sarton. Book review: Avicennae De congelatione et conglutinatione lapidum by E. J. Holmyard; D. C. Mandeville. Isis, 11(1):134–135, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224767.

Wright:1928:BRB

[756] J. K. Wright. Book review: Il Milione by Marco Polo; Luigi Foscolo Benedetto. Isis, 11(1):135–138, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224768.

Sarton:1928:BRBm

[757] George Sarton. Book review: The Opus Majus by Roger Bacon; Robert Belle Burke. Isis, 11(1):138–141, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224769.

Sarton:1928:BRBn

[758] George Sarton. Book review: The Cipher of Roger Bacon by William Romaine Newbold; Roland Grubb Kent. Isis, 11(1):141–145, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224770.

Sarton:1928:BRBo

[759] George Sarton. Book review: Forerunners of Saint Francis and Other Studies by Ellen Scott Davison; Gertrude R. B. Richards. Isis, 11(1):145– 147, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224771.

Finkel:1928:BRB

[760] Joshua Finkel. Book review: Ozar Massaoth by J. D. Eisenstein. Isis, 11 (1):147–149, September 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224772.

Metzger:1928:BRBc

[761] Hélène Metzger. Book review: Une nouvelle philosophie des sciences — le causalisme de M. Emile Meyerson by André Metz. Isis, 11(1):149–150, September 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224773.

Sarton:1928:TTC

[762] George Sarton. Twenty third critical bibliography of the history and philosophy of science and of the history of civilization (to October 1927). Isis, 11(1):151–272, September 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224774.

Waterman:1928:AITb

[763] Dorothy Waterman. Author’s index to the twenty third bibliography. Isis, 11(1):273–282, September 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224775.

Anonymous:1928:FMe

[764] Anonymous. Front matter. Isis, 11(2):283–284, December 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224422.

Stephenson:1928:ZSN

[765] J. Stephenson. The zoological section of the Nuzhatu-l-Qulûb. Isis, 11 (2):285–315, December 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224423.

Cajori:1928:RHB

[766] Florian Cajori. A revaluation of Harriot’s Artis Analyticae Praxis. Isis, 11(2):316–324, December 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224424.

Winthrop:1928:SNB

[767] John Winthrop, Geo Starkey, and C. A. Browne. Scientific notes from the books and letters of John Winthrop, Jr., (1606–1676). Isis, 11(2):325– 342, December 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224425.

Johnsson:1928:PDF

[768] J. W. S. Johnsson and A. Krekich. La peste de 1763 en Dalmatie. (French) [The plague of 1763 in Dalmatia]. Isis, 11(2):343–363, December 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224426.

Struik:1928:CBP

[769] D. J. Struik and Ruth Struik. Cauchy and Bolzano in Prague. Isis, 11 (2):364–366, December 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224427.

Raeder:1928:JLH

[770] Hans Ræder. Johan Ludvig Heiberg. 27/11 1854–4/1 1928. Isis, 11(2): 367–374, December 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224428.

Browne:1928:EFS

[771] C. A. Browne. Edgar fahs Smith. Isis, 11(2):375–384, December 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224429.

Sarton:1928:BRBp

[772] George Sarton. Book review: Le comput manuel de Magister Anianus by David Eugene Smith. Isis, 11(2):385–387, December 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224430.

Sarton:1928:BRBq

[773] George Sarton. Book review: [Sir Isaac Newton (1727–1927)]. A Bicentenary Evaluation of His Work; F. E. Brasch; [Isaac Newton (1727–1927)]. A Memorial Volume Edited for the Mathematical Association by W. J. Greenstreet; Un double centenaire, Newton et Laplace by Emile Picard. Isis, 11(2):387–393, December 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224431.

G:1928:BRBc

[774] L. G. Book review: Jean Sylvain Bailly, premier maire de Paris by Fernand Laurent. Isis, 11(2):393–395, December 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224432.

Archibald:1928:BRB

[775] Raymond Clare Archibald. Book review: La Science Égyptienne. L’Arithmétique au Moyen Empire by O. Gillain. Isis, 11(2):395–398, December 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224433.

Transehe:1928:BRB

[776] N. Transehe. Book review: Istoria geograficheskago oznakomleniya s Yakutskim Kraem (History of the Geographical Acquaintance with the Yakutsk Region) by L. S. Berg. Isis, 11(2):398–399, December 1928. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/224434.

Sarton:1928:BRBr

[777] George Sarton. Book review: The Coral Reef Problem by William Morris Davis. Isis, 11(2):399–401, December 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224435.

Sarton:1928:BRBs

[778] George Sarton. Book review: Orokaiva Magic by F. E. Williams. Isis, 11 (2):401–403, December 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224436.

Sarton:1928:TFC

[779] George Sarton. Twenty fourth critical bibliography of the history and philosophy of science and of the history of civilization (to may 1928). Isis, 11(2):404–568, December 1928. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224437.

Waterman:1928:AITc

[780] Dorothy Waterman. Authors’ index to the twenty fourth bibliography. Isis, 11(2):569–580, December 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224438.

Anonymous:1928:BTCc

[781] Anonymous. Brief table of contents of volume XI (1928-2). Isis, 11(2):581– 582, December 1928. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224439.

Anonymous:1929:VIa

[782] Anonymous. Volume information. Isis, 12(1):1, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224403.

Anonymous:1929:FMa

[783] Anonymous. Front matter. Isis, 12(1):3–4, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224404.

Sarton:1929:PVT

[784] George Sarton. Preface to volume twelve. Isis, 12(1):5–9, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224405.

ofCremona:1929:PCT

[785] Gerard of Cremona and Robert Steele. Practical chemistry in the Twelfth Century Rasis de aluminibus et salibus. Isis, 12(1):10–46, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224406.

Johnsson:1929:PDIa

[786] J. W. S. Johnsson. La peste de 1763 en dalmatie. II. (French) [The plague of 1763 in Dalmatia. II]. Isis, 12(1):47–87, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224407.

Rome:1929:RPH

[787] A. Rome. Le R. P. Henri Bosmans, S.J. (1852–1928). Isis, 12(1):88– 112, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224408.

Meyerhof:1929:NQM

[788] Max Meyerhof. Notes sur quelques médecins juifs égyptiens qui se sont illustrés à l’époque arabe. (French) [Notes on some Egyptian Jewish physicians who are known from the Arab era]. Isis, 12(1):113–131, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224409.

Ganguli:1929:NIM

[789] Sâradâkânta Gâṅguli. Notes on Indian mathematics. A criticism of George Rusby Kaye’s interpretation. Isis, 12(1):132–145, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224410.

Sarton:1929:MIHa

[790] George Sarton. Medallic illustrations of the history of science (first series: XIXth and XXth Centuries) fifth article. Isis, 12(1):146–148, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224411.

Barnes:1929:NC

[791] Harry Elmer Barnes, George Sarton, and Edward Kremers. Notes and correspondance. Isis, 12(1):149–151, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224412.

Sarton:1929:BSM

[792] George Sarton. The ≪Book of Surgeries≫ of Master Thomas Scellinck from Thienen. Isis, 12(1):152–153, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224413.

Sarton:1929:BRBa

[793] George Sarton. Book review: The rôle of Scientific Societies in the Seventeenth Century by Martha Ornstein. Isis, 12(1):154–156, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224414.

Bevenot:1929:BRBa

[794] Hugh G. Bévenot. Book review: Sibyllinischer Sternkampf und Phaëton, in naturgeschichtlicher Beleuchtung by Franz Xaver Kugler. Isis, 12(1): 156–157, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224415.

Sarton:1929:BRBb

[795] George Sarton. Book review: The Bakhshâlî Manuscript by George Rusby Kaye. Isis, 12(1):157–161, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224416.

S:1929:BRBa

[796] G. S. Book review: Anatomie mandchoue by Abr. Clod-Hansen; Vilhelm Thomsen. Isis, 12(1):161, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224417.

Sarton:1929:BRBc

[797] George Sarton. Book review: Disertaciones y opúsculos by Julián Ribera; Julián Tarragó. Isis, 12(1):161–163, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224418.

Karpinski:1929:BRB

[798] Louis C. Karpinski. Book review: The Early Mathematical Sciences in North and South America by Florian Cajori. Isis, 12(1):163–165, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224419.

Sarton:1929:BRBd

[799] George Sarton. Book review: Catalogue des Manuscrits Alchimiques Grecs by J. Bidez; F. Cumont; A. Delatte; O. Lagercrantz; J. Ruska; C. O. Zuretti; A. Severyns; Michel Psellus; Joseph Bidez. Isis, 12(1):165–168, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224420.

Sarton:1929:BRBe

[800] George Sarton. Book review: Catalogue of Latin and Vernacular Alchemical Manuscripts in Great Britain and Ireland, Dating from before the XVIth Century. Volume I by Dorothea Waley Singer; Annie Anderson. Isis, 12(1):168–169, February 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224421.

Anonymous:1929:FMb

[801] Anonymous. Front matter. Isis, 12(2):171–172, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224781.

Sarton:1929:RFA

[802] George Sarton and Frederic E. Brasch. Report of the Fifth Annual Meeting of the History of Science Society. Isis, 12(2):173–178, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224782.

Anonymous:1929:FLM

[803] Anonymous. Fourth list of members of the History of Science Society. Isis, 12(2):179–183, May 1929. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224783.

Thompson:1929:IAS

[804] James Westfall Thompson. The introduction of Arabic science into lorraine in the Tenth Century. Isis, 12(2):184–193, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224784.

Waters:1929:FCF

[805] E. G. R. Waters. A Fifteenth Century French algorism from Liége. Isis, 12(2):194–236, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/224785.

Pelseneer:1929:LIN

[806] Jean Pelseneer. Une lettre inédite de Newton. Isis, 12(2):237–254, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224786.

Johnsson:1929:PDIb

[807] J. W. S. Johnsson. La peste de 1763 en Dalmatie. III. Appendice. (French) [The plague of 1763 in Dalmatia. III. Appendix]. Isis, 12(2):255–286, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224787.

Rainoff:1929:WLF

[808] T. J. Rainoff. Wave-like fluctuations of creative productivity in the development of West-European physics in the Eighteenth and Nineteenth Centuries. Isis, 12(2):287–319, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224788.

Sarton:1929:MIHb

[809] George Sarton. Medallic illustrations of the history of science (first series: XIXth and XXth Centuries) sixth article. Isis, 12(2):320–322, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224789.

Loria:1929:NC

[810] Gino Loria and George Sarton. Notes and correspondence. Isis, 12(2): 323–325, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224790.

Sarton:1929:BRBf

[811] George Sarton. Book review: Further Correspondence of John Ray by Robert W. Theodore Gunther. Isis, 12(2):326–328, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224791.

Sarton:1929:BRBg

[812] George Sarton. Book review: Anecdota atheniensa. Tome I by Armand Delatte. Isis, 12(2):328–330, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224792.

Laufer:1929:BRB

[813] B. Laufer. Book review: A Study of Chinese Alchemy by Obed Simon Johnson. Isis, 12(2):330–332, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224793.

Sarton:1929:BRBh

[814] George Sarton. Book review: A History of Mathematical Notations by Florian Cajori. Isis, 12(2):332–336, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224794.

Sarton:1929:BRBi

[815] George Sarton. Book review: The History of Biology by Erik Nordenskiöld; Leonard Bucknall Eyre. Isis, 12(2):336–340, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224795.

Sarton:1929:BRBj

[816] George Sarton. Book review: Barnacles in Nature and in Myth by Edward Heron-Allen. Isis, 12(2):340–341, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224796.

Sarton:1929:BRBk

[817] George Sarton. Book review: This Puzzling Planet by Edwin Tenney Brewster. Isis, 12(2):341–343, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224797.

Metzger:1929:BRB

[818] Hélène Metzger. Book review: Les fonctions mentales chez les sociétés inférieures by Lucien Lévy Bruhl; Les non-civilisés et nous, différence irréductible ou identité foncière by Raoul Allier; La raison primitive by Olivier Leroy. Isis, 12(2):343–347, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224798.

Sarton:1929:TFC

[819] George Sarton. Twenty fifth critical bibliography of the history and philosophy of science and of the history of civilization (to October 1928). Isis, 12(2):348–441, May 1929. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224799.

Siegel:1929:AITa

[820] Frances Siegel. Author’s index to the twenty fifth bibliography. Isis, 12(2): 442–450, May 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224800.

Anonymous:1929:FMc

[821] Anonymous. Front matter. Isis, 12(3):451, December 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224468.

Gandz:1929:OAG

[822] Solomon Gandz. The origin of angle-geometry. Isis, 12(3):452–481, December 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/224469.

Cajori:1929:HDH

[823] Florian Cajori. History of determinations of the heights of mountains. Isis, 12(3):482–514, December 1929. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224470.

Guinet:1929:IAV

[824] Leon Guinet. Index of authors to volumes X, XI and XII. Isis, 12(3):515– 585, December 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224471.

Anonymous:1929:BTC

[825] Anonymous. Brief table of contents of volume XII (1929). Isis, 12(3):586– 587, December 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224472.

Anonymous:1929:FMd

[826] Anonymous. Front matter. Isis, 13(1):1–4, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224592.

Singer:1929:MSA

[827] Dorothea Waley Singer. Michael Scot and alchemy. Isis, 13(1):5–15, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224593.

Jenner:1929:ULE

[828] Edward Jenner and G. S. An unpublished letter of Edward Jenner (1809). Isis, 13(1):16–17, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

224594.

Sarton:1929:DLC

[829] G. Sarton, J. R. Mayer, J. P. Joule, and Sadi Carnot. The discovery of the law of conservation of energy. Isis, 13(1):18–44, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224595.

Rosenfeld:1929:EHG

[830] L. Rosenfeld. Edmund Hoppe (geb. 1854, 02, 25, gest. 1928, 08, 12.). (German). Isis, 13(1):45–50, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224596.

Thorndike:1929:CCH

[831] Lynn Thorndike. Of the cylinder called the horologe of travelers. Isis, 13(1):51–52, September 1929. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224597.

Thorndike:1929:VLM

[832] Lynn Thorndike. Vatican Latin manuscripts in the history of science and medicine. Isis, 13(1):53–102, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224598.

Thorndike:1929:NC

[833] Lynn Thorndike and George Sarton. Notes and correspondence. Isis, 13 (1):103–105, September 1929. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224599.

Sarton:1929:BRBl

[834] George Sarton. Book review: The book of the ten treatises on the eye ascribed to H .unain ibn Ish.âq (809–877 A.D.) by H .unain ibn Ish.âq; Max Meyerhof. Isis, 13(1):106–109, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224600.

Thorndike:1929:BRBa

[835] Lynn Thorndike. Book review: Fasc. IX, De retardatione accidentium senectutis cum aliis opusculis de rebus medicinalibus by Rogeri Baconi; A. G. Little; E. Withington. Isis, 13(1):110–111, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224601.

Sarton:1929:BRBm

[836] George Sarton. Book review: Die Instrumente der Sternwarte zu Marâgha nach den Mitteilungen von Al ‘Urd.î by Hugo J. Seemann. Isis, 13(1):111– 113, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224602.

Sarton:1929:BRBn

[837] George Sarton. Book review: Zwölf Bücher vom Berg- und Hüttenwesen in denen die ämter, Instrumente, Maschinen und alle Dinge, die zum Bergund Hüttenwesen gehören, nicht nur aufs deutlichste beschrieben, sondern auch durch Abbildungen, die am gehörigen Orte eingefügt sind, unter Angabe der lateinischen und deutschen Bezeichnungen aufs klarste vor Augen gestellt werden by Georg Agricola; Carl Schiffner; Ernst Darmstaedter; Paul Knauth; Wilhelm Pieper; Friedrich Schumacher; Victor Tafel; Emil Treptow; Erich Wandhoff; De re metallica by Georgius Agricola; Herbert Clark Hoover; Lou Henry Hoover. Isis, 13(1):113–116, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224603.

W:1929:BRB

[838] J. K. W. Book review: Richard Hakluyt and the English Voyages by George Bruner Parks; James A. Williamson. Isis, 13(1):116–118, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/224604.

S:1929:BRBb

[839] G. S. Book review: Neue Astronomie by Johannes Kepler; Max Caspar. Isis, 13(1):118, September 1929. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224605.

S:1929:BRBc

[840] G. S. Book review: Early Science in Oxford. Vol. IV by Robert Theodore Gunther. Isis, 13(1):118–119, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224606.

Sarton:1929:BRBo

[841] George Sarton. Book review: The Chequered Career of Ferdinand Rudolph Hassler, First Superintendent of the United States Coast Survey by Florian Cajori. Isis, 13(1):119–121, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224607.

Bevenot:1929:BRBb

[842] H. G. Bévenot. Book review: Grundriss der Antiken Zeitrechnung by Wm. Kubitschek. Isis, 13(1):121–122, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224608.

Sarton:1929:BRBp

[843] George Sarton. Book review: The Greek Atomists and Epicurus by Cyril Bailey. Isis, 13(1):123–125, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224609.

Sarton:1929:BRBq

[844] George Sarton. Book review: Das Geschütz im Mittelalter by Bernhard Rathgen. Isis, 13(1):125–127, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224610.

S:1929:BRBd

[845] G. S. Book review: Introduction à l’astronomie nautique arabe by Gabriel Ferrand. Isis, 13(1):127–128, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/224611.

S:1929:BRBe

[846] G. S. Book review: Las matemáticas en la Biblioteca del Escorial by José A. Sánchez Pérez. Isis, 13(1):128–129, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/224612.

S:1929:BRBf

[847] G. S. Book review: A History of Mathematical Notations. Vol. 2. Notations Mainly in Higher Mathematics by Florian Cajori. Isis, 13(1):129– 130, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/224613.

Sarton:1929:BRBr

[848] George Sarton. Book review: A Source Book in Astronomy by Harlow Shapley; Helen E. Howarth. Isis, 13(1):130–134, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224614.

May:1929:BRB

[849] R. M. May. Book review: Études sur la chimie physiologique de la peau by L. H. Dejust; J. Verne; Raoul Combes; Maurice Parat; Achille Urbain; R. Dijarric de la Rivière; L. de Saint-Rat. Isis, 13(1):134–135, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/224615.

Sarton:1929:BRBs

[850] George Sarton. Book review: Civilization by James Harvey Robinson. Isis, 13(1):136–137, September 1929. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224616.

Sarton:1929:BRBt

[851] George Sarton. Book review: An Introduction to the History of Medicine by Fielding H. Garrison. Isis, 13(1):137–138, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224617.

Thorndike:1929:BRBb

[852] Lynn Thorndike. Book review: Witchcraft in Old and New England by George Lyman Kittredge. Isis, 13(1):138–141, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224618.

Sarton:1929:TSC

[853] George Sarton. Twenty-sixth critical bibliography of the history and philosophy of science and of the history of civilization (to March 1, 1929). Isis, 13(1):142–259, September 1929. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224619.

Siegel:1929:AITb

[854] Frances Siegel. Authors’ index to the twenty-sixth bibliography. Isis, 13 (1):260–268, September 1929. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/224620.

Anonymous:1929:BM

[855] Anonymous. Back matter. Isis, 13(1):??, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/224621.

Anonymous:1929:VIb

[856] Anonymous. Volume information. Isis, 13(1):??, September 1929. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/224591.